### v1

# Conditional regression for the Nonlinear Single-Variable Model

Abstract
Several statistical models for regression of a function $F$ on $\mathbb{R}^{d}$ without the statistical and computational curse of dimensionality exist, for example by imposing and exploiting geometric assumptions on the distribution of the data (e.g. that its support is low-dimensional), or strong smoothness assumptions on $F$, or a special structure $F$. Among the latter, compositional models assume $F=f\circ g$ with $g$ mapping to $\mathbb{R}^{r}$ with $r\ll d$, have been studied, and include classical single- and multi-index models and recent works on neural networks. While the case where $g$ is linear is rather well-understood, much less is known when $g$ is nonlinear, and in particular for which $g$’s the curse of dimensionality in estimating $F$, or both $f$ and $g$, may be circumvented. In this paper, we consider a model $F(X):=f(\Pi_{\gamma}X)$ where $\Pi_{\gamma}:\mathbb{R}^{d}\to[0,\textrm{len}_{\gamma}]$ is the closest-point projection onto the parameter of a regular curve $\gamma:[0,\textrm{len}_{\gamma}]\to\mathbb{R}^{d}$ and $f:[0,\textrm{len}_{\gamma}]\to\mathbb{R}^{1}$. The input data $X$ is not low-dimensional, far from $\gamma$, conditioned on $\Pi_{\gamma}(X)$ being well-defined. The distribution of the data, $\gamma$ and $f$ are unknown. This model is a natural nonlinear generalization of the single-index model, which corresponds to $\gamma$ being a line. We propose a nonparametric estimator, based on conditional regression, and show that under suitable assumptions, the strongest of which being that $f$ is coarsely monotone, it can achieve the one-dimensional optimal min-max rate for non-parametric regression, up to the level of noise in the observations, and be constructed in time $\mathcal{O}(d^{2}n\log n)$. All the constants in the learning bounds, in the minimal number of samples required for our bounds to hold, and in the computational complexity are at most low-order polynomials in $d$.
Keywords: Nonparametric regression; compositional models; single-index model.
1 Introduction
We consider the standard regression problem of estimating a function $F:\mathbb{R}^{d}\to\mathbb{R}^{1}$ from $n$ samples $\{(X_{i},Y_{i})\}_{i=1}^{n}$, where $X_{i}$’s are i.i.d. realizations of a predictor variable $X\in\mathbb{R}^{d}$ with distribution $\rho_{X}$, and $\zeta_{i}$ are realizations (independent among themselves and of the $X_{i}$’s), of a random variable $\zeta$ modeling observational noise, and
$Y_{i}=F(X_{i})+\zeta_{i}\,.$
In the general nonparametric and distribution-free setup where we only know that $F\in\mathcal{C}^{s}(\mathbb{R}^{d})$ is Hölder continuous with exponent $s>0$ and, say, compacted supported, the min-max nonparametric rate for estimating $F$ in $L^{2}(\rho)$ is $n^{-\frac{s}{2s+d}}$, see (Binev et al., 2023; Györfi et al., 2002) and references therein. Kernel estimators attain this learning rate with properly chosen bandwidth and kernel, as so do a variety of other well-understood estimators based on Fourier or multiscale decompositions, see (Györfi et al., 2002; Binev et al., 2005, 2007) and the numerous references therein. This rate deteriorates dramatically as the dimension $d$ of the ambient space increases: this is an instance of the curse of dimensionality. As no estimator can achieve a faster learning rate, in the min-max sense, for all such functions, and yet in many applications $d$ is very large, it is of interest to consider special classes of regression problems where the curse of dimensionality may be avoided.
Intrinsically low-dimensional models. In one direction, one may make geometric assumptions about the distribution $\rho_{X}$ of the input data $X_{i}$, for example that $\rho_{X}$ is supported on a low-dimensional manifold $\mathcal{M}$, say of dimension $r\ll d$, while $F$ is, say, a generic $s$-Hölder function on $\mathcal{M}$. In these settings, estimators exist that converge at an optimal rate concerning the intrinsic dimension $r$, see (Bickel and Li, 2007; Kpotufe, 2011; Kpotufe and Garg, 2013; Liao et al., 2016, 2022; Liu et al., 2024a), with at least some of them with associated fast algorithms, for example the estimators in (Liao et al., 2016, 2022) can be constructed in time $\mathcal{O}(C^{r}d^{s}n\log n)$ and they achieve, adaptively, min-max optimal rate on large function spaces with unknown regularity possibly varying across locations and scales.
Functions with high degree of regularity. In a different direction, one may consider function classes with smaller complexity (e.g., as measured by metric entropy) than $s$-Hölder functions on $\mathbb{R}^{d}$. A straightforward but highly limiting assumption is that the Hölder exponent $s$ is proportional to $d$ so that the min-max rate above is independent of $d$. Larger function classes are obtained by imposing integrability of the Fourier transform (Barron, 1993), or strong mixed-smoothness (Strömberg, 1998), for example by using highly anisotropic Besov spaces. These choices however still require a degree of smoothness increasing with $d$ in order to avoid the curse of dimensionality. For example in these anisotropic Besov spaces, functions need to have $\mathcal{O}(1)$ axis-oriented directions in which they have “regular, $\mathcal{O}(1)$ smoothness”, and have “$\mathcal{O}(d)$ smoothness” in all other directions. In the context of deep neural networks, Suzuki and Nitanda (2024) demonstrate that functions in this class can be well-approximated well, with suitable architectures, in a way that avoids the curse of dimensionality if the degree of anisotropic smoothness is sufficiently large, while leaving open the aspects of learning and optimization.
Compositional models. Yet other functional classes are obtained by imposing structural assumptions, which are often in the form of compositional models. For example assume that
$F=f\circ g\quad,\text{ with }g:\mathbb{R}^{d}\to\mathbb{R}^{r}\text{ and }f:\mathbb{R}^{r}\to\mathbb{R}^{1}$
having some Hölder regularity, and $r\leq d$. For general function compositions, (Juditsky et al., 2009) proves that in some situations, when $g$ is sufficiently smooth, there is an improved min-max nonparametric rate of estimation of $F=f\circ g$, but still subject to the curse of dimensionality. Recent works combine anisotropic smoothness with composability, especially in the context of approximators or estimators constructed with deep neural networks. In this direction, (Schmidt-Hieber, 2020; Shen et al., 2021) consider spaces of functions that are compositions of low-dimensional functions with anisotropic smoothness conditions, studying the dependence of approximability and learning risk on the dimensionality of the spaces involved and the smoothness of the corresponding functions. Unfortunately the inner-most such function is defined on the original high-dimensional space $\mathbb{R}^{d}$, so in general these results do not circumvent the curse of dimensionality, unless again the (anisotropic, in general) regularity of such function scales with $d$. Another direction is to design different types of neural networks that exploit nonlinear compositions, see (Lai and Shen, 2021; Liu et al., 2024b). Most of these results only address whether a function can well-approximated by a neural network, and do not cover the learning problem nor address computational aspects, in particular whether some optimization algorithm can find (efficiently?) the parameters of the desired estimator. Another aspect that is often left unaddressed is the dependency of constants on the ambient dimension, which is unfortunately often exponential: see Shamir (2020) for a discussion of some of these aspects, and Shen et al. (2021) for bounds on such constants in some of these cases. In the parametric setting, models where $g$ is a polynomial and $f$ is a function in a known finite-dimensional space have been considered; for example Wang et al. (2024); Lee et al. (2024) (and related work referenced therein) consider special classes of polynomials $g$ for which a customized layer-by-layer training of a neural network leads to estimators of $F$ that are not cursed by the dimensionality.
Single- and multi-index models. Classical examples of structural assumptions based on function composition for which the curse of dimensionality can provably be avoided include single- and multi-index models, as well as generalized linear models (Stone, 1982; Hastie and Tibshirani, 2014; Breiman and Friedman, 1985; Horowitz and Mammen, 2007). In the case when $g=G:\mathbb{R}^{d}\to\mathbb{R}^{r}$ is a linear operator, this is called the multi-index model, and it implies that the function $F\in\mathcal{C}^{s}(\mathbb{R}^{d})$ only depends on a small number $r$ of linear features: $F(x)=f(Gx)$ for some link function $f\in\mathcal{C}^{s}(\mathbb{R}^{r})$, matrix $G:\mathbb{R}^{d}\rightarrow\mathbb{R}^{r}$, and the projection of $X$ on the range of $G$ is sufficient for regression. The particular case $r=1$ is called the single-index model, where the function $F$ has the structure $F(x)=f(\langle v,x\rangle)$ for some unknown index vector $v\in\mathbb{S}^{d-1}$ and unknown link function $f\in\mathcal{C}^{s}(\mathbb{R}^{1})$. These models have been intensively studied: Stone (1982) conjectured that the min-max rate for regression of single-index models is $n^{-\frac{s}{2s+1}}$ (resp. $n^{-\frac{s}{2s+r}}$ for multi-index models), coinciding with the one-dimensional min-max nonparametric rate, thereby escaping the curse of dimensionality. This rate is achieved with kernel estimators in (Hardle and Stoker, 1989, Theorem 3.3) and (Horowitz, 1998, Section 2.5), and the index $v$ can be estimated at the parametric rate $\mathcal{O}(n^{-\frac{1}{2}})$ under suitable assumptions. The existence of an estimator converging at rate $n^{-\frac{s}{2s+1}}$ is shown in (Györfi et al., 2002, Corollary 22.1), and (Gaïffas and Lecué, 2007, Theorem 2) demonstrates that $n^{-\frac{s}{2s+1}}$ is indeed the min-max rate. These models have also received attention recently both for their connections to “feature learning” in neural networks, see for example Lee et al. (2024) and references therein for learning single-index models with neural networks (trained in a suitable layer-wise fashion), as well as Radhakrishnan et al. (2024), and references therein, showing that deep neural networks appear to exploit low-dimensional linear subspaces for classification and prediction.
Several methods for estimating the index vector $v$ (or the matrix $G$) and $f$ were developed over the years, with varying tradeoffs between the assumptions needed to produce optimal estimators and the computational cost for constructing them. One category includes semiparametric methods based on maximum likelihood estimation (Ichimura, 1993; Hardle et al., 1993; Delecroix et al., 1997; Delecroix and Hristache, 1999; Delecroix et al., 2006; R. J. Carroll and Wand, 1997) and M-estimators that produce $\sqrt{n}$-consistent index estimates under general assumptions, but with computationally demanding implementations, relying on sensitive bandwidth selections for kernel smoothing and on high-dimensional joint optimization of $f$ and $v$. Another category includes direct methods: for example Average Derivative Estimation (Stoker, 1986; Hardle and Stoker, 1989) estimates the index vector $v$ by exploiting its proportionality to $\nabla F$. Early implementations of this idea suffer from the curse of dimensionality due to the use of kernel estimation for the gradient. The technique of (Hristache et al., 2001) uses an iterative scheme to gradually adapt an elongated neighbourhood window. Gradient-based methods (stochastic and non-stochastic) have been studied in the context of neural networks (Lee et al., 2024).
A category of techniques particularly relevant to the present work includes conditional, or inverse regression methods, which derive their estimators from statistics of the conditional distribution of the explanatory variable $X$ given observations of the (noisy) response variable $Y$. Prominent examples include sliced inverse regression (Duan and Li, 1991; Li, 1991), sliced average variance estimation (Cook, 2000), simple contour regression (Li et al., 2005; Coudret et al., 2013), with its analysis (in the multi-index case) in the work (Li and Wang, 2007), which yields an estimator with near-optimal rates for the multi-index model. Conditional methods partition the range of the response variable $Y$ into small intervals and consider their pre-images that are, when $f$ is monotone, slices with the thinnest side along the index direction $v$. They are often straightforward in implementation, consisting of the computation of conditional empirical moments, or other statistics related to the level sets of $f$, and can have only one parameter to tune: the width of slices. Unfortunately, several of the works above, including (Hristache et al., 2001; Li and Wang, 2007), while obtaining asymptotic learning rates that avoid the curse of dimensionality and are in some cases min-max optimal, require a minimal number of samples that is exponential in the ambient dimension $d$ (either explicitly, or through constants exponential in $d$ in the bounds).
Besides the statistical view that the single- and multi-index model do not entail a statistical cost cursed by the ambient dimension, another critical problem is to find algorithms with reasonable computational cost to implement a statistically optimal estimator for these models. For example, even in the case of single-index models, $M$-estimators typically required high-dimensional non-convex optimization in $v$ and $f$; methods based on optimizing over $v$, even when combined with random sampling, may scale exponentially in $d$ due to the need of obtaining points inside a narrow cone around the unknown $v$. Conditional methods can often be implemented in a computationally efficient way, running in time $\mathcal{O}(C_{d}n)$ (up to $\log n$ factors, and with the exception of contour regression methods, which scale quadratically in $n$), with a constant $C_{d}$ a low-order polynomial in $d$. Lanteri et al. (2022) discuss in detail these techniques for the single-index model, and introduce of a variation, called Smallest Vector Regression, which enjoys optimal statistical guarantees up to log factors, provides a theoretically optimal choice for selecting of the slice width parameter (a crucial parameter whose choice is often not discussed (Coudret et al., 2013, p. 75)), and is amenable to a computationally-efficient implementation with cost $\mathcal{O}(d^{2}n\log n)$, all without the curse of dimensionality in the exponents and in the constants, for the rate, the minimum sample requirements and the computational cost.
In conclusion, while the situation is rather well-understood for compositional models $F=f\circ g$ with $g$ linear, much is open when $g$ is nonlinear, in which case it is not clear when the compositional structure allows one to circumvent the curse of dimensionality, except in cases where $g$ has high regularity (global or anisotropic), increasing with $d$, or belong to certain families of polynomials. We propose a simple model that, perhaps surprisingly, without such strong regularity assumptions, still is amenable to nonparametric estimators that avoid the curse of dimensionality, both statistically and computationally.
1.1 The Nonlinear Single-Variable Model
Our contributions in this work are the following:
  1. (i)
    we introduce a model, called the Nonlinear Single-Variable Model, that is intermediate between the semi-parametric single-index model and the nonparametric function composition model: both the outer and inner functions $f$ and $g$ are nonlinear, but both the distribution $\rho_{X}$ and the inner function $g$ have a special structure related to the geometry of the problem, with the range of $g$ being an unknown curve $\gamma$ in $\mathbb{R}^{d}$;
  2. (ii)
    we construct an efficient estimator that provably defeats the curse of dimensionality by achieving a dimension-independent optimal (up to log factors) learning rate, with no constants nor minimum sample requirements cursed by the ambient dimension $d$;
  3. (iii)
    an efficient, near linear time algorithm that constructs the estimator given data. All the constants, in both the learning bounds and in the computational costs, scale as low-order polynomials in the dimension $d$, making the estimator practical.
This model generalizes the single-index model by allowing for $g$ nonlinear (but with a certain geometric structure), while still being amenable to estimation with strong statistical and computational guarantees that are not cursed by the ambient dimension, and with no regularity assumptions that scale with the dimension. To our knowledge, this may be the first (nontrivial) statistical model that is shown to possess all these properties.
Definition 1 (Nonlinear Single-Variable Model)
The regression function $F$ has the following decomposition:
$\mathbb{E}[Y|X]=F(X)=f(\Pi_{\gamma}X)$
where the underlying curve $\gamma:[0,\mathrm{len}_{\gamma}]\to\mathbb{R}^{d}$ is parametrized by arc-length and has length $\mathrm{len}_{\gamma}$ , and the link function $f\,:\,[0,\mathrm{len}_{\gamma}]\to\mathbb{R}^{1}$ depends only on one variable. The random vector $X$ is supported in some domain $\Omega_{\gamma}\subseteq\mathbb{R}^{d}$ containing the image of $\gamma$, such that the closest-point projection $\Pi_{\gamma}:\Omega_{\gamma}\to[0,\mathrm{len}_{\gamma}]$, with
$\Pi_{\gamma}(x):=\operatorname*{{arg\,min}}_{t\in[0,\mathrm{len}_{\gamma}]}\left\|x-\gamma(t)\right\|\,,$
is well-defined on $\Omega_{\gamma}$, i.e. the minimizer is unique.
We express the random vector $X$ as its position along the curve and its displacement away from the curve:
$X=\gamma(t)+M_{\gamma^{\prime}(t)}\begin{pmatrix}Z_{d-1}\\ 0\end{pmatrix}$ (1)
with $t=\Pi_{\gamma}X$, $t$ a random variable on $[0,\mathrm{len}_{\gamma}]$ and $Z_{d-1}$ a centered random vector in $\mathbb{R}^{d-1}$. For each unit vector $v\in\mathbb{S}^{d-1}$, $M_{v}\in\mathbb{O}(d)$ is an orthogonal matrix that maps the $d$-th canonical basis vector $\hat{e}_{d}$ to the vector $v$. Therefore, conditioned on $t=t_{0}\in[0,\mathrm{len}_{\gamma}]$, the random vector $M_{\gamma^{\prime}(t_{0})}\begin{pmatrix}Z_{d-1}\\ 0\in\end{pmatrix}\in\operatorname{span}\{\gamma^{\prime}(t_{0})\}^{\perp}$ is the displacement of $X$ away from $\gamma$.
The regression problem for the Nonlinear Single-Variable Model: Given pairs $(X_{i},Y_{i})_{i=1}^{n}$ with $X_{i}$’s independent copies of $X$ as above, and $Y_{i}=F(X_{i})+\zeta_{i}$, with $F$ as in Definition 1 and observational noise $\zeta_{i}$, we goal is to estimate the regression function $F$ on the support of $X$. The $\zeta_{i}$’s are assumed sub-Gaussian, independent among themselves and from the $X_{i}$’s. Note that the distribution of $X$ is unknown, as are both the link function $f$ and the underlying curve $\gamma$. We illustrate one example in Fig.1.
Figure 1: One example of a Nonlinear Single-Variable Model (1): the underlying curve $\gamma$, plotted in black, is a Meyer helix in $\mathbb{R}^{36}$ (details in Appendix C) with $\sigma_{\gamma}=0.5$ and the link function $f\in\mathcal{C}^{0.7}(\mathbb{R}^{1})$ is strictly monotone, and $\zeta\equiv 0$. We generate $n=5000$ samples $X_{i}$ scattered the curve in a tube of radius 6, colored by $Y_{i}=F(X_{i})=f(\Pi_{\gamma}X_{i})$. Left: Random projection of the data onto $\mathbb{R}^{3}$. Right: Orthogonal projection of data onto the first $3$ principal components. The distribution around this curve does not appear to be linearly embeddable in low dimensions without increasing its complexity, see Appendix C.
The domain $\Omega_{\gamma}\subseteq\mathbb{R}^{d}$ needs to be such that the closest-point projection $\Pi_{\gamma}$ is well-defined on it: this condition is connected with the concept of reach of $\gamma$ (Federer, 1959). Starting with the distance function ${\mathrm{dist}}_{\gamma}(x):\mathbb{R}^{d}\to[0,\infty)$, ${\mathrm{dist}}_{\gamma}(x)=\inf\{\left\|x-\gamma(t_{0})\right\|:\ t_{0}\in[0,\mathrm{len}_{\gamma}]\}$, the domain $Unp_{\gamma}$ is defined as the set of all points $x\in\mathbb{R}^{d}$ for which there is a unique point $\gamma(t_{0})$ closest to $x$. The map $\Pi_{\gamma}:Unp_{\gamma}\to[0,\mathrm{len}_{\gamma}]$ maps $x\in Unp_{\gamma}$ to the unique $t\in[0,\mathrm{len}_{\gamma}]$ such that ${\mathrm{dist}}_{\gamma}(x)=\left\|x-\gamma(t)\right\|$. For $t_{0}\in[0,\mathrm{len}_{\gamma}]$, the local reach is $\mathrm{reach}_{\gamma}(t_{0}):=\sup\{r:\ B(\gamma(t_{0}),r)\subseteq Unp_{\gamma}\}$, and the global reach is $\mathrm{reach}_{\gamma}:=\inf\{\mathrm{reach}_{\gamma}(t_{0}):\ t_{0}\in[0,\mathrm{len}_{\gamma}]\}$. Both $\mathrm{reach}_{\gamma}(t_{0})$ and $\mathrm{reach}_{\gamma}$ take values in $[0,\infty]$. With these definitions, $\Omega_{\gamma}$ could be as large as $\bigcup_{t_{0}\in[0,\mathrm{len}_{\gamma}]}B(\gamma(t_{0}),\mathrm{reach}_{\gamma}(t_{0}))$. Note that, therefore, the distribution $\rho_{X}$ is neither supported on $\gamma$, nor highly concentrated on $\gamma$, unlike the aforementioned models for regression on manifolds, and similarly to single-index models, where the distribution of the data typically has large variance in the directions normal to the direction of the index vector: here we are letting this variance to be as large as possible, constrained on keeping $\Pi_{\gamma}$, the natural generalization of the orthogonal projection onto the line spanned by the single index, well-defined.
For any positive integer $m\in\mathbb{N}$ and any positive real number $s\in\mathbb{R}$, we define the function space $\mathcal{C}^{s}(\mathbb{R}^{m})$ via the following semi-norm: A function $f:\mathbb{R}^{m}\to\mathbb{R}$ is called $s$-smooth, for $s>0$, with $s=s_{1}+s_{2}$ for nonnegative $s_{1}\in\mathbb{N}_{0}$, chosen to be the largest integer strictly less than $s$, and positive $s_{2}\in(0,1]$, if there exists a positive value $\left[f\right]_{\mathcal{C}^{s}}>0$ such that for every nonnegative $m$-tuple $(\alpha_{1},\dots,\alpha_{m})\in\mathbb{N}_{0}^{m}$ with $\sum_{i}\alpha_{i}=s_{1}$, we have $\left|\frac{\partial^{s_{1}}f(x_{1})}{\partial x_{1}^{\alpha_{1}}\cdots x_{m}^{\alpha_{m}}}-\frac{\partial^{s_{1}}f(x_{2})}{\partial x_{1}^{\alpha_{1}}\cdots x_{m}^{\alpha_{m}}}\right|\leqslant\left[f\right]_{\mathcal{C}^{s}}\left\|x_{1}-x_{2}\right\|^{s_{2}}$ for any $x_{1},x_{2}\in\mathbb{R}^{m}$, i.e. all the $s_{1}$-th derivatives of $f$ are Hölder continuous with exponent $s_{2}$. The smallest constant $\left[f\right]_{\mathcal{C}^{s}}$ that can be chosen in the above inequalities is called the $s$-smooth semi-norm of $f$. The function space $\mathcal{C}^{s}(\mathbb{R}^{m})$ consists of functions $f\in L^{\infty}(\mathbb{R}^{m})$ with finite semi-norm $\left[f\right]_{\mathcal{C}^{s}}$.
We ready for an informal version of the main Theorems 3 and 4:
Theorem 2 (Informal)
Suppose that $f\in\mathcal{C}^{s}(\mathbb{R}^{1})$ for some $s\in[\frac{1}{2},2]$ and that $f$ is coarsely monotone. With some assumptions on the underlying curve $\gamma$, the distribution $\rho_{X}$ of the random variable $X$, and the variance $\sigma_{\zeta}^{2}$ of the noise $\zeta$, if the number of training data $n$ satisfies $n\geqslant\operatorname{poly}(d,\mathrm{len}_{\gamma})$, then the estimator ${\widehat{F}}$ constructed by Algorithm 1 satisfies
$\mathbb{E}\left[\left|{\widehat{F}}(X)-F(X)\right|^{2}\right]\lesssim C_{1}(f,\gamma,\rho_{X},\sigma_{\zeta},d)n^{-\frac{2s}{2s+1}}\log n+C_{2}(f,\gamma,\rho_{X},\sigma_{\zeta},d)\max(\sigma_{\zeta},\omega_{f})^{2(s\wedge 1)}\,.$
If we further suppose that there is no observational noise, i.e., $\zeta=0$ almost surely, and that $f$ is strictly monotone, then the estimator ${\widehat{F}}$ satisfies
$\mathbb{E}\left[\left|{\widehat{F}}(X)-F(X)\right|^{2}\right]\lesssim C_{3}(f,\gamma,\rho_{X},d)\left(\frac{\log^{3}n}{n^{2}}\right)^{s\wedge 1}\,.$
The dependency of the constants $C_{1},C_{2},C_{3}$ on $d$ is a low-order polynomial, and ${\widehat{F}}$ can be constructed by an algorithm that runs in time $\mathcal{O}(d^{2}n\log n)$.
The bound on the expected $L^{2}(\rho_{X})$ error of our estimator contains two terms: the first one shows that the learning rate is near optimal (up to $\log n$ factors) as it is the min-max rate of one-dimensional non-parametric regression for $f$, as if we knew the curve $\gamma$; the second term is a constant at which our estimator saturates, due to the estimator producing a piecewise linear approximation to $\gamma$, but only at scales above $\sigma_{\zeta}$. The estimator in fact produces an estimate of the underlying curve $\gamma$, of the (nonlinear) closest-point projection $\Pi_{\gamma}$, and of the link function $f$, thereby providing an interpretable result in terms of all the terms in the compositional structure to the regression function $F$. Further remarks may be found after the formal statement of the main Theorems, in Section 2.
We conclude this section by reporting in Table 1 several math symbols used throughout this paper.
Table 1: Notation
symbol definition symbol definition
$C,c$ positive absolute constants $\|A\|$ spectral norm of a matrix $A$
$a\lesssim b$ $a\leq Cb$ for some positive absolute constant $C$ $a\asymp b$ $a\lesssim b$ and $b\lesssim a$
$a\wedge b$ minimum of $\{a,b\}$ $|I|$ Lebesgue measure of an interval $I$
$\lambda_{m}(A)$ $m$-th largest eigenvalues of a square matrix $A$ $v_{m}(A)$ singular vector corresponding to $\lambda_{m}(A)$
$B(x,r)$ Euclidean ball of center $x$ and radius $r$ $\mathbbm{1}(E)$ indicator function of an event $E$
$\operatorname{span}\{\mathcal{S}\}$ linear span of a set $\mathcal{S}$ $P_{u}$ orthogonal projection onto $\operatorname{span}\{u\}$
$\{\mathcal{S}\}^{\perp}$ orthogonal complement of a set $\mathcal{S}$ $\Pi_{\gamma}$ nearest point projection onto the position along $\gamma$
$\mathrm{len}_{\gamma}$ length of the curve $\gamma$ $\mathrm{reach}_{\gamma}$ reach of the curve $\gamma$
$F=f\circ\Pi_{\gamma}$ unknown regression function from $\mathbb{R}^{d}$ to $\mathbb{R}^{1}$ $f$ unknown link function from $\mathbb{R}^{1}$ to $\mathbb{R}^{1}$
$\left[f\right]_{\mathcal{C}^{s}}$ semi-norm of an $s$-smooth function $\left|f\right|_{L^{\infty}}$ $L^{\infty}$-norm for function $f$
$X$ random vector in $\mathbb{R}^{d}$ with density function $\rho_{X}$$\{X_{i}\}_{i=1}^{n}$ samples of $X$
$Y$ random variable dependent on $X$$\{Y_{i}\}_{i=1}^{n}$ samples of $Y$
$\zeta$ random variable modeling noise $\zeta_{i}$ samples of $\zeta$
$R$ bounded interval, range of $\{Y_{i}\}_{i=1}^{n}$ $l$ total number of partitions of interval $R$
$\{R_{l,h}\}_{h=1}^{l}$ partition intervals of $R$ indexed by $h$ $\widehat{R}_{l,h}$ the set of sample $Y_{i}$ such that $Y_{i}\in{R_{l,h}}$
${S_{l,h}}$ slice, conditional distribution $X|Y\in{R_{l,h}}$ $\widehat{S}_{l,h}$ empirical slice, the set of sample $X_{i}$ such that $Y_{i}\in{R_{l,h}}$
$n$ number of sample points $\{(X_{i},Y_{i})\}_{i=1}^{n}$ ${n_{l,h}}$ number of samples in empirical slice $\widehat{S}_{l,h}$
${\mu_{l,h}}$ center of slice ${S_{l,h}}$, i.e. $\mathbb{E}[X|Y\in{R_{l,h}}]$ ${\widehat{\mu}_{l,h}}$ empirical mean of points in $\widehat{S}_{l,h}$
${v_{l,h}}$ significant vector of slice ${S_{l,h}}$ ${\widehat{v}_{l,h}}$ empirical significant vector of points in $\widehat{S}_{l,h}$
$\Sigma_{l,h}$ covariance matrix for slice ${S_{l,h}}$ ${\widehat{\Sigma}_{l,h}}$ empirical covariance matrix for points in $\widehat{S}_{l,h}$
$H_{l,h}$ geometric quantity of slice ${S_{l,h}}$ $\widehat{H}_{l,h}$ estimated geometric quantity for points in $\widehat{S}_{l,h}$
${\mathrm{dist}}(x,h)$ distance from point $x$ to slice ${S_{l,h}}$ ${\widehat{\mathrm{dist}}}(x,h)$ estimated distance from point $x$ to empirical slice $\widehat{S}_{l,h}$
${h_{x}}$ nearest index for point $x$ $I$ bounded subinterval of one-dimensional projection of $X$
$j$ number of partition of subinterval $I$ ${\widehat{f}_{j|v}}$ estimator of $f$ by local polynomial fitting
2 An estimator for the Nonlinear Single-Variable Model
We propose an estimator for $F$, and a corresponding efficient algorithm, for the Nonlinear Single-Variable Model based on inverse (or conditional) regression, which also produces a sketch of the curve $\gamma$ and an estimator of the closest-point projection $\Pi_{\gamma}$.
Step 1: extract geometric features of the underlying curve $\gamma$. Given data $\{(X_{i},Y_{i})\}_{i=1}^{n}$, let $R$ be an interval containing all the $Y_{i}$’s, and $\{{R_{l,h}}\}_{h=1}^{l}$ a partition of $R$, either uniform or based on empirical quantiles (so that all ${R_{l,h}}$ contain the same number of points). We partition the data $\{(X_{i},Y_{i})\}_{i=1}^{n}$ into pairs $\{(\widehat{S}_{l,h},\widehat{R}_{l,h})\}_{h=1}^{l}$ where $\widehat{R}_{l,h}:=\{Y_{i}:Y_{i}\in{R_{l,h}}\}$ and $\widehat{S}_{l,h}:=\{X_{i}:Y_{i}\in{R_{l,h}}\}$. Each empirical slice $\widehat{S}_{l,h}$ is the empirical pre-image of the interval $R_{l,h}$ in the output variable $Y$. We perform the Principal Component Analysis (PCA) of each $\widehat{S}_{l,h}$ to obtain its mean ${\widehat{\mu}_{l,h}}$ and its “significant vector” ${\widehat{v}_{l,h}}$: ${\widehat{\mu}_{l,h}}$ will be approximately on $\gamma$, and ${\widehat{v}_{l,h}}$ will be approximately tangent to $\gamma$ at ${\widehat{\mu}_{l,h}}$, yielding a local first-order approximation to $\gamma$.
Step 2: design a distance function ${\mathrm{dist}}(x,h)$ (and its empirical version ${\widehat{\mathrm{dist}}}(x,h)$), based on the geometric shape of ${S_{l,h}}$ (and $\widehat{S}_{l,h}$ respectively), that measure how far the point $x$ is away from the slice ${S_{l,h}}$ ($\widehat{S}_{l,h}$, respectively). By assigning each $x\in\Omega_{\gamma}$ to the “nearest” slice according to this distance function, we partition the domain $\Omega_{\gamma}$ into several local neighborhoods and we use the significant vector ${\widehat{v}_{l,h}}$ to project the points $X_{i}\in\widehat{S}_{l,h}$ onto a one-dimensional interval $I^{(l,h)}$ in each local neighborhood. Because the significant vector ${\widehat{v}_{l,h}}$ is approximately tangential to the curve $\gamma$, this linear projection approximates the nonlinear projection $\Pi_{\gamma}$ in each local neighborhood.
Step 3: perform a one-dimensional piecewise polynomial regression on the projected points in $I^{(l,h)}$ and obtain a local estimator of the regression function $F$ in a local neighborhood. This, together with the “nearest” slice assignment in the second step, allows us to construct a global estimator of $F$ in the whole domain $\Omega_{\gamma}$.
2.1 Extracting the geometric features of the underlying curve $\gamma$
For each $t\in[0,\mathrm{len}_{\gamma}]$, let $\hat{n}(t):=\frac{\gamma^{\prime\prime}(t)}{\left\|\gamma^{\prime\prime}(t)\right\|}$ be the unit normal vector to $\gamma$, pointing inwards the circle of curvature. For $x\in\Omega_{\gamma}$ we define the signed projected distance from $x$ to $\gamma$ to be $\tilde{d}(x,\gamma):=\langle x-\gamma(\Pi_{\gamma}x),\hat{n}(\Pi_{\gamma}x)\rangle_{\mathbb{R}^{d}}$, which is zero if $x$ is on the curve $\gamma$, positive if $x$ is inside the circle of curvature, and negative otherwise. A direct calculation shows that the compositional structure $F=f\circ\Pi_{\gamma}$ in Definition 1 implies that for $x\in\Omega_{\gamma}$
${\nabla F}(x)=\frac{f^{\prime}(\Pi_{\gamma}x)}{1-\left\|\gamma^{\prime\prime}(\Pi_{\gamma}x)\right\|\tilde{d}(x,\gamma)}{\gamma^{\prime}}(\Pi_{\gamma}x)\,.$ (2)
Therefore, the gradient vectors of points in each level set $\Pi_{\gamma}^{-1}(t_{0})=\{x:\Pi_{\gamma}x=t_{0}\}$ are all parallel to ${\gamma^{\prime}}(t_{0})$, albeit with magnitude depending on the relative position of $x$ and $\gamma$. As a consequence, each $\Pi_{\gamma}^{-1}(t_{0})$ is contained in a hyperplane. If we could perform singular value decomposition for points on each level set $\Pi_{\gamma}^{-1}(t_{0})$, the singular vector corresponding to the 00 singular value would be parallel to ${\gamma^{\prime}}(t_{0})$.
The geometry of the level sets will play a prominent role in constructing our estimator of the Nonlinear Single-Variable Model, as it did in some of the earlier works on the single-index model including (Stoker, 1986; Hardle and Stoker, 1989), and recent refinements such as the Smallest Vector Regression estimator of Lanteri et al. (2022). In the single-index model, where the underlying curve $\gamma$ is a line segment with direction $v$, any level set $\{x:F(x)=c\}$ is a hyperplane perpendicular to $v$. As in (Lanteri et al., 2022), we shall take a uniform partition on the empirical range $R:=[\min(Y_{i}),\max(Y_{i})]$, consisting of suitably small intervals $\{{R_{l,h}}\}_{h=1}^{l}$ where $l\in\mathbb{N}$ is the total number of partitioning intervals, indexed by $h=1,\dots,l$; for each partitioning interval ${R_{l,h}}$, we consider the empirical slice
$\widehat{S}_{l,h}:=\{X_{i}:Y_{i}\in{R_{l,h}}\}\,,$
which is the pre-image of ${R_{l,h}}$ at the level of samples, i.e., a sample from the conditional distribution $X|Y\in{R_{l,h}}$. Each empirical slice $\widehat{S}_{l,h}$ is utilized in Lanteri et al. (2022) as an approximation to a level set $\Pi_{\gamma}^{-1}(s_{0})$, and $\widehat{S}_{l,h}$ should be “thin” along the direction of $v$ (and therefore $\nabla F$) and “wide” on directions orthogonal to $v$, at least under suitable assumptions on $F$ and noise level $\sigma_{\zeta}$. Lanteri et al. (2022) then perform local PCA on each $\widehat{S}_{l,h}$ and use the smallest principal component to approximate the direction of $v$: since $\gamma$ is a line with direction $v$, these smallest principal components should all be independent estimators of the index vector $v$, and these per-slice estimates can be suitably averaged across all slices to obtain an estimator for $v$. Once $v$ is estimated, the input data is projected onto this line, and one-dimensional non-parametric regression yields an estimator for $f$.
In the Nonlinear Single-Variable Model, we are still going to use empirical slices, but here we cannot perform an aggregation of estimated vectors because the tangent vectors to $\gamma$ are not constant due to the curvature of $\gamma$. Instead, we can only rely on local information to estimate the nearest-point projection onto $\gamma$, which is now a nonlinear function. Each $\widehat{S}_{l,h}$ still approximates some level set $\Pi_{\gamma}^{-1}(t_{0})$ under the Assumption (LCV) below: roughly speaking, when ${R_{l,h}}$ has a sufficiently small diameter, the $\widehat{S}_{l,h}$ is “thin” along the tangential direction $\gamma^{\prime}(t_{0})$ and “wide” along directions orthogonal to $\gamma^{\prime}(t_{0})$. Consequently, PCA on each empirical slice $\widehat{S}_{l,h}$ locally will yield:
  1. (i)
    the empirical mean ${\widehat{\mu}_{l,h}}$, which approximates the conditional expectation ${\mu_{l,h}}:=\mathbb{E}[X|Y\in{R_{l,h}}]$ and be approximately on the curve $\gamma$;
  2. (ii)
    the empirical covariance matrix ${\widehat{\Sigma}_{l,h}}$, which should approximate the conditional covariance matrix $\Sigma_{l,h}:=\mathbb{E}[(X-{\mu_{l,h}})(X-{\mu_{l,h}})^{\intercal}|Y\in{R_{l,h}}]$;
  3. (iii)
    the smallest principal component ${\widehat{v}_{l,h}}$ of ${\widehat{\Sigma}_{l,h}}$, which should approximate the smallest principal component ${v_{l,h}}$ of $\Sigma_{l,h}$ and be approximately tangential to the curve.
These yield a first-order approximation of underlying curve $\gamma$ in a local neighborhood of $\widehat{S}_{l,h}$. Figure 2 depicts an example where the sample mean is approximately on the curve, and the smallest principal component vector is approximately tangential to the curve.
[…]
Figure 2: In the same setup as in Fig.1, we partition the range uniformly into $l=800$ intervals, and consider two slices. Top: a visualization of the two empirical slices, where we only plot 2000 samples per slice (in green and blue), with $\gamma$ in black. The red circles and vectors are the sample means and smallest principal components of the two empirical slices. Bottom: bar plots with the largest, second largest, average, second smallest, and smallest singular value of the empirical slices. The smallest singular value is significantly smaller than the others.
However, the above argument relies on the Assumption (LCV), which imposes some restrictions on the underlying curve $\gamma$ and the “radius” of $\Omega_{\gamma}$ around $\gamma$, the regression function $F$, and the noise level $\sigma_{\zeta}$. Without such an assumption, it might not be the case that slices are “thin” in the direction tangent to the curve for various reasons: (i) with an insufficient amount of samples, choosing very fine partitions of the range $R$ is not optimal: smaller intervals will contain fewer samples, leading to a high variance in the estimation of the mean and principal component(s) of the slices; (ii) the observational noise $\zeta$ in the outputs forces a lower bound on the diameter of the partitioning intervals ${R_{l,h}}$ of the range $R$: subdividing the range into intervals with size smaller than the noise level $\sigma_{\zeta}$ does not improve estimation, nor would make the slices thinner; (iii) the reach $\mathrm{reach}_{\gamma}$ may be small because of the complexity of the underlying curve $\gamma$; (iv) the distribution of $X$ might strongly concentrate around $\gamma$, i.e. $\sigma_{\gamma}$, the standard deviation of $Z_{d-1}$, is small. In these situations, we might encounter slices that are “wide”, i.e., much wider in the curve’s local tangent direction than in all other directions—essentially the opposite of being “thin”. In this case, the largest principal component can be significantly larger than the remaining components and is the one that aligns with the tangent direction of $\gamma$. Figure 3 illustrates how, in these situations, the sample mean and the largest principal component of a slice may provide an approximation of the underlying curve $\gamma$. In this setting, the largest singular value is significantly larger than the other ones.
[…]
Figure 3: In the same setup and visual conventions of Fig.1, but with the range $R$ split uniformly into 80 intervals, and a different pair of slices. Top: the slices are now elongated along the curve, rather than perpendicularly to it. Bottom: the largest singular value is now significantly larger than the remaining ones.
We want our algorithm to adapt to both “thin” and “wide” slice scenarios discussed above: consider a slice ${S_{l,h}}$ with conditional mean ${\mu_{l,h}}$ and conditional covariance matrix $\Sigma_{l,h}$, and their empirical counterparts $\widehat{S}_{l,h}$, ${\widehat{\mu}_{l,h}}$, ${\widehat{\Sigma}_{l,h}}$. We define $H_{l,h}$ and $\widehat{H}_{l,h}$ as
$H_{l,h}:=\log\left(\frac{{\lambda_{\mathrm{mid}}(\Sigma_{l,h})}^{2}}{{\lambda_{1}(\Sigma_{l,h})}{\lambda_{d}(\Sigma_{l,h})}}\right),\ \ \widehat{H}_{l,h}:=\log\left(\frac{{\lambda_{\mathrm{mid}}(\widehat{\Sigma}_{l,h})}^{2}}{{\lambda_{1}(\widehat{\Sigma}_{l,h})}{\lambda_{d}(\widehat{\Sigma}_{l,h})}}\right)\,,$
where for any $d\times d$ square matrix $M$, $\lambda_{1}(M)\geqslant\lambda_{2}(M)\geqslant\cdots\geqslant\lambda_{d-1}(M)\geqslant\lambda_{d}(M)$ are the eigenvalues of $M$ in descending order, and $\lambda_{\mathrm{mid}}(M)=(\lambda_{2}(M)\times\cdots\times\lambda_{d-1}(M))^{{1}/(d-2)}$ is the geometric mean of the eigenvalues excluding the largest and the smallest ones. A slice ${S_{l,h}}$ is “thin” when ${\lambda_{d}(\Sigma_{l,h})}\ll{\lambda_{1}(\Sigma_{l,h})}\approx{\lambda_{2}(\Sigma_{l,h})}\approx\cdots\approx{\lambda_{d-1}(\Sigma_{l,h})}$, and hence $H_{l,h}>0$; the larger $H_{l,h}$ is, the “thinner” ${S_{l,h}}$ is. A slice ${S_{l,h}}$ is “wide” when ${\lambda_{1}(\Sigma_{l,h})}\gg{\lambda_{2}(\Sigma_{l,h})}\approx\cdots\approx{\lambda_{d}(\Sigma_{l,h})}$, and hence $H_{l,h}<0$; the more negative $H_{l,h}$ is, the “wider” ${S_{l,h}}$ is. If $H_{l,h}$ is close to zero, then the geometric shape of the slice ${S_{l,h}}$ is undetermined: it may be roughly isotropic or may have both very large ${\lambda_{1}(\Sigma_{l,h})}$ and very small ${\lambda_{d}(\Sigma_{l,h})}$. We define the significant vector ${v_{l,h}}$ and the empirical significant vector ${\widehat{v}_{l,h}}$ as
${v_{l,h}}:=\begin{cases}v_{d}(\Sigma_{l,h})&\text{ if }H_{l,h}>0\\ v_{1}(\Sigma_{l,h})&\text{ if }H_{l,h}<0\end{cases},\ \ \ \ {\widehat{v}_{l,h}}:=\begin{cases}v_{d}({\widehat{\Sigma}_{l,h}})&\text{ if }\widehat{H}_{l,h}>0\\ v_{1}({\widehat{\Sigma}_{l,h}})&\text{ if }\widehat{H}_{l,h}<0\end{cases}\,.$
The significant vector ${v_{l,h}}$ is used to estimate the tangent vector $\gamma^{\prime}$ in both the “thin” and “wide” slice scenarios. When $H_{l,h}\approx 0$, we only expect to use the sample mean ${\mu_{l,h}}$ as a local 00-th order approximation to $\gamma$, as the slice has no preferred direction. Crucially, while in the “thin” scenario we expect ${\widehat{v}_{l,h}}$ to be a good approximation of the tangential vector, in the “wide” scenario the curvature of $\gamma$ can have a significant effect in our estimation of the local direction of the curve.
2.2 Estimating the nonlinear projection $\Pi_{\gamma}$ by assigning points to the “nearest” slice
Before regressing $f$, we construct an estimator for $\Pi_{\gamma}(x)$, which is the value of the parameter in $[0,\mathrm{len}_{\gamma}]$ of the curve such that $\gamma(\Pi_{\gamma}(x))$ is the closest point on the curve to a point $x\in\Omega_{\gamma}$. We design a distance function ${\mathrm{dist}}(x,h)$ between $x$ and slice ${S_{l,h}}$ and assign $x$ to the “nearest” slice under this distance function. This assignment maps $\Omega_{\gamma}$ onto $\{1,\dots,l\}$ and thus can be interpreted as a zeroth-order approximation of $\Pi_{\gamma}$. After this assignment, we use ${\mu_{l,h}}$ and ${v_{l,h}}$ to obtain a first-order approximation of $\Pi_{\gamma}(\cdot)$ locally on each slice ${S_{l,h}}$.
Our choice of distance function ${\mathrm{dist}}(x,h)$ is dictated by two purposes. First, it should be “local”, i.e., the distance between $x$ and the center of a slice should play a role. Second, it should be anisotropic: on any level set $\{x:\Pi_{\gamma}x=t_{i}\}$ we have $\langle x-\gamma(t_{i}),\gamma^{\prime}(t_{i})\rangle=0$, so the distance to the hyperplane normal to $\gamma^{\prime}(t_{i})$ should play a prominent role, but cannot be too dominant, as there may be multiple slices, even far away from each other, with $\langle x-\gamma(t_{i}),\gamma^{\prime}(t_{i})\rangle\approx 0$. We therefore consider a distance function of the form
$\left|\langle x-{\mu_{l,h}},{v_{l,h}}\rangle\right|+c\left\|x-{\mu_{l,h}}\right\|$
for some $c>0$. The value of $c$ cannot be too small, or this distance function would fail for highly self-entangled curves, nor too large because we want a small distance ${\mathrm{dist}}(x,h)$ if the point $x$ is close to the slice ${S_{l,h}}$. The optimal choice of $c$ depends on the curve and the distribution of data around it. We define the distance function ${\mathrm{dist}}(x,h)$ separately in the “thin” slice scenario and the “wide” slice scenario:
$\begin{split}{\mathrm{dist}}(x,h)&=\begin{cases}|\langle x-{\mu_{l,h}},{v_{l,h}}\rangle|^{2}+\frac{{\lambda_{d}(\Sigma_{l,h})}}{{\lambda_{1}(\Sigma_{l,h}^{b})}}\left\|x-{\mu_{l,h}}\right\|^{2}\ \text{ if }H_{l,h}>0\\ \left\|x-{\mu_{l,h}}\right\|^{2}+\frac{{\lambda_{d}(\Sigma_{l,h})}}{{\lambda_{1}(\Sigma_{l,h})}}|\langle x-{\mu_{l,h}^{b}},{v_{l,h}}\rangle|^{2}\ \ \text{ if }H_{l,h}<0\end{cases}\,,\end{split}$ (3)
and similarly for its empirical counterpart ${\widehat{\mathrm{dist}}}(x,h)$. In the “thin” slice scenario, this distance function focuses on measuring the displacement between $x$ and ${S_{l,h}}$ along the direction ${v_{l,h}}$, and is less sensitive to the displacement orthogonal to ${v_{l,h}}$. In the “wide” scenario, ${\mathrm{dist}}(x,h)$ instead pays less attention to the displacement along ${v_{l,h}}$ and focuses on the displacement orthogonal to ${v_{l,h}}$. Note that when the shape of the slice ${S_{l,h}}$ (and $\widehat{S}_{l,h}$) is roughly isotropic, $\lambda_{d}/\lambda_{1}$ is roughly one, so the two cases in (3) are consistent with each other and the distance above varies regularly as $H_{l,h}$ changes sign. This distance function ${\mathrm{dist}}(x,h)$ has the following advantages: (i) it focuses on local slices while incorporating information about the geometry of each slice; (ii) inside the local neighborhood, it pays special attention to displacement along the tangential direction; (iii) it is distribution-adaptive, allowing, for example, to handle in robust fashion heteroscedasticity in the distribution of $X$ around the curve; (iv) its performance is amenable to mathematical analysis.
We will use equation (3) in the proof of convergence of our algorithm. We remark that this distance is a simplification of the Mahalanobis distance from $x$ to a slice ${S_{l,h}}$ ${\mathrm{dist}}(x,h):=||\Sigma_{l,h}^{-{1/2}}(x-{\mu_{l,h}})||$, which also puts heavier weight on the displacement along the tangential direction in the “thin” slice scenario and lighter weight in the “wide” slice scenario. As it is a bit harder to analyze mathematically than the distance in (3), we use it only in numerical simulations.
Slices with little data yield estimators with large variance, and will be disregarded: let ${n_{l,h}}:=\#{S_{l,h}}$ be the number of samples in ${S_{l,h}}$, and define the index set of “heavy” slices
$\mathcal{H}_{l}:=\{h\in\{1,\dots,l\}:{n_{l,h}}\geqslant n/l\}\,.$
For $x\in\Omega_{\gamma}$, we define the “nearest” index for the slice that $x$ belongs to as
${h_{x}}:=\operatorname*{{arg\,min}}_{h\in\mathcal{H}_{l}}{\mathrm{dist}}(x,h)\,,$
and the “correct” index for the slice that $x$ belongs to as the unique index ${h^{\prime}_{x}}$ such that $F(x)\in R_{l,{h^{\prime}_{x}}}$. For almost all $x\in\Omega_{\gamma}$, the minimizer ${h_{x}}$ is unique, since the set of points that cannot be uniquely assigned is a subset of $\{x\in\Omega_{\gamma}\subseteq\mathbb{R}^{d}:\exists h^{\prime},h^{\prime\prime}\text{ s.t. }{\mathrm{dist}}(x,h^{\prime})={\mathrm{dist}}(x,h^{\prime\prime})\}$, which is a finite union of hyper-surfaces in $\mathbb{R}^{d}$ and thus has measure zero. We will show that in the “thin” slice scenario, the nearest index ${h_{x}}$ is almost the correct index ${h^{\prime}_{x}}$ under suitable assumptions, i.e., $|{h^{\prime}_{x}}-{h_{x}}|\leqslant 1$ (Section 3). The possibility of an error $|{h_{x}}-{h^{\prime}_{x}}|=1$ stems from the possibility that points near the boundary of a slice can be misclassified to one of its adjacent slices. Given sample data, the “nearest” slice is estimated using the empirical counterpart of the distance, yielding
${\widehat{h}_{x}}:=\operatorname*{{arg\,min}}_{h\in\mathcal{H}_{h}}{\widehat{\mathrm{dist}}}(x,h)\,.$
We shall prove that $|{\widehat{h}_{x}}-{h_{x}}|\leqslant 1$ with high probability, with the possibility of $|{h^{\prime}_{x}}-{h_{x}}|=1$ and $|{\widehat{h}_{x}}-{h_{x}}|=1$ both stemming from points near the boundary of a slice being misclassified to one of its adjacent slices. Nonetheless, the probability of adjacent misclassification (i.e., ${\widehat{h}_{x}}={h^{\prime}_{x}}\pm 1$) is relatively small but does not decrease to zero as the sample size increases: to avoid artifacts in the regression stage, we include data from adjacent slices to estimate the regression function in each local neighborhood.
2.3 Local estimator of the link function $f$ and global estimator of the regression function $F$
After assigning a point $x\in\mathbb{R}^{d}$ to an estimated slice $\widehat{S}_{l,{\widehat{h}_{x}}}$ with index ${\widehat{h}_{x}}$, we use piecewise polynomial estimators to regress the link function $f$ on the corresponding slice. As noted above, we also need to consider empirical slices adjacent to $\widehat{S}_{l,{\widehat{h}_{x}}}$. First, we project the data from $\bigcup_{|h-{\widehat{h}_{x}}|\leqslant 1}\widehat{S}_{l,h}$ orthogonally onto the line with direction ${\widehat{v}_{l,\widehat{h}_{x}}}$. We consider the projected data $\{\langle{\widehat{v}_{l,\widehat{h}_{x}}},X_{i}\rangle:X_{i}\in\bigcup_{|h-{\widehat{h}_{x}}|\leqslant 1}\widehat{S}_{l,h}\}$, cover it with an interval $I^{(l,{\widehat{h}_{x}})}$, and further partition $I^{(l,{\widehat{h}_{x}})}$ uniformly into $j$ smaller intervals $\{I^{(l,{\widehat{h}_{x}})}_{j,k}\}_{k=1}^{j}$. We use the sample values $\{Y_{i}:X_{i}\in\bigcup_{|h-{\widehat{h}_{x}}|\leqslant 1}\widehat{S}_{l,h}\}$ to construct a local polynomial ${\widehat{f}_{j,k|{\widehat{v}_{l,\widehat{h}_{x}}}}}$ on each interval $I^{(l,{\widehat{h}_{x}})}_{j,k}$ of the partition by solving a least squares fitting problem, obtaining a piecewise polynomial estimator ${\widehat{f}_{j|{\widehat{v}_{l,\widehat{h}_{x}}}}}$ of the link function. The degree of the local polynomials needed to obtain optimal estimation rates depends on the regularity of the function. A proper partition (or scale) is then chosen to minimize the expected mean squared error (MSE) using classical bias-variance trade-off arguments. Composing this local estimator of $f$ with the “nearest” slice assignment gives us a global estimator $F(x)={\widehat{f}_{j|{\widehat{v}_{l,\widehat{h}_{x}}}}}(\langle{\widehat{v}_{l,\widehat{h}_{x}}},x\rangle)$.
Input :  Samples $\{(X_{i},Y_{i})\}_{i=1}^{n}\subseteq\mathbb{R}^{d}\times\mathbb{R}$, number of partitions $l,j\in\mathbb{N}$, polynomial degree $m\in\mathbb{N}$ , truncation level $M\in(0,\infty]$.
Output :  ${\widehat{F}}$ estimate of $F$.
  1. 1.a
    Compute interval $R$ of range of samples $\{Y_{i}\}_{i=1}^{n}$. Construct $\{{R_{l,h}}\}_{h=1}^{l}$, the uniform partition of $R$ into $l$ intervals whose preimages are $\{\widehat{S}_{l,h}\}_{h=1}^{l}$ where $\widehat{S}_{l,h}=\{X_{i}:Y_{i}\in{R_{l,h}}\}$;
  2. 1.b
    Denote ${n_{l,h}}=\#\{Y_{i}\in{R_{l,h}}\}$. For each $h\in\mathcal{H}_{l}=\{h:{n_{l,h}}\geq n/l\}$, compute:
${\widehat{\mu}_{l,h}}=\frac{1}{{n_{l,h}}}\sum_{i}X_{i}\mathbbm{1}\{X_{i}\in\widehat{S}_{l,h}\}$, the empirical mean
${\widehat{\Sigma}_{l,h}}=\frac{1}{{n_{l,h}}}\sum_{i}(X_{i}-{\widehat{\mu}_{l,h}})(X_{i}-{\widehat{\mu}_{l,h}})^{\intercal}\mathbbm{1}\{X_{i}\in\widehat{S}_{l,h}\}$, the empirical covariance matrix
${\widehat{\lambda}_{l,h,m}}$, the $m$-th eigenvalue of ${\widehat{\Sigma}_{l,h}}$, in descending order
${\widehat{v}}_{l,h,m}$, the $m$-th eigenvector of ${\widehat{\Sigma}_{l,h}}$ corresponding to eigenvalue ${\widehat{\lambda}_{l,h,m}}$
$H_{l,h}=\log\left(\frac{({\widehat{\lambda}_{l,h,2}}\times\cdots\times{\widehat{\lambda}_{l,h,d-1}})^{{2}/{(d-2)}}}{{\widehat{\lambda}_{l,h,1}}{\widehat{\lambda}_{l,h,d}}}\right)$
the empirical significant vector ${\widehat{v}_{l,h}}$, equals ${\widehat{v}}_{l,h,d}$ if $H_{l,h}>0$ and to ${\widehat{v}}_{l,h,1}$ otherwise
  • 2.a
    Given $x\in\mathbb{R}^{d}$, for each $h\in\mathcal{H}_{l}$ compute the estimated distance between $x$ and $\widehat{S}_{l,h}$
    ${\widehat{\mathrm{dist}}}(x,h)=\begin{cases}|\langle x-{\widehat{\mu}_{l,h}},{\widehat{v}_{l,h}}\rangle|^{2}+\frac{{\lambda_{d}(\widehat{\Sigma}_{l,h})}}{{\lambda_{1}(\widehat{\Sigma}_{l,h})}}\left\|x-{\widehat{\mu}_{l,h}}\right\|^{2}\ \text{ if }\widehat{H}_{l,h}>0\\ \left\|x-{\widehat{\mu}_{l,h}}\right\|^{2}+\frac{{\lambda_{d}(\widehat{\Sigma}_{l,h})}}{{\lambda_{1}(\widehat{\Sigma}_{l,h})}}|\langle x-{\widehat{\mu}_{l,h}},{\widehat{v}_{l,h}}\rangle|^{2}\ \ \text{ if }\widehat{H}_{l,h}<0\end{cases}$
  • 2.b
    compute the estimated nearest slice index ${\widehat{h}_{x}}=\operatorname*{{arg\,min}}_{h\in\mathcal{H}_{l}}{\widehat{\mathrm{dist}}}(x,h)$
  • 3.a
    For each $h\in\mathcal{H}_{l}$ compute: the interval $I^{(l,h)}$ containing $\{\langle{\widehat{v}_{l,h}},X_{i}\rangle:X_{i}\in\bigcup_{|h-{\widehat{h}_{x}}|\leqslant 1}\widehat{S}_{l,h}\}$; the uniform partition $\{{{I^{(l,h)}_{j,k}}}\}_{k=1}^{j}$ of $I^{(l,h)}$; $n^{(l,h)}_{j,k}=\#\{X_{i}:\langle{\widehat{v}_{l,h}},X_{i}\rangle\in{I^{(l,h)}_{j,k}}\}$.
  • 3.b
    For each $h\in\mathcal{H}_{l}$ and each $k\in\mathcal{K}_{j}=\{k:n^{(l,h)}_{j,k}\geqslant{n_{l,h}}/j\}$ compute
    ${\widehat{f}_{j,k|{\widehat{v}_{l,h}}}}=\operatorname*{{arg\,min}}_{\deg(p)\leqslant m}\sum_{X_{i}\in\bigcup_{|h-{\widehat{h}_{x}}|\leqslant 1}\widehat{S}_{l,h}}\left|Y_{i}-p(\langle{\widehat{v}_{l,h}},X_{i}\rangle)\right|^{2}\mathbbm{1}_{{I^{(l,h)}_{j,k}}}(\langle{\widehat{v}_{l,h}},X_{i}\rangle)$
  • 3.c
    For each $h\in\mathcal{H}_{l}$, compute ${\widehat{f}_{j|{\widehat{v}_{l,h}}}}(r)=\sum_{k\in\mathcal{K}_{j}}{\widehat{f}_{j,k|{\widehat{v}_{l,h}}}}(r)\mathbbm{1}_{{I^{(l,h)}_{j,k}}}(r)$ and return the estimator
    ${\widehat{F}}(x)={\widehat{f}_{j|{\widehat{v}_{l,\widehat{h}_{x}}}}}(\langle{\widehat{v}_{l,\widehat{h}_{x}}},x\rangle)$
  • Algorithm 1 Significant Vector Regression
    2.4 The main algorithm and guarantees on the estimator it produces
    Algorithm 1 summarizes the construction of the proposed estimator of the regression function $F$. We prove that the mean squared error of our estimator is the sum of the estimation error and the curve approximation error. The estimation error decays the one-dimensional min-max optimal nonparametric learning rate, up to a saturation level dependent on quantities expected to be small, or even zero, as we discuss in detail below.
    Theorem 3 (MSE of the Estimator constructed by Algorithm 1)
    Assume that (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), ($\bm{\omega_{f}}$), and (P) hold true, and that $f\in\mathcal{C}^{s}(\mathbb{R}^{1})$ with $s\in\left[\frac{1}{2},2\right]$. Let $C_{\gamma,f},M^{*},l_{\max}$ be the constants specified in (4) below, and $C$ a universal constant. Then, for $n$ such that $\frac{n}{\log^{{3/2}}n}\gtrsim C_{\gamma,f}\frac{C_{f}\mathrm{len}_{\gamma}}{C_{f}^{\prime}\sigma_{\gamma}}$, if we choose $l^{*},j^{*}$ as
    $(l^{*},j^{*}):=\begin{cases}\left(C_{\gamma,f}^{-1}n\log^{-2}n,\,C\right)&\text{ if }\frac{n^{\frac{2s}{2s+1}}}{\log^{2}n}\lesssim C_{\gamma,f}M^{*}\\ \left(l_{\max},\,C\frac{M^{*}}{l_{\max}}n^{\frac{1}{2s+1}}\right)&\text{ if }n^{\frac{1}{2s+1}}\gtrsim\frac{CC_{Y}R_{0}}{M^{*}\max(\sigma_{\zeta},\omega_{f})}\\ \left(n^{\frac{1}{2s+1}}M^{*},\,C\right)&\text{ otherwise}\end{cases}\,,$
    the estimator constructed by Algorithm 1, in running time $\mathcal{O}(d^{2}n\log n)$, satisfies
    $\mathbb{E}\left[\left|{\widehat{F}}(X)-F(X)\right|^{2}\right]\lesssim C_{1}(f,\gamma,\rho_{X},\sigma_{\zeta},d)n^{-\frac{2s}{2s+1}}{\log n}+C_{2}(f,\gamma,\rho_{X},\sigma_{\zeta},d)$
    We specify the constants in the Theorem above:
    $C_{1}(f,\gamma,\rho_{X},\sigma_{\zeta},d)$ $:=\left(\frac{C_{f}}{C_{f}^{\prime}}\mathrm{len}_{\gamma}\sigma_{\zeta}^{2}\right)^{\frac{2s}{2s+1}}\!\!\!\!\!\!\!\!\!(\left[f\right]_{\mathcal{C}^{s}}\!+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s}})^{\frac{2}{2s+1}}\,,$ (4)
    $C_{2}(f,\gamma,\rho_{X},\sigma_{\zeta},d)$ $:=\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\!\!\left(\frac{\sigma_{\gamma}C_{f}}{\mathrm{reach}_{\gamma}}\max(\sigma_{\zeta},\omega_{f})\right)^{2(s\wedge 1)}\!\!\!\!\!,\,\,C_{\gamma,f}:=\frac{R_{0}^{3}C_{f}^{2}}{C_{f}^{\prime}}\max\left(\frac{\mathrm{len}_{\gamma}d^{3/2}}{\sigma_{\gamma}^{4}},\frac{R_{0}^{5}C_{f}^{2}d^{4}}{C_{f}^{\prime 3}\sigma_{\gamma}^{8}}\right)\,,\,$
    $M^{*}$ $:=\left(\sigma_{\zeta}^{-1}(C_{f}C_{Y}R_{0})^{s}(\left[f\right]_{\mathcal{C}^{s}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s}})\right)^{\frac{2}{2s+1}}\quad,\quad l_{\max}:=\frac{CC_{Y}R_{0}}{\max(\sigma_{\zeta},\omega_{f})}\,.$
    The above theorem states that our estimator achieves the min-max optimal rate for the one-dimensional nonparametric regression, with an additional approximation error of magnitude $\mathcal{O}(\sigma_{\zeta}^{2})$ for functions that are both Lipschitz and monotone, therefore defeating the curse of dimensionality by exploiting the compositional structure of $F$, even if the inner function is nonlinear (unlike the single- and multi-index model) and not particularly smooth (and its regularity does not scale with the ambient dimension). It is worth mentioning that the result of Theorem 3 is scaling invariant in $X$ and $Y$.
    This result is satisfactory in the following respects:
    1. (i)
      it avoids the curse of dimensionality, with $d$ not appearing in the learning rate $\frac{2s}{2s+1}$;
    2. (ii)
      the learning rate matches the min-max rate for $1$-d nonparametric regression;
    3. (iii)
      the minimal number of samples $n$ required is only a low-order polynomial in $d$, $\mathrm{len}_{\gamma}$, and $\left\|\gamma^{\prime\prime}\right\|$. This is not exponential in ambient dimension $d$ (unless in the extreme case where $\gamma$ has length or curvature growing exponentially in $d$, e.g. a space-filling curve).
    4. (iv)
      Algorithm 1 in fact also estimates $\gamma$, the nonlinear projection $\Pi_{\gamma}$, and $f$. In each local neighborhood of $\widehat{S}_{l,h}$, the empirical mean ${\widehat{\mu}_{l,h}}$ and significant vector ${\widehat{v}_{l,h}}$ give a line segment that approximates $\gamma$, and $\Pi_{\gamma}$ is estimated, up to a translation, by the piecewise linear approximation $x\mapsto\langle x,{\widehat{v}_{l,h}}\rangle$. This provides interpretability to our estimator, in the sense that both $F$ and its structure are resolved; this is not the case for neural networks, even in the case of single-index models, where the index can be identified only by suitably averaging over multiple weights in multiple nodes.
    There are two apparent shortcomings in our results:
    1. (i)
      the assumption that $f$ is coarsely monotone. While similar conditions have appeared in other works using conditional regression (see Lanteri et al., 2022, for a discussion), and even in recent approaches based on gradient descent for the single- and multi-index models (see e.g., Damian et al., 2024; Arous et al., 2021; Bietti et al., 2023, and references therein) the “amount of oscillation” of $f$ appears prominently and imposes additional sampling requirements (e.g., scaling as $d^{\mathcal{O}(L)}$ where $L$ is the number of vanishing moments in the Hermite polynomial basis, albeit it is not clear to us that such assumptions and sampling requirements are sharp);
    2. (ii)
      the second shortcoming is that we have a second additive term, which we call curve approximation error, which does not vanish as $n$ tends to infinity. It is typically small, e.g., if external noise $\sigma_{\zeta}$ or the curvature $\left\|\gamma^{\prime\prime}\right\|$ is small, and in fact it will vanish when external noise $\sigma_{\zeta}$ or curvature $\left\|\gamma^{\prime\prime}\right\|$ vanishes. In the limiting case where the underlying curve $\gamma$ is a straight line segment, we have no curve approximation error, recovering the results of (Lanteri et al., 2022). Otherwise, the nonzero approximation error exists because we are using first-order approximations in the estimation of the local directions of the curve, and with noisy observations our technique for constructing the estimator does not allow us to consider local pieces below the scale of the noise. While it seems possible to use approximations of higher order, they would require significant additional computational complexity and more refined statistical analysis, possibly with minimal impact in practical applications, and are left to future work.
    In the case of a strictly monotone Lipschitz function with zero external noise, we obtain the rate $\mathcal{O}(n^{-2})$. Here, the curve approximation error vanishes because there is no upper bound for the number of slice $l$, contrary to the noisy case in Theorem 3:
    Theorem 4 (MSE of Algorithm 1 in the noiseless case)
    Assume that (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), ($\bm{\omega_{f}}$), and (P) hold true. Assume that there is no observational noise, i.e., $\zeta\equiv 0$ almost surely, that the link function $f$ is perfectly monotone, i.e., $\omega_{f}=0$, and that $f\in\mathcal{C}^{s}$ with $s\in\left[\frac{1}{2},2\right]$. With $C_{\gamma,f}$ in (4), when $\frac{n}{\log^{3/2}n}\gtrsim C_{\gamma,f}\frac{C_{f}\mathrm{len}_{\gamma}}{C_{f}^{\prime}\sigma_{\gamma}}$, if we choose $l^{*}=\frac{1}{C_{\gamma,f}}\frac{n}{\log^{3/2}n}$ and $j^{*}=C$, then the estimation error of Algorithm 1 satisfies
    $\mathbb{E}\left[\left|{\widehat{F}}(X)-F(X)\right|^{2}\right]\lesssim(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}\left(C_{\gamma,f}(C_{f}C_{Y}R_{0})^{2}\frac{\log^{3}n}{n^{2}}\right)^{s\wedge 1}\,.$
    The better rate $\mathcal{O}((n^{-2}\log^{3}n)^{(s\wedge 1)})$ is a consequence of having zero observational noise. Bauer et al. (2017) proves that, in the case of the $L^{\infty}$-norm, the min-max rate of nonparametric regression for functions in $\mathcal{C}^{s}(\mathbb{R}^{d})$ with noiseless observations is $(\log n/n)^{s/d}$. The rate on the mean squared error in Theorem 4 agrees with this when $s\leqslant 1$ and $d=1$, as if $\gamma$ (and therefore $\Pi_{\gamma}$) were known. Our rate is suboptimal for $s>1$, as we only perform a first-order approximation of the underlying curve and do not estimate higher-order parameters such as the curvature. Our estimator avoids the curse of dimensionality by exploiting the compositional structure of $F$, even if the inner function is (unknown and) nonlinear (unlike in the single- and multi-index model) and not particularly smooth (and its regularity does not scale with the ambient dimension). A key difference between Theorem 3 and Theorem 4 is that there is no curve approximation error in the latter: noiseless observations allow us to perform an unlimited amount of partitioning, obtaining “thin” slices, provided that enough samples are available (in order to control the variance of the objects we estimate in each slice).
    When (LCV) is not satisfied, our estimator quickly saturates at the level of the curve approximation error:
    Theorem 5 (NVM without (LCV))
    Assume that (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), ($\bm{\omega_{f}}$), and (SC) hold true. When $\frac{n}{\log^{1/2}n}\gtrsim\frac{C_{Y}R_{0}^{7}d^{3}}{C_{f}^{6}\max(\sigma_{\zeta},\omega_{f})^{7}}$, if we choose $l^{*}=\frac{C_{Y}R_{0}}{\max(\sigma_{\zeta},\omega_{f})}$ and $j^{*}=C$, the estimation error of Algorithm 1 satisfies
    $\mathbb{E}\left[\left|{\widehat{F}}(X)-F(X)\right|^{2}\right]\lesssim\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\left(\frac{\sigma_{\gamma}C_{f}}{\mathrm{reach}_{\gamma}}\max(\sigma_{\zeta},\omega_{f})\right)^{2(s\wedge 1)}\,.$
    3 Analysis of the Estimator
    We introduce several properties that our model may have and will be assumed, in various combinations, in our results. We start by collecting a few conditions on the distribution of $X,Y$, and $\zeta$ that are fairly standard:
    1. (X subG)
      $X$ has sub-Gaussian distribution with variance proxy $R_{0}^{2}$, and has a density function $\rho_{X}$ with $\left[\rho_{X}\right]_{\mathcal{C}^{2}}<\infty$.
    1. (Y subG)
      $Y$ has sub-Gaussian distribution with variance proxy $C_{Y}^{2}R_{0}^{2}$.
    1. ($\bm{\zeta}$ subG)
      $\zeta$ is sub-Gaussian with variance proxy $\sigma_{\zeta}^{2}$.
    Recall that $X$ can be decomposed as position along the underlying curve and deviation away from the curve $\gamma$ as in (1). The following assumption on $Z_{d-1}$ considers how random vector $X$ deviates off the underlying curve $\gamma$:
    1. ($\bm{\gamma_{1}}$)
      $\gamma$ has a Lipschitz derivative $\gamma^{\prime}:[0,\mathrm{len}_{\gamma}]\to\mathbb{R}^{d}$. For each $t_{0}\in[0,\mathrm{len}_{\gamma}]$, the conditional random vector $Z_{d-1}|t=t_{0}$ is mean zero, isotropic with variance $\sigma_{\gamma}(t_{0})^{2}\mathbb{I}_{d-1}$, and supported in an Euclidean ball $B(\gamma(t_{0}),c\ \mathrm{reach}_{\gamma}(t_{0}))\subseteq\mathbb{R}^{d-1}$ for some $c<1$.
    Assumption ($\bm{\gamma_{1}}$) implies that for any $t_{0}\in[0,\mathrm{len}_{\gamma}]$, the conditional mean is on the curve
    $\mu_{t_{0}}:=\mathbb{E}[X\mid t=t_{0}]=\gamma(t_{0})\,,$
    and the conditional covariance matrix has eigenvalue 0 on the eigenspace $\operatorname{span}\{\gamma^{\prime}(t_{0})\}$ and eigenvalue $\sigma_{\gamma}(t_{0})$ on eigenspace $\operatorname{span}\{\gamma^{\prime}(t_{0})\}^{\perp}$, since
    $\Sigma_{t_{0}}$ $:=\mathbb{E}[(X-\mu_{t_{0}})(X-\mu_{t_{0}})^{\intercal}\mid t=t_{0}]=\sigma_{\gamma}(t_{0})^{2}\mathbb{I}_{d}-\sigma_{\gamma}(t_{0})^{2}\gamma^{\prime}(t_{0})\gamma^{\prime}(t_{0})^{\intercal}\,.$
    Because the underlying curve $\gamma$ is unknown, we cannot condition on $t=\Pi_{\gamma}X$. Since we condition over the sample value $Y_{i}$ from data $(X_{i},Y_{i})$, we need a property that partially reveals the “one-to-one” correspondence between $t=\Pi_{\gamma}(X)$ and $F(X)=f(\Pi_{\gamma}X)$:
    1. ($\bm{\omega_{f}}$)
      There exist constants $\omega_{f}\geqslant 0$ and $C_{f}>C_{f}^{\prime}>0$ that only depend on the link function $f$ such that, for every interval $T$ with $|T|\geqslant\omega_{f}$, we have
      $C_{f}^{\prime}|T|\leqslant\left|\left[\min f^{-1}(T),\max f^{-1}(T)\right]\right|\leqslant C_{f}|T|\,.$
    Assumption ($\bm{\omega_{f}}$) may be regarded as a large-scale sub-Lipschitz property. If $f$ is bi-Lipschitz, and therefore in particular monotone, then ($\bm{\omega_{f}}$) is satisfied with $\omega_{f}=0$. However, ($\bm{\omega_{f}}$) for $\omega_{f}>0$ does not imply that $f$ is monotone: it relaxes monotonicity to monotonicity “at scales larger than $\omega_{f}$”, and thus we say that $f$ is “coarsely monotone”.
    The following assumption gives a lower bound on the conditional variance:
    1. (LCV)
      Define $\sigma_{\gamma}:=\min_{t_{0}\in[0,\mathrm{len}_{\gamma}]}\sigma_{\gamma}(t_{0})$ as the minimum value of $\sigma_{\gamma}(t_{0})$. We assume that $\sigma_{\gamma}\geqslant 2C_{f}\max(\sigma_{\zeta},\omega_{f})$.
    The purpose of Assumption (LCV) is that for any interval $T\subseteq[0,\mathrm{len}_{\gamma}]$, it allows us to compute the conditional mean
    $\mu_{T}:=\mathbb{E}[X\mid t\in T]=\mathbb{E}_{t}[\gamma(t)\mid t\in T]$
    and conditional covariance matrix
    $\Sigma_{T}$ $:=\mathbb{E}[(X-\mu_{T})(X-\mu_{T})^{\intercal}\mid t\in T]$
    $=\mathbb{E}\left[(\gamma(t)-\mu)(\gamma(t)-\mu)^{\intercal}\mid t\in T\right]+\mathbb{E}[\sigma_{\gamma}(t)^{2}\mid t\in T]\mathbb{I}_{d-1}-\mathbb{E}[\sigma_{\gamma}(t)^{2}\gamma^{\prime}(t)\gamma^{\prime}(t)^{\intercal}\mid t\in T]$
    The above identity illustrates that when slices are thin enough, Significant Vector Regression is approximately estimating the tangential vector $\gamma^{\prime}$: if $T$ is small enough such that the first term has a negligible spectral norm, compared with the second and the third term, because the second term is a multiple of identity matrix and the third term has negative sign, the smallest principal component of $\Sigma_{T}$ is roughly the largest principal component of $\mathbb{E}[\sigma_{\gamma}(t)^{2}\gamma^{\prime}(t)\gamma^{\prime}(t)^{\intercal}\mid t\in T]$, yielding an estimate of the direction of $\mathbb{E}[\sigma_{\gamma}(t)\gamma^{\prime}(t)\mid t\in T]$.
    Given the above, it is natural that the “thinness” of the slice is desirable: this requires a small interval $T$ and a lower bound on $\sigma_{\gamma}$: this motivates Assumption (LCV), since scales below the noise level $\sigma_{\zeta}$ of the observed $Y_{i}$ or below the rough monotonicity scale $\omega_{f}$ are not valuable for our inverse regression approach.
    Finally, we consider the following property, which is satisfied by various regular densities on compact normal domains with no geometrical symmetry required:
    1. (P)
      The density $\rho_{X}$ has compact support. Let $\rho_{v}$ denote the push-forward measure of the distribution of $X$ along the map $x\mapsto\langle x,v\rangle$. Then there exists constant $c>0$ such that, for every $v\in\mathbb{S}^{d-1}$ and for every interval $I\subseteq\operatorname{\mathrm{supp}}\rho_{v}$, we have $\rho_{v}(I)\geqslant c|I|$ and $\mathrm{Var}[\langle v,X\rangle|\langle v,X\rangle\in I]\geqslant c|I|^{2}$.
    Assumption ($\bm{\gamma_{1}}$) implies that for $t_{0}\in[0,\mathrm{len}_{\gamma}]$, $\sigma_{\gamma}(t_{0})<\min(R_{0},\mathrm{reach}_{\gamma}/\sqrt{d})$; (LCV) implies that $\sigma_{\gamma}\geqslant 2C_{f}\sigma_{\zeta}$; ($\bm{\omega_{f}}$) implies that ${C_{f}^{\prime}}C_{Y}R_{0}\leqslant\mathrm{len}_{\gamma}\leqslant C_{f}C_{Y}R_{0}$.
    3.1 Estimation of slice parameters with Assumption (LCV)
    Consider the event of bounded data
    :=$\mathcal{B}:=\left\{\left\|X\right\|\leqslant C_{X}\sqrt{d}R_{0},|Y|\leqslant C_{Y}R_{0}\right\}$
    for some $C_{X},C_{Y}\geqslant 1$ fixed constant from now on. We define the following bounded version of ${\mu_{l,h}}$ and $\Sigma_{l,h}$:
    ${\mu_{l,h}^{b}}:=\mathbb{E}\left[X|Y\in{R_{l,h}},\mathcal{B}\right],\ \ \ {\Sigma_{l,h}^{b}}:=\operatorname{{Cov}}\left[X|Y\in{R_{l,h}},\mathcal{B}\right]$
    Given the event
    $\mathcal{B}_{i}:=\{\left\|X_{i}\right\|\leqslant C_{X}\sqrt{d}R_{0},|Y_{i}|\leqslant C_{Y}R_{0}\}$
    we define
    $n^{b}:=\sum_{i}\mathbbm{1}_{\mathcal{B}_{i}}(X_{i}),\ \ \ {n_{l,h}^{b}}:=\sum_{i}\mathbbm{1}(\left\{Y_{i}\in{R_{l,h}}\right\}\cap\mathcal{B}_{i})\,.$
    The random variable $n^{b}$, assuming (X subG) and (Y subG), is larger than a constant fraction of $n$ with high probability (see Lemma 21). The sample counterparts of ${\mu_{l,h}^{b}}$ and ${\Sigma_{l,h}^{b}}$ are
    ${\widehat{\mu}_{l,h}^{b}}:=\frac{1}{{n_{l,h}^{b}}}\sum_{i:\left\{Y_{i}\in{R_{l,h}}\right\}\cap\mathcal{B}_{i}}X_{i}\,,\,{\widehat{\Sigma}_{l,h}^{b}}:=\frac{1}{{n_{l,h}^{b}}}\sum_{i:\left\{Y_{i}\in{R_{l,h}}\right\}\cap\mathcal{B}_{i}}(X_{i}-{\widehat{\mu}_{l,h}^{b}})(X_{i}-{\widehat{\mu}_{l,h}^{b}})^{\intercal}\,.$
    We denote by ${v_{l,h}^{b}}$ and ${\widehat{v}_{l,h}^{b}}$ the significant vector of ${\Sigma_{l,h}^{b}}$ and ${\widehat{\Sigma}_{l,h}^{b}}$, respectively. Finally, let ${\mathcal{H}_{l}^{b}}:=\{h:{n_{l,h}^{b}}\geqslant\frac{n^{b}}{l}\}$. Given $\mathcal{B}$, it is natural to pick the interval $R$ in Significant Vector Regression as
    $R:=[-C_{Y}R_{0},C_{Y}R_{0}]$
    Taking slices $Y\in{R_{l,h}}$ is equivalent to conditioning on $Y\in{R_{l,h}}$. In this procedure, we obtain information on conditional random variables such as the slice center ${\mu_{l,h}^{b}}:=\mathbb{E}[X|Y\in{R_{l,h}},\mathcal{B}]$, the slice covariance matrix ${\Sigma_{l,h}^{b}}$, and the slice significant vector ${v_{l,h}^{b}}$. Moreover, the eigenvalues $\lambda_{1}({\Sigma_{l,h}^{b}}),\dots,\lambda_{d}({\Sigma_{l,h}^{b}})$ determine features of the geometric shape for the slice ${\Sigma_{l,h}^{b}}$; in particular, the smallest eigenvalue $\lambda_{d}({\Sigma_{l,h}^{b}})$ determines the “width” of slice ${S_{l,h}^{b}}$. We will show that these parameters can be estimated with small errors given a moderate sample size. Recall that $\Pi_{\gamma}:\mathbb{R}^{d}\to[0,L]$ maps points to the one-dimensional interval $[0,L]$, that encodes the position along curve $\gamma$. We start with a proposition on estimating slice position on curve $\Pi_{\gamma}X|Y\in T,\mathcal{B}$.
    Proposition 6
    Suppose ($\bm{\zeta}$ subG) and ($\bm{\omega_{f}}$) hold true. Let $T\subseteq R$ be a bounded interval with $|T|\geqslant\max(\sigma_{\zeta},\omega_{f})$. Then:
    1. (a)
      For every $i=1,\dots,n$ and every $\tau\geqslant 1$
      $\mathbb{P}\left\{\left|\Pi_{\gamma}X_{i}-\mathbb{E}\left[\Pi_{\gamma}X|Y\in T,\mathcal{B}\right]\right|\gtrsim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})\mid Y_{i}\in T,\mathcal{B}_{i}\right\}\leqslant 2n^{-\tau}\,.$
    2. (b)
      $\mathrm{Var}\left[\Pi_{\gamma}X|Y\in T,\mathcal{B}\right]\lesssim C_{f}^{2}(|T|^{2}+\sigma_{\zeta}^{2})$.
    We now bound the estimation error for the tangential direction and the smallest eigenvalue $\lambda_{d}({\Sigma_{l,h}^{b}})$:
    Proposition 7
    Suppose ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$) and ($\bm{\omega_{f}}$) hold true. Let $T\subseteq R$ be a bounded interval with $|T|\geqslant\max(\sigma_{\zeta},\omega_{f})$. Then:
    1. (a)
      For every $i=1,\dots,n$ and every $\tau\geqslant 1$, we have
      $\mathbb{P}\left\{\left|\langle{v_{l,h}^{b}},X_{i}\rangle-\mathbb{E}[\langle{v_{l,h}^{b}},X\rangle|Y\in T,\mathcal{B}]\right|\gtrsim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})\mid Y_{i}\in T,\mathcal{B}_{i}\right\}\leqslant 2n^{-\tau}\,;$
    2. (b)
      $\lambda_{d}({\Sigma_{l,h}^{b}})=\mathrm{Var}\left[\langle{v_{l,h}^{b}},X\rangle\mid Y\in T,\mathcal{B}\right]\lesssim C_{f}^{2}(|T|^{2}+\sigma_{\zeta}^{2})$.
    We now show under which assumptions $\lambda_{d}({\Sigma_{l,h}^{b}})$ is small compared to $\lambda_{d-1}({\Sigma_{l,h}^{b}})$ and the other eigenvalues, yielding the “thin slice scenario”:
    Corollary 8
    Suppose (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), and ($\bm{\omega_{f}}$) hold true. Then, for every $l$ such that $l\gtrsim C_{f}C_{Y}R_{0}/\sigma_{\gamma}$, and $|{R_{l,h}^{b}}|\geqslant\max(\sigma_{\zeta},\omega_{f}),h\in\mathcal{H}_{l}^{b}$, we have
    $\lambda_{d-1}({\Sigma_{l,h}^{b}})-\lambda_{d}({\Sigma_{l,h}^{b}})\gtrsim\sigma_{\gamma}^{2}\,.$
    In part 1. b) of Algorithm 1, we compute on each slice its sample mean ${\widehat{\mu}_{l,h}^{b}}$, sample covariance matrix ${\widehat{\Sigma}_{l,h}^{b}}$, eigenvalues ${\widehat{\lambda}_{l,h,m}}$ and eigenvectors ${\widehat{v}}_{l,h,m}$ of the sample covariance matrix. It is natural to ask how accurately these parameters can be estimated: we address this in Proposition 17 and Lemma 16. These results are technical and postponed to the appendix; here, we record that they yield the following corollary that gives the expected near $\sqrt{n_{loc}}$-consistency in estimating the parameters in each slice, where $n_{loc}$ is the number of samples per slice:
    Corollary 9
    Suppose (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), and ($\bm{\omega_{f}}$) hold true. Then, for every $l$ such that $l\gtrsim C_{f}C_{Y}R_{0}/\sigma_{\gamma}$, and $|{R_{l,h}^{b}}|\geqslant\max(\sigma_{\zeta},\omega_{f})$ for all $h\in\mathcal{H}_{l}^{b}$, for every $\epsilon>0$ and $\tau\geqslant 1$, if $n$ is sufficiently large so that $\frac{n}{\sqrt{\tau\log n}}\gtrsim\left(\frac{C_{f}C_{Y}R_{0}}{\sigma_{\gamma}}\right)^{2}dl(t+\log d+\log l)$, we have
    $\mathbb{P}\left\{\exists h\in{\mathcal{H}_{l}^{b}}:\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|\gtrsim C_{Y}C_{f}R_{0}\sqrt{t+\log l+\log d}\sqrt{\frac{\sqrt{\tau\log n}}{nl}}\ \bigg{|}\ {n_{l,h}^{b}}\right\}\lesssim e^{-t}+n^{-\tau}$
    $\mathbb{P}\left\{\exists h\in{\mathcal{H}_{l}^{b}}:\left\|{{\widehat{v}_{l,h}^{b}}}-{{v_{l,h}^{b}}}\right\|\gtrsim C_{Y}C_{f}R_{0}^{2}\sigma_{\gamma}^{-2}\sqrt{t+\log l+\log d}\sqrt{\frac{d\sqrt{\tau\log n}}{nl}}\ \bigg{|}\ {n_{l,h}^{b}}\right\}\lesssim e^{-t}+n^{-\tau}$
    Moreover, if $\frac{n}{\log^{3/2}n}\gtrsim C_{Y}^{2}C_{f}^{2}R_{0}^{4}\sigma_{\gamma}^{-4}dl(\log d+\log l)$, then for any $h\in{\mathcal{H}_{l}^{b}}$ and $p\geq 1$,
    $\mathbb{E}\left[\left\|{\widehat{v}_{l,h}^{b}}-{v_{l,h}^{b}}\right\|^{2p}\right]\lesssim C(p)(C_{Y}C_{f}R_{0}^{2}\sigma_{\gamma}^{-2})^{2p}(d\log d)^{p}\left(\frac{\log n\log l}{nl}\right)^{p}$
    3.2 Estimation of the distance function and classification accuracy
    Here we assume (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), and ($\bm{\omega_{f}}$): by Corollary 8 we are in “thin” slice scenario, i.e. $H_{l,h}>0$, and the distance function simplifies to ${\widehat{\mathrm{dist}}}(x,h)=|\langle x-{\widehat{\mu}_{l,h}^{b}},{\widehat{v}_{l,h}^{b}}\rangle|^{2}+\frac{{\lambda_{d}(\widehat{\Sigma}_{l,h}^{b})}}{{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}}\left\|x-{\widehat{\mu}_{l,h}^{b}}\right\|^{2}$. In Algorithm 1, we take the slice ${\widehat{h}_{x}}$, which has the smallest estimated distance, to $x$, as an estimator of the true correct ${h^{\prime}_{x}}$. The following proposition states that the population counterpart ${h_{x}}$ of the nearest index almost equals the correct index ${h^{\prime}_{x}}$.
    Proposition 10 (nearest index is almost correct)
    Let ${h_{x}}=\operatorname*{{arg\,min}}_{h\in\mathcal{H}_{l}}{\mathrm{dist}}(x,h)$ be the nearest index and define the correct index ${h^{\prime}_{x}}$ be the unique $h\in\mathcal{H}_{l}$ such that $F(x)\in{R_{l,h}}$. Suppose ($\bm{\zeta}$ subG), ($\bm{\omega_{f}}$), and (LCV) hold true. Suppose that $|{R_{l,h}}|\geqslant\max(\sigma_{\zeta},\omega_{f})$ for all $h\in\mathcal{H}_{l}$. Then $|{h_{x}}-{h^{\prime}_{x}}|\leqslant 1$. Moreover, the phenomenon of adjacent misclassification (i.e., $|{h_{x}}-{h^{\prime}_{x}}|=1$) only occurs for points near the boundary of some slices.
    In Algorithm 1, we use the sample slice with index ${\widehat{h}_{x}}$, which has the smallest estimated distance to $x$, to estimate the correct ${h^{\prime}_{x}}$: this gives the correct/adjacent classification w.h.p.:
    Proposition 11 (classification accuracy)
    Assume (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), and ($\bm{\omega_{f}}$) hold true. If $l\gtrsim\frac{C_{f}\mathrm{len}_{\gamma}}{C_{f}^{\prime}\sigma_{\gamma}}$, and $|{R_{l,h}}|\geqslant\max(\sigma,\omega_{f})$ for all $h\in\mathcal{H}_{l}^{b}$, then the probability of misclassification by at least two slices in part 2.b of Algorithm 1 satisfies
    $\mathbb{P}\left(\left|{\widehat{h}_{x}}-{h_{x}}\right|\geqslant 2\right)\lesssim ld\exp\left(-c\frac{nl^{-1}}{\sqrt{\tau\log n}}\min\left(\frac{C_{f}^{\prime}\sigma_{\gamma}^{4}}{C_{f}^{2}R_{0}^{3}d^{3/2}\mathrm{len}_{\gamma}},\frac{\sigma_{\gamma}^{8}C_{f}^{\prime 4}}{R_{0}^{8}C_{f}^{4}d^{4}}\right)\right)+ln^{-\tau}\,.$
    As we see from the above inequality, the probability of misclassification decays at least linearly in $\frac{n}{\sqrt{\tau\log n}}$, as long as $n$ and $l$ satisfy
    $\frac{n/l}{\log^{3/2}n}\gtrsim C_{\gamma,f}:=\frac{R_{0}^{3}C_{f}^{2}}{C_{f}^{\prime}}\max\left(\frac{d^{3/2}\mathrm{len}_{\gamma}}{\sigma_{\gamma}^{4}},\frac{R_{0}^{5}C_{f}^{2}d^{4}}{C_{f}^{\prime 3}\sigma_{\gamma}^{8}}\right)\quad\text{ and }\quad l\gtrsim\frac{C_{f}\mathrm{len}_{\gamma}}{C_{f}^{\prime}\sigma_{\gamma}}\,.$ (5)
    We conclude that for $n$ large the estimation error corresponding to misclassification by at least two slices is negligible compared with the error corresponding to correct classification. Similar to Proposition 11, the phenomenon of misclassification to adjacent slices seems inevitable for points near the boundary of some slices. To prevent this effect from undermining the performance of the estimator ${\widehat{F}}$, we will include data from adjacent slices in the regression of the link function $f$ in each local neighborhood after the projection onto the local tangent to $\gamma$ (as in step 3.b of Algorithm 1).
    3.3 Function estimation error corresponding to almost correct classification
    In this subsection, we work on the event that we classify $x\in\Omega_{\gamma}$ to the empirical slice $\widehat{S}_{l,h}$ and this classification is almost correct, i.e. $|h-{h^{\prime}_{x}}|\leqslant 1$. We include data from adjacent slices and perform the linear projection of the samples in $\bigcup_{|h-{\widehat{h}_{x}}^{\prime}|\leqslant 1}\widehat{S}_{l,h}$ onto the one-dimensional line with direction ${\widehat{v}_{l,h}}$. Let interval $I^{(l,h)}$ denote the range of the projected data $\{\langle{\widehat{v}_{l,h}^{b}},x\rangle:x\in\bigcup_{|h-{h^{\prime}_{x}}|\leqslant 1}\widehat{S}_{l,h}\}$. We use a piecewise polynomial estimator to learn the link function $f$ in a local neighborhood. Consider the mean-squared error of the estimated function ${\widehat{F}}(x)={\widehat{f}_{j|{\widehat{v}_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)$ vs. the true function $F$. There are several sources for such error: the curve approximation error between the nonlinear projection $\Pi_{\gamma}(\cdot)$ and the local linear projection $\langle{v_{l,h}^{b}},\cdot\rangle$, the estimation error for ${v_{l,h}^{b}}$, and the bias and variance in the nonparametric estimator ${\widehat{f}_{j|{v_{l,h}^{b}}}}$. Let $\widetilde{f}$ denote the nearest approximation of $F(x)=f(\Pi_{\gamma}x)$ onto the space of one-dimensional functions dependent only $\langle{v_{l,h}^{b}},X\rangle)$:
    $\widetilde{f}:=\operatorname*{{arg\,min}}_{g:\mathbb{R}^{1}\to\mathbb{R}^{1}}\mathbb{E}\left[\left|Y-g(\langle{v_{l,h}^{b}},X\rangle)\right|^{2}\mathbbm{1}_{I^{(l,h)}}(\langle{v_{l,h}^{b}},X\rangle)\ \big{|}\ X\in{S_{l,h}}\right]\,.$
    One can show that $\widetilde{f}$ has the following explicit formula
    $\widetilde{f}(s)=\mathbb{E}\left[F(X)\ \big{|}\ X\in{S_{l,h}},\ \langle{v_{l,h}^{b}},X\rangle=s\right]$
    and hence $\widetilde{f}$ is Hölder continuous with semi-norm $\left[\widetilde{f}\right]_{\mathcal{C}^{s}}\leqslant\left[f\right]_{\mathcal{C}^{s}}+\frac{C_{Y}R_{0}}{l}|\rho_{X}|_{\mathcal{C}^{s}}$.
    We partition the interval $I^{(l,h)}$ uniformly into $j$ intervals ${I^{(l,h)}_{j,k}}$, $k=1,\dots,j$. On each ${I^{(l,h)}_{j,k}}$, we let ${f_{j,k|v}}$ be the $m$-order polynomial population estimator of the function $\widetilde{f}$, conditioned on projecting the data of slice ${S_{l,h}}$ onto the unit vector $v\in\mathbb{S}^{d-1}$. Joining the ${f_{j,k|v}}$’s together over $k=1,\dots,j$, we obtain ${f_{j|v}}$ as a population estimator of $\widetilde{f}$ conditioned on projecting the data of slice ${S_{l,h}}$ onto the unit vector $v\in\mathbb{S}^{d-1}$. This ${f_{j|v}}$ is a piecewise $m$-order polynomial with $j$ pieces:
    ${f_{j,k|v}}$ $=\operatorname*{{arg\,min}}_{p:\mathbb{R}^{1}\to\mathbb{R}^{1},\deg(p)\leqslant m}\mathbb{E}\left[\left|Y-p(\langle v,X\rangle)\right|^{2}\mathbbm{1}_{{I^{(l,h)}_{j,k}}}(\langle v,X\rangle)\ \big{|}\ X\in{S_{l,h}}\right]\quad,\quad$
    ${f_{j|v}}(s)$ $=\sum_{k=1}^{j}{f_{j,k|v}}(s)\mathbbm{1}_{I_{j,k}}(s)\,,$
    Conditioned on the event of almost correct classification $|{\widehat{h}_{x}}-{h_{x}}|\leqslant 1$, we decompose the estimation error $|F-{\widehat{f}_{j|{\widehat{v}_{l,h_{x}}^{b}}}}|$ into the following terms, which we call respectively, nonlinear curve approximation error, direction error, bias, variance, and projection error:
    $\left|F(x)-{\widehat{f}_{j|{\widehat{v}_{l,h_{x}}^{b}}}}\right|$ $\leqslant\underbrace{\left|f(\Pi_{\gamma}x)-\widetilde{f}(\langle{v_{l,h_{x}}^{b}},x\rangle)\right|}_{(NCA)}+\underbrace{\left|\widetilde{f}(\langle{v_{l,h_{x}}^{b}},x\rangle)-\widetilde{f}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|}_{(\Phi)}$
    $+\underbrace{\left|\widetilde{f}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{f_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|}_{(B)}+\underbrace{\left|{f_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{\widehat{f}_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|}_{(V)}$
    $+\underbrace{\left|{\widehat{f}_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{\widehat{f}_{j|{\widehat{v}_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|}_{(\Psi)}$
    Proposition 12 (Mean Squared Error conditioned on almost correct classification)
    Consider assumptions (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), ($\bm{\omega_{f}}$), and (P). Suppose $f\in\mathcal{C}^{s}$ with $s\in\left[\frac{1}{2},2\right]$. Conditioned on almost correct classification in part 2. b) in Algorithm 1, then we have the following estimates for $n$ is sufficiently large,
    $\operatorname{\mathrm{MSE}}_{(NCA)}\lesssim$ $\left[f\right]_{\mathcal{C}^{{s\wedge 1}}}^{2}(\sigma_{\gamma}\left\|\gamma^{\prime\prime}\right\|C_{f})^{{2(s\wedge 1)}}\max\left(\sigma_{\zeta},\omega_{f},\frac{\mathrm{len}_{\gamma}}{C_{f}^{\prime}l}\right)^{{2(s\wedge 1)}}$
    $\operatorname{\mathrm{MSE}}_{(\Phi)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{{s\wedge 1}}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{{s\wedge 1}}})^{2}(\mathrm{len}_{\gamma}R_{0}^{2}\sigma_{\gamma}^{-2})^{{2(s\wedge 1)}}(d^{2}\log d)^{{s\wedge 1}}\left(\frac{\log n\log l}{nl}\right)^{{2(s\wedge 1)}}$
    $\operatorname{\mathrm{MSE}}_{(B)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})^{2}C_{f}^{2s}\max\left(\sigma_{\zeta},\omega_{f},\frac{\mathrm{len}_{\gamma}}{C_{f}^{\prime}l}\right)^{2s}j^{-2s}$
    $\operatorname{\mathrm{MSE}}_{(V)}\lesssim$ $\sigma_{\zeta}^{2}\frac{lj\log j}{n}$
    $\operatorname{\mathrm{MSE}}_{(\Psi)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}\left(\mathrm{reach}_{\gamma}^{2}\frac{C_{f}^{2}C_{Y}^{2}R_{0}^{4}}{\sigma_{\gamma}^{4}}d\log d\frac{\log n\log l}{nl}\right)^{\frac{1}{2-(s\wedge 1)}}$
    $+\epsilon_{0}^{2}+j\exp\left(-\frac{c\epsilon_{0}^{2}n(lj)^{-1}}{(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}\mathrm{reach}_{\gamma}^{2(s\wedge 1)}}\right)$
    $+C_{Y}^{2}R_{0}^{2}d\exp\left(-\frac{c\sigma_{\gamma}^{4}nl^{-1}j^{-2}(\tau\log n)^{-\frac{1}{2}}}{C_{Y}^{2}C_{f}^{2}R_{0}^{4}d(l^{-2}+l^{-1}j^{-1})}\right)+C_{Y}^{2}R_{0}^{2}d\exp\left(-\frac{c\sigma_{\gamma}^{4}nl^{-1}}{R_{0}^{4}d^{2}}\right)+C_{Y}^{2}R_{0}^{2}n^{-\tau}$
    where free variable $\epsilon_{0}$ can be taken arbitrarily.
    Putting together these bounds and optimizing, yields our main Theorems.
    4 Numerical Experiments
    We test the performance of Algorithm 1 on synthetic data to demonstrate its performance and scalability, consistently with the main Theorems. We let the number of samples $n$ increases from $10^{4}$ to $10^{6}$. For each $n$, we randomly pick $n$ points from the underlying curve, and we use 90% for constructing the estimator and 10% for testing in Algorithm 1. The algorithm requires two key scale parameters, and we will use the values $l^{*}$ and $j^{*}$ dictated by the main Theorems. We want to study the mean squared error $\mathbb{E}[{\widehat{F}}_{n}(X)-F(X)|^{2}]$, the estimation error of the center along the tangential direction $\mathbb{E}[|\langle{\hat{\mu}}_{l^{*},h}-\mu_{l^{*},h},\gamma^{\prime}_{l^{*},h}\rangle|]$, and the difference between the significant vector and the tangential direction $\mathbb{E}[||{\widehat{v}}_{l^{*},h}-\gamma^{\prime}_{l^{*},h}||]$. For each $n$, we run the numerical estimation with five independent repetitions.
    In each example in this section, we randomly generate $n_{\max}:=2\times 10^{6}$ points from the underlying curve $\gamma$, for which we will have a complete parametrization. We use all these $2\times 10^{6}$ data to compute an approximation to the center ${\mu_{l,h}}:=\mathbb{E}\left[X\mid Y\in{R_{l,h}}\right]$ and the average tangential vector $\gamma^{\prime}_{l,h}:=\frac{\mathbb{E}\left[\gamma^{\prime}(\Pi_{\gamma}X)\mid Y\in{R_{l,h}}\right]}{\left\|\mathbb{E}\left[\gamma^{\prime}(\Pi_{\gamma}X)\mid Y\in{R_{l,h}}\right]\right\|}$ on each slice. Here, the unit vector $\gamma^{\prime}_{l,h}$ is parallel to the average tangential direction $\mathbb{E}\left[\gamma^{\prime}(\Pi_{\gamma}X)\ \mid\ X\in{S_{l,h}}\right]$ on the slice ${S_{l,h}}$.
    To obtain a good estimation of the nonlinear curve approximation error $\operatorname{\mathrm{MSE}}_{(NCA)}$, which we indicate as responsible for the additive term that does not go to 00 as $n$ increases in the bound in Theorem 3, we replace the estimated parameters $\{({\widehat{\mu}_{l,h}},{\widehat{v}_{l,h}})\}_{h=1}^{l}$ by the “oracle” parameters $\{({\mu_{l,h}},\gamma^{\prime}_{l,h})\}_{h=1}^{l}$ (computed on the $n_{\max}$ points as described above) when performing the local linear projection and the local polynomial regression on each sample slice. We choose $(l,j)$ to obtain the minimum value of the MSE when the number of samples is $n_{\max}$, denoted by “MSE at $n=2\times 10^{6}$” in Fig.s 4,5,6,7,8. In this way we aim at reducing the effect of any errors originating from the estimation of the parameters $({\widehat{\mu}_{l,h}},{\widehat{v}_{l,h}})_{h=1}^{l}$ and from the particular choice $l=l^{*}$ and $j=j^{*}$, thereby imputing the reported “MSE at $n=2\times 10^{6}$” mainly to the nonlinear curve approximation error $\operatorname{\mathrm{MSE}}_{(NCA)}$.
    For all figures in this section, we use loglog plots to study how the following quantities decay with the number of samples $n$: the mean squared error $\mathbb{E}[|{\widehat{F}}_{n}(X)-F(X)|^{2}]$, the estimation error of the center along the tangential direction $\mathbb{E}[|\langle{\hat{\mu}}_{l^{*},h}-\mu_{l^{*},h},\gamma^{\prime}_{l^{*},h}\rangle|]$, and the difference between the significant vector and the tangential direction $\mathbb{E}[||{\widehat{v}}_{l^{*},h}-\gamma^{\prime}_{l^{*},h}||]$. On the interval $n\in[10^{5},10^{6}]$, we use least squares linear regression to estimate the decaying rate. We add a dashed vertical line $n=10^{5}$ for those figures to emphasize that the learning rate only corresponds to $n\in[10^{5},10^{6}]$. Because the mean squared error has a term due to the curve approximation error, one should keep in mind that the learning rate for mean squared error in numerical tests here still only indicative of the learning rate for $n\in[10^{5},10^{6}]$ and may be smaller than the min-max optimal exponent $\frac{2s}{2s+1}$ when the mean squared error is dominated by the curve approximation error and the overall mean squared error starts to get saturated, which happens for large $n$. To complicate things further, such curve approximation error depends on curvature, which in turn can depend on the dimension.
    4.1 Example 1: Circular Arcs and verifying Theorem 3
    We consider the simplest nonlinear curves as arcs of circles because they have constant curvatures and can be linearly embedded in $\mathbb{R}^{2}$. We consider a collection of curves where each is an arc of a circle. These curves are embedded in a fixed ambient dimension 20, with a fixed length $\mathrm{len}_{\gamma}=1$. Again, we fix $\sigma_{\gamma}=0.5$ and the random vector $Z_{d-1}$ follows the normal distribution $\mathcal{N}(0,\sigma_{\gamma}^{2}\mathbb{I}_{d-1})$ with truncation at $\left\|Z_{d-1}\right\|<0.9\,\mathrm{reach}_{\gamma}$. The curvature of these curves is varied across realizations, ranging from $0.04$ to $0.4$ We set our upper bound for the curvature to be 0.4 because otherwise $\mathrm{reach}_{\gamma}$ becomes too small and the variance of $Z_{d-1}$ will no longer be approximately $\sigma_{\gamma}^{2}$.
    Figure 4: Numerical Tests in Section 4.1: Circular arcs embedded in $\mathbb{R}^{d}$, $d=20$, with unit length and curvature varying in $[0.04,0.4]$. We fix the noise level $\sigma_{\zeta}=0.03$. Upper row: MSE for ${\widehat{F}}$ (left) and MSE at $n=2\times 10^{6}$ as a function of curvature (right); Bottom row: estimation error for the center along tangential direction (left) and difference between estimated significant vector and the tangential direction (right) over $5$ runs.
    Theorem 3 claims that the curve approximation error is proportional to $\left\|\gamma^{\prime\prime}\right\|^{{2(s\wedge 1)}}$, which is smaller for curve with smaller curvature: we verify that claim by the following numerical test. We choose $f(t)=\exp(t)$ for $t\in[0,1]$, which has smoothness exponent $s\geq 2$. The external noise follows the normal distribution $\mathcal{N}(0,\sigma_{\zeta}^{2})$ with $\sigma_{\zeta}=0.03$. Figure 4 supports the following conclusions, consistent with Theorem 3 and our analysis: (i) when $\sigma_{\zeta}>0$, then the mean squared error has a nonzero lower bound independent of $n$; (ii) the curve approximation error is proportional to $\left\|\gamma^{\prime\prime}\right\|^{{2(s\wedge 1)}}$; (iii) Corollary 9 states that the decaying rate is $\mathcal{O}(n^{-\frac{1}{2}})$ for $\mathbb{E}[|\langle{\hat{\mu}}_{l^{*},h}-\mu_{l^{*},h},\gamma^{\prime}_{l^{*},h}\rangle|]$ and $\mathbb{E}[||{\hat{\mu}}_{l^{*},h}-\gamma^{\prime}_{l^{*},h}||]$.
    4.2 Meyer helix
    We perform numerical tests for a collection of curves called “Meyer helix” in $\mathbb{R}^{d}$, which we construct so that the curve complexity grows with the ambient dimension $d$. This is inspired by a curve called the Meyer staircase (named after Y. Meyer), defined by the map $[0,1]\rightarrow L^{p}(\mathbb{R})$, for some $p\geq 1$, given by $t\mapsto\mathbf{1}_{[0,\frac{1}{2}]}(\cdot-t)$. We smooth out this original example by considering translations of a Gaussian, and we induce further twists in the curve to increase its complexity by introducing the Meyer helix as a variation. We measure the growing complexity of this family of curves as a function of $d$, and show that $\mathrm{len}_{\gamma}\asymp d^{1.5},\mathrm{diam}_{\gamma}\asymp d^{0.5},|\gamma^{\prime\prime}|\asymp d^{-0.5},\mathrm{reach}_{\gamma}\asymp d^{0.5}$, and effective linear dimension $\asymp d^{1}$: see details in Appendix C. This collection of curves allows us to verify the performance of our algorithm when varying $d$: we fix $\sigma_{\gamma}=0.5$ and let the random vector $Z_{d-1}$ follow the normal distribution $\mathcal{N}(0,\sigma_{\gamma}^{2}\mathbb{I}_{d-1})$ with truncation at $|Z_{d-1}|<0.9\,\sqrt{d}$.
    4.2.1 Example: Verifying Theorem 3 and Corollary 9
    We let the underlying curve $\gamma$ to be the Meyer helix in $d=7$ dimensions, which has $\mathrm{len}_{\gamma}=53.78$ and $\mathrm{reach}=2.65$. Consider the link function $f(t)=\mathrm{len}_{\gamma}\cdot\exp(t/\mathrm{len}_{\gamma})$ for $t\in[0,\mathrm{len}_{\gamma}]$ which has smoothness exponent $s=2$. The observational noise $\zeta$ follows the normal distribution $\mathcal{N}(0,\sigma_{\zeta}^{2})$ where the noise level $\sigma_{\zeta}$ varies from 0.05 to 0.2.
    Figure 5: Numerical Tests in Section 4.2.1: Meyer helix in $d=7$ dimensions, with $f$ of smoothness exponent $s=2$, and noise level $\sigma_{\zeta}$ varying in $[0.05,0.2]$. Top row: MSE for ${\widehat{F}}$ (left) and MSE at $n=2\times 10^{6}$ as a function of $\sigma_{\zeta}$ (right); Bottom row: estimation error of center along tangential direction (left) and difference between estimated significant vector and tangential direction (right), over five independent runs.
    Figure 5 supports the following theoretical conclusions: (i) When $\sigma_{\zeta}>0$, then the mean squared error has a nonzero lower bound. (ii) The curve approximation error is proportional to $\sigma_{\zeta}^{{2(s\wedge 1)}}$. (iii) consistently with Corollary 9, the learning rate is $\mathcal{O}(n^{-\frac{1}{2}})$ for $\mathbb{E}[|\langle{\hat{\mu}}_{l^{*},h}-\mu_{l^{*},h},\gamma^{\prime}_{l^{*},h}\rangle|]$ and $\mathbb{E}[||{\hat{\mu}}_{l^{*},h}-\gamma^{\prime}_{l^{*},h}||]$.
    4.2.2 Example: Verifying Theorem 3 and Corollary 9
    We consider a collection of Meyer helix curves with the following ambient dimensions and geometric parameters:
    $d$ 3 4 5 6 7 8 9
    $\mathrm{len}_{\gamma}$ 20.86 33.39 35.35 43.89 53.78 90.20 96.65
    $\mathrm{reach}_{\gamma}$ 1.73 2.00 2.24 2.45 2.65 2.83 3.00
    Consider the link function $f(t)=\mathrm{len}_{\gamma}\cdot\exp(t/\mathrm{len}_{\gamma})$ for $t\in[0,\mathrm{len}_{\gamma}]$ which has smoothness exponent $s=2$. The observational noise $\zeta$ follows the normal distribution $\mathcal{N}(0,\sigma_{\zeta}^{2})$ where the magnitude of the noise is $\sigma_{\zeta}=0.2$.
    Figure 6: Numerical Tests in Section 4.2.2: Meyer helix curves in $\mathbb{R}^{d}$, $d\in\{3,\dots,9\}$, $f$ with smoothness exponent $s=2$, and $\sigma_{\zeta}=0.2$. Left: Mean Squared Error for ${\widehat{F}}$. Middle: $L^{2}$-relative Mean Squared Error, $\mathbb{E}[|{\widehat{F}}-F|^{2}]/\mathbb{E}[|F(X)-\mathbb{E}(F(X))|^{2}]$, computed over five independent runs. Right: Mean Squared Error of ${\widehat{F}}$ at $n=2\times 10^{6}$ as a function of $d$.
    In the upper-left plot of Figure 6, we observe that, for fixed $n=10^{5}$, the mean squared error is larger for Meyer helix curves with in ambient dimension $d$. This is because the coefficient before $n^{-\frac{2s}{2s+1}}$ in Theorem 3 is larger for curves with larger $\mathrm{len}_{\gamma}$. In upper right plot, we observe that, by contrast, the mean squared error at $n=2\times 10^{6}$ is smaller for Meyer helix curves in larger ambient dimension $d$. This verifies the statement in Theorem 3 that the curve approximation error is proportional to $\mathrm{reach}_{\gamma}^{-2{s\wedge 1}}$. Some observations about this example: the Meyer helix in smaller $d$ has smaller $\mathrm{len}_{\gamma},\mathrm{reach}_{\gamma}$, and larger curvature, and therefore: (i) the requirement for the number of samples for Theorem 3 is smaller; (ii) the first term in the mean squared error, which is $\mathcal{O}(n^{-\frac{2s}{2s+1}}\log n)$, has a smaller coefficient; (iii) the second term in the bound for the mean squared error in Theorem 3, which is the curve approximation error, has larger magnitude; (iv) if we denote by $n_{1}$ the number of samples such that first term balances the second term in the upper bound for the mean squared error, $n_{1}$ increases with $d$. As a consequence, on the particular interval $n=[10^{5},10^{6}]$ where the decaying rate is calculated, one need to pay attention to the value of $n_{1}$ that determines which term dominates the mean squared error. In particular, the Meyer helix in $d=3$ dimensions has $n_{1}\approxeq 10^{5}$, with the MSE exhibiting good decay for $n\leq n_{1}$, and saturating for $n\geq n_{1}$. In contrast, the Meyer helix in $d=7$ deimsnsions has $n_{1}\approxeq 10^{6}$, and thus we observe a good decay rate on the interval $n\in[10^{5},10^{6}]$. For the Meyer helix in higher dimension, the requirement for the number of samples by Theorem 3 is even larger. There is another phenomenon appearing in this numerical test: we notice that for $d=8,9$, the error increases when $n$ increase from $10^{4}$ to $2\times 10^{4}$ and decreases when $n\geqslant 4\times 10^{4}$. First, this is not a contradiction with Theorem 3 because the requirement for the minimal number of samples is not satisfied for the Meyer helix in dimension $d=8,9$ when $n\leqslant 3\times 10^{4}$. Second, this increase of error for small $n$ is due to the transition from the “wide” slice scenario to the “thin” slice scenario. Further investigation on the average empirical geometric quantity $\frac{1}{l^{*}}\sum_{h=1}^{l^{*}}\widehat{H}_{l^{*},h}$ shows that we have the “wide” slice scenario when $n\leqslant 3\times 10^{3}$ and have the “thin” slice scenario when $n\geqslant 4\times 10^{4}$. For $n\in[3\times 10^{4},4\times 10^{4}]$, the average empirical geometric quantity $\frac{1}{l^{*}}\sum_{h=1}^{l^{*}}\widehat{H}_{l^{*},h}\approx 0$ and thus the slices are roughly isotropic which, as discussed in Section 2.1, prevents an accurate estimate of the significant vector.
    Figure 6 supports the following conclusions, in line with our theoretical analysis: (i) the constants in $\mathcal{O}(n^{-2s/(2s+1)})$ are bigger for curves with bigger $\mathrm{len}_{\gamma}$; (ii) the requirement for the number of samples so that Theorem 3 holds is larger for more complex curves; (iii) the significant vector cannot be estimated well when the geometric shape of a slice is roughly isotropic; (iv) the value of $n_{1}$ is larger for curves with larger length and larger reach, where $n_{1}$ is the number of samples such that the first term balances the second term in the upper bound for the MSE; (v) the curve approximation error is proportional to $\mathrm{reach}_{\gamma}^{-{2(s\wedge 1)}}$;
    Figure 7: Numerical Tests in Section 4.2.3: the collection “Meyer helix” of curves with varying dimension. Left: $d=5,\dots,10$. Right: $d=6,12,24,36,48$. For all tests, we choose zero observational noise $\zeta\equiv 0$, and the link function is Lipschitz (i.e., smoothness exponent $s=1$). These errors are computed as the average values in five independent repetitions.
    4.2.3 Example: Verifying Theorem 4
    We consider a collection of Meyer helix curves in the following ambient dimensions and information:
    $d$ 5 6 7 8 9 10 12 24 36 48
    $\mathrm{len}_{\gamma}$ 35.35 43.89 53.78 90.20 96.65 93.88 135.76 435.48 730.62 1306.78
    $\mathrm{reach}_{\gamma}$ 2.24 2.45 2.65 2.83 3.00 3.16 3.46 4.90 6.00 6.93
    Note that for Meyer helix curves with $d=5,\dots,12$, the number of data is sufficient when $n\geq 10^{4}$, while for $d=24,36,48$, the requirement for the number of data is substantially larger. Figure 7 supports the following theoretical conclusions: (i) when $\sigma_{\zeta}=0$, then the mean squared error has no lower bound. The curve approximation error here is zero; (ii) the decaying rate is $\mathcal{O}(n^{-2})$ if the link function $f$ is monotone and Lipschitz; (iii) The requirement for the number of data is larger for more complex curves.
    5 Robustness: Performance of Algorithm 1 in general setting
    Recall that Assumption (LCV) gives a quantitative requirement for the “thin” slice scenario. If we relax this assumption and consider instead the “wide” slice scenario, we expect that the largest principal component on each slice gives a proper approximation of the tangential direction under some assumption. Small curvature will do:
    1. (SC)
      Define $\sigma_{\gamma}:=\min_{t_{0}\in[0,\mathrm{len}_{\gamma}]}\sigma_{\gamma}(t_{0})$ as the minimum value of $\sigma_{\gamma}(t_{0})$. We assume that there exist $c_{1},c_{2}>0$ such that $\sigma_{\gamma}<c_{1}C_{f}\max(\sigma_{\zeta},\omega_{f})$ and $C_{f}\max(\sigma_{\zeta},\omega_{f})\leqslant c_{2}\mathrm{reach}_{\gamma}$.
    Proposition 13 ((SC) implies “wide” slices)
    Suppose (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), and ($\bm{\omega_{f}}$) hold true, together with (SC), with $c_{1},c_{2}$ smaller than a small-enough universal constant. Then, for every $l$ such that $|{R_{l,h}}|\asymp\max(\sigma_{\zeta},\omega_{f}),h\in\mathcal{H}_{l}$, we have
    $\lambda_{1}({\Sigma_{l,h}^{b}})-\lambda_{2}({\Sigma_{l,h}^{b}})\gtrsim C_{f}^{2}|{R_{l,h}}|^{2}$
    Recall that we have the following distance function in this situation of “wide” slices, i.e. $H_{l,h}<0$, ${\mathrm{dist}}(x,h)=\left\|x-{\mu_{l,h}^{b}}\right\|^{2}+\frac{{\lambda_{d}(\Sigma_{l,h}^{b})}}{{\lambda_{1}(\Sigma_{l,h}^{b})}}|\langle x-{\mu_{l,h}^{b}},{v_{l,h}^{b}}\rangle|^{2}$, and similarly for its empirical counterpart. With a proof as in Proposition 11, we conclude that
    Proposition 14 (classification accuracy without (LCV))
    Assume that (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), ($\bm{\omega_{f}}$), and (SC) hold true. Fix $l=\frac{C_{Y}R_{0}}{\max(\sigma,\omega_{f})}$. Then, the probability of misclassification by at least two slices in part 2. b) of Algorithm 1 can be bounded by
    $\mathbb{P}\left(\left|{\widehat{h}_{x}}-{h_{x}}\right|\geqslant 2\right)\lesssim ld\exp\left(-c\frac{C_{f}^{6}\max(\sigma_{\zeta},\omega_{f})^{7}n}{C_{Y}R_{0}^{7}d^{3}\sqrt{\log n}}\right)+ln^{-\tau}$
    5.1 Example: Verifying Theorem 5
    Recall that in the “wide” slice scenario, we have a nontrivial error that does not vanish as $n$ grows and is at the magnitude of the curve approximation error. We will verify this in the following numerical test. Let the underlying curve $\gamma$ to be the Meyer helix in $d=21$ dimension, which has $\mathrm{len}_{\gamma}=370.63$ and $\mathrm{reach}_{\gamma}=4.58$. The link function $f(t)=\mathrm{len}_{\gamma}\exp(t/\mathrm{len}_{\gamma})$ for $t\in[0,\mathrm{len}_{\gamma}]$ has smoothness exponent $s=2$. The external noise $\zeta$ follows the normal distribution $\mathcal{N}(0,\sigma_{\zeta}^{2})$ where the noise level varies from 2 to 6. Note that assumption (LCV) is not satisfied in this case.
    Figure 8: Numerical Tests in Section 5.1: Meyer helix in $d=21$ dimensions. The link function has smoothness exponent $s=2$, and the noise level $\sigma_{\zeta}$ varies from 2 to 6. Left: mean squared error; Right: mean squared error at $n=2\times 10^{6}$; These errors are computed as the average values in five independent repetitions.
    Figure 8 verifies that when the assumption (LCV) is not satisfied, then the mean squared error saturates at the level of the curve approximation error.
    6 Conclusion
    We introduced the Nonlinear Single-Variable Model, which is a compositional model $F=f\circ g$ for functions in high-dimensions where both $f$ and $g$ are nonlinear, but $g$ has a one-dimensional range, and $f$ is therefore a function of only one dimension. Thanks for the geometric structure inherent in $g$, using techniques based on inverse regression, at least when $f$ is roughly monotone, we are able to efficiently estimate the level sets of $g$ and then learn $f$, both in a nonparametric fashion, with learning rates and sample requirements not cursed by the ambient dimension, and with computationally efficient algorithms for constructing the estimators that scale linearly in the number of samples and with constants moderately depending on the ambient dimension.
    Future directions include the extension to functions $f$ that are not roughly monotone presents challenges for inverse regression method, but it is also so for other techniques, for example for stochastic gradient methods in the single- and multi-index models (Arous et al., 2021; Bietti et al., 2023), and there are still gaps in the understanding of the sharp statistical and computational tradeoffs. Another extension of interest is to a Nonlinear Multi-Variable Model, where the curve $\gamma$ is replaced by a higher-dimensional manifold $\mathcal{M}$, also presents challenges, as the geometry of the level sets of $F$ becomes significantly more complicated. Both these extensions are subjects of current investigation and left to future work.
    Understanding how compositional structure affects the design of estimators, or specific “general purpose” algorithms for constructing them (such as SGD applied to a suitable loss function), especially in the case of multiple compositions, beyond just two, and when and how such structure can help avoid the curse of dimensionality, is an interesting area of research with many open questions.
    Acknowledgments and Disclosure of Funding
    The authors are grateful to A. Lanteri, S. Vigogna and T. Klock (during his visit to Johns Hopkins) for early discussions and experiments on this model, and to Fei Lu and Xiong Wang and their feedback on an early version of this manuscript. This research was partially supported by AFOSR FA9550-21-1-0317 and FA9550-23-1-0445. Prisma Analytics Inc. provided and managed the computing resources for the numerical experiments.
    Appendix A Proof of Propositions
    Lemma 15 (Matrix Bernstein Inequality (Vershynin, 2018, Theorem 5.4.1) )
    Let
    $M_{1}$$\dots$, $M_{n}$ be a sequence of i.i.d. $d_{1}\times d_{2}$ dimensional random matrices with $\mathbb{E}M_{i}=0,\left\|M_{i}\right\|\leqslant B$ (bounded operator norm). Denote sample mean $\widehat{M}=\frac{1}{n}\sum_{i\leqslant n}M_{i}$ and denote covariance norm $\sigma^{2}=\max\left(\left\|\mathbb{E}[X_{i}X_{i}^{\intercal}]\right\|,\left\|\mathbb{E}[X_{i}^{\intercal}X_{i}]\right\|\right)$. Then
    $\mathbb{P}\left(\left\|\widehat{M}\right\|\gtrsim t\right)\lesssim(d_{1}+d_{2})\exp\left(-\frac{cnt^{2}}{Bt+\sigma^{2}}\right)$
    Lemma 16 (Concentration Inequalities for means, covariances and eigenvalues)
    Suppose that $X_{i}$ are iid bounded by $\left\|X_{i}\right\|\leqslant R_{0}\sqrt{d}$, let ${\mu_{l,h}^{b}}=\mathbb{E}[X]$ and ${\widehat{\mu}_{l,h}^{b}}=\frac{1}{{n_{l,h}^{b}}}\sum_{i\leqslant{n_{l,h}^{b}}}X_{i}$ denote the mean and sample mean. Let ${\Sigma_{l,h}^{b}}=\mathbb{E}[(X-{\mu_{l,h}^{b}})(X-{\mu_{l,h}^{b}})^{\intercal}]$ denote the covariance matrix, ${\widehat{\Sigma}_{l,h}^{b}}=\frac{1}{{n_{l,h}^{b}}}\sum_{i\leqslant{n_{l,h}^{b}}}(X_{i}-{\widehat{\mu}_{l,h}^{b}})(X_{i}-{\widehat{\mu}_{l,h}^{b}})^{\intercal}$ the sample covariance matrix, and ${\widetilde{\Sigma}_{l,h}^{b}}=\frac{1}{{n_{l,h}^{b}}}\sum_{i\leqslant{n_{l,h}^{b}}}(X_{i}-{\mu_{l,h}^{b}})(X_{i}-{\mu_{l,h}^{b}})^{\intercal}$ be the augmented covariance matrix. Then,
    $\mathbb{P}\left(\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|\gtrsim R_{0}\sqrt{\frac{\delta}{\alpha}}\right)\lesssim d\exp\left(-\frac{c{n_{l,h}^{b}}\delta}{\alpha d}\right),\ \ \text{ where }\delta<\alpha d$
    $\mathbb{P}\left(\left\|{\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|\gtrsim\beta\frac{R_{0}^{2}}{\alpha}\right)\lesssim d\exp\left(-\frac{c\beta^{2}{n_{l,h}^{b}}}{\alpha^{2}d^{2}}\right),\ \ \text{ where }\beta<\alpha d$
    $\mathbb{P}\left(\left|{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{1}(\Sigma_{l,h}^{b})}\right|\gtrsim\beta\frac{R_{0}^{2}}{\alpha}\right)\lesssim d\exp\left(-\frac{c\beta^{2}{n_{l,h}^{b}}}{\alpha^{2}d^{2}}\right),\ \ \text{ where }\beta<\alpha d$
    Proof  The first two inequalities can be shown directly by Lemma 15. We can split ${\widehat{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}$ into ${\widehat{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}={\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}+({\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}})({\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}})^{\intercal}$. To bound deviation in ${\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}$, one can directly use Weyl’s inequality
    $\left|{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{1}(\Sigma_{l,h}^{b})}\right|\leqslant\left\|{\widehat{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|\leqslant\left\|{\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|+\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|^{2}\,.$
    The third inequality follows from the first two inequalities.
    A.1 Proof of Proposition 6
    Proof  Define intervals $I_{k}:=(-\sqrt{2(k+1)}\sigma_{\zeta},\sqrt{2k}\sigma_{\zeta}]\cup[\sqrt{2k}\sigma_{\zeta},\sqrt{2(k+1)}\sigma_{\zeta})$ for $k=0,1,2,\dots$. We first note that, thanks to ($\bm{\zeta}$ subG), we have $\zeta_{i}\in\bigcup_{k\leqslant\tau\log n}I_{k}$ with probability higher than $1-2n^{-\tau}$, for every $i=1,\dots,n$. Conditioned on this event and on $Y_{i}\in T$, $\Pi_{\gamma}X_{i}\in f^{-1}(T+\bigcup_{k\leq\tau\log n}I_{k})$. Meanwhile, $\mathbb{E}\left[\Pi_{\gamma}X|Y\in T,\mathcal{B},\zeta\in I_{k}\right]\in[\min f^{-1}(T+I_{k}),\max f^{-1}(T+I_{k})]$. It follows from assumption ($\bm{\omega_{f}}$) that we have an absolute bound upon conditioning on $Y_{i}\in T,\mathcal{B},\zeta_{i}\in\bigcup_{k\leq\tau\log n}I_{k}$:
    $\left|\Pi_{\gamma}X_{i}-\mathbb{E}\left[\Pi_{\gamma}X|Y\in T,\mathcal{B},\zeta\in I_{k}\right]\right|\leqslant C_{f}(|T|+\sqrt{\max(k,\tau\log n)}\sigma_{\zeta})\,.$
    By the law of total expectation,
    $\left|\Pi_{\gamma}X_{i}-\mathbb{E}\left[\Pi_{\gamma}X|Y\in T,\mathcal{B}\right]\right|\leqslant\sum_{k=0}^{\infty}\left|\Pi_{\gamma}X_{i}-\mathbb{E}\left[\Pi_{\gamma}X|Y\in T,\mathcal{B},\zeta\in I_{k}\right]\right|\mathbb{P}(\zeta\in I_{k})$
    $\lesssim$ $C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta}+\sigma_{\zeta}\sum_{k>\tau\log n}\sqrt{k}e^{-k})\lesssim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})$
    which finishes the proof of (a). For (b), we use the law of total expectation to write
    $\mathrm{Var}\left[\Pi_{\gamma}X|Y\in T,\mathcal{B}\right]$ $=\mathbb{E}\left[\left(\Pi_{\gamma}X-\mathbb{E}\left[\Pi_{\gamma}X|Y\in T,\mathcal{B}\right]\right)^{2}\mid Y\in T,\mathcal{B}\right]$
    $=\sum_{k=0}^{\infty}\mathbb{E}\left[\left(\Pi_{\gamma}X-\mathbb{E}\left[\Pi_{\gamma}X|Y\in T,\mathcal{B}\right]\right)^{2}\mid Y\in T,\mathcal{B},\zeta\in I_{k}\right]\mathbb{P}\left\{\zeta\in I_{k}\right\}\,.$
    To bound each term, we follow the same approach as in part (a) to bound conditional random variable $\Pi_{\gamma}X\mid Y\in T,\mathcal{B},\zeta\in I_{k}$
    $\left|\Pi_{\gamma}X\mid Y\in T,\mathcal{B},\zeta\in I_{k}-\mathbb{E}\left[\Pi_{\gamma}X|Y\in T,\mathcal{B}\right]\right|\lesssim$ $C_{f}(|T|+\sqrt{k}\sigma_{\zeta})\,,$
    whence
    $\mathrm{Var}\left[\Pi_{\gamma}X|Y\in T,\mathcal{B}\right]\lesssim C_{f}^{2}\bigg{(}|T|^{2}+\sigma_{\zeta}^{2}\sum_{t=0}^{\infty}te^{-t}\bigg{)}\lesssim C_{f}^{2}(|T|^{2}+\sigma_{\zeta}^{2})\,.$
    A.2 Proof of Proposition 7 and Corollary 8
    Proof  Recall that that $X_{i}=\gamma(t_{i})+W_{i}$ where $t_{i}=\Pi_{\gamma}X_{i}$ denotes position of $X_{i}$ along curve and $W_{i}=M_{\gamma^{\prime}(t_{i})}\begin{pmatrix}W_{i}^{\prime}\\ 0\end{pmatrix}$ denotes the deviation off the curve. Here each $M_{v}\in\mathbb{O}(d)$ is a rotation matrix on $\mathbb{R}^{d}$ that maps $d$-th canonical unit vector $\hat{e}_{d}=(0,\dots,0,1)$ to the unit vector $v\in\mathbb{S}^{d-1}$. Observe that $\langle{v_{l,h}^{b}},X_{i}\rangle=\langle{v_{l,h}^{b}},\gamma(\Pi_{\gamma}X_{i})\rangle+\langle{v_{l,h}^{b}},W_{i}\rangle$, so we will show a high probability bound for each term.
    We will utilize the contraction property of $\gamma$ for the first term. Notice that $\gamma:[0,L]\to\mathbb{R}^{d}$ is a contraction map, i.e., it has Lipschitz constant 1: $\left\|\gamma(t_{1})-\gamma(t_{2})\right\|\leqslant|t_{1}-t_{2}|$. Recall that in Proposition 6 part (a) we show that conditioned on event that $\zeta_{i}\in\bigcup_{t\leqslant\tau\log n}Z_{t}$ and $Y_{i}\in T$, we have $\Pi_{\gamma}X_{i}\in f^{-1}(T+\bigcup_{t\leqslant\tau\log n}Z_{t})$. As a consequence, the contraction property of $\gamma$ shows that conditioned on the same event, we also have $\gamma(\Pi_{\gamma}X_{i})\in\gamma(f^{-1}(T+\bigcup_{t\leqslant\tau\log n}Z_{t}))$ whose diameter is bounded by diameter of $f^{-1}(T+\bigcup_{t\leqslant\tau\log n}Z_{t})$. Following the proof of Proposition 6, we have
    $\mathbb{P}\left\{\left\|\gamma(\Pi_{\gamma}X_{i})-\mathbb{E}[\gamma(\Pi_{\gamma}X_{i})\mid Y\in T,\mathcal{B}]\right\|\gtrsim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})\mid Y_{i}\in T,\mathcal{B}_{i}\right\}\leqslant 2n^{-\tau}$
    $\left\|\operatorname{{Cov}}[\gamma(\Pi_{\gamma}X)|Y\in T,\mathcal{B}]\right\|\lesssim C_{f}^{2}(|T|^{2}+\sigma_{\zeta}^{2})$
    and as a consequence,
    $\mathbb{P}\left\{\left|\langle{v_{l,h}^{b}},\gamma(\Pi_{\gamma}X_{i})\rangle-\mathbb{E}[\langle{v_{l,h}^{b}},\gamma(\Pi_{\gamma}X)\rangle\mid Y\in T,\mathcal{B}]\right|\gtrsim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})\mid Y_{i}\in T,\mathcal{B}_{i}\right\}\leqslant 2n^{-\tau}$
    $\mathrm{Var}[\langle{v_{l,h}^{b}},\gamma(\Pi_{\gamma}X)\rangle|Y\in T,\mathcal{B}]\lesssim C_{f}^{2}(|T|^{2}+\sigma_{\zeta}^{2})$
    By construction, $W_{i}^{\prime}\in\mathbb{R}^{d-1}$ are independent, identical, and centered distribution on $B(0,\mathrm{reach}_{\gamma})\subseteq\mathbb{R}^{d-1}$ (each $W_{i}^{\prime}$ may be dependent of $t_{i}=\Pi_{\gamma}X_{i}$). Moreover, conditioned on $Y_{i}\in T$ and $\mathcal{B}_{i}$, the geometric assumption ($\bm{\gamma_{1}}$) implies that
    $\left|\langle W_{i},{v_{l,h}^{b}}\rangle\right|\leqslant 2\left|\Pi_{\gamma}X_{i}-\mathbb{E}\left[\Pi_{\gamma}X|Y\in T,\mathcal{B}\right]\right|\,.$
    Follow the same procedure as in Proposition 6 part (a) we conclude that
    $\mathbb{P}\left\{\left|\langle{v_{l,h}^{b}},W_{i}\rangle-\mathbb{E}[\langle{v_{l,h}^{b}},W\rangle\mid Y\in T,\mathcal{B}]\right|\gtrsim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})\mid Y_{i}\in T,\mathcal{B}_{i}\right\}\leqslant 2n^{-\tau}\,,$
    $\mathrm{Var}[\langle{v_{l,h}^{b}},W\rangle|Y\in T,\mathcal{B}]\lesssim C_{f}^{2}(|T|^{2}+\sigma_{\zeta}^{2})\,,$
    We combine the above high probability bounds together and conclude that
    $\mathbb{P}\left\{\left|\langle{v_{l,h}^{b}},X_{i}\rangle-\mathbb{E}[\langle{v_{l,h}^{b}},X\rangle\mid Y\in T,\mathcal{B}]\right|\gtrsim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})\mid Y_{i}\in T,\mathcal{B}_{i}\right\}\leqslant 2n^{-\tau}\,,$
    and
    $\lambda_{d}({\Sigma_{l,h}^{b}}):=\mathrm{Var}[\langle{v_{l,h}^{b}},X\rangle|Y\in T,\mathcal{B}]\lesssim C_{f}^{2}(|T|^{2}+\sigma_{\zeta}^{2})\,.$
    By assumption in the Corollary, we have $l\gtrsim C_{f}C_{Y}R_{0}/\sigma_{\gamma}$ and $\frac{C_{Y}R_{0}}{l}\cong|{R_{l,h}^{b}}|\gtrsim\max\left(\sigma_{\zeta},\omega_{f}\right)$. This implies that the above right-hand side is bounded by $C_{f}^{2}|{R_{l,h}^{b}}|^{2}\lesssim\sigma_{\gamma}^{2}$, and on the other hand, assumption (LCV) implies that $\lambda_{d-1}({\Sigma_{l,h}^{b}})\gtrsim\sigma_{\gamma}^{2}$.
    A.3 Proof of Proposition 17 and Corollary 9
    Proposition 17 (local NVM)
    Suppose (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), and ($\bm{\omega_{f}}$) hold true. Let ${\mu_{l,h}^{b}}$ be the mean of the $h$-th slice and ${v_{l,h}^{b}}$ be the significant vector of the $h$-th slice. Then, for every $l$ such that $l\gtrsim C_{f}C_{Y}R_{0}/\sigma_{\gamma}$, $|{R_{l,h}^{b}}|\geqslant\max\{\sigma_{\zeta},\omega_{f}\}$ for all $h\in\mathcal{H}_{l}^{b}$, for every $\epsilon\in(0,1)$ and $\tau\geqslant 1$, we have:
    1. (a)
      for any $h\in\mathcal{H}_{l}$ and any $\epsilon>0$, the estimation error of the slice mean along the tangential direction can be bounded as
      $\mathbb{P}\left\{\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|>\frac{\sigma_{\gamma}^{2}\epsilon}{R_{0}\sqrt{d}}\ \bigg{|}\ {n_{l,h}^{b}}\right\}\lesssim d\exp\left(-\frac{c\sigma_{\gamma}^{4}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{C_{Y}^{2}C_{f}^{2}R_{0}^{4}d(l^{-2}+l^{-1}\epsilon)}\right)+n^{-\tau}\,;$
    2. (b)
      for any $h\in\mathcal{H}_{l}$ and any $\epsilon>0$, for $l\gtrsim C_{f}C_{Y}R_{0}/\sigma_{\gamma}$, the estimation error of the significant vector can be bounded as,
      $\mathbb{P}\left\{\left\|{{\widehat{v}_{l,h}^{b}}}-{{v_{l,h}^{b}}}\right\|>\epsilon\ \big{|}\ {n_{l,h}^{b}}\right\}\lesssim d\exp\left(-\frac{c\sigma_{\gamma}^{4}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{C_{Y}^{2}C_{f}^{2}R_{0}^{4}d(l^{-2}+l^{-1}\epsilon)}\right)+d\exp\left(-\frac{c\sigma_{\gamma}^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)+n^{-\tau}\,;$
    3. (c)
      for any $h\in\mathcal{H}_{l}$, $u>0$, for $l\gtrsim C_{f}C_{Y}R_{0}/\sigma_{\gamma}$, the estimation error of the width of the slice can be bounded as,
      $\mathbb{P}\big{(}\left|{\lambda_{d}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{d}(\Sigma_{l,h}^{b})}\right|>$ $C_{f}^{2}C_{Y}^{2}R_{0}^{2}u^{2}/l^{2}\ \bigg{|}\ {n_{l,h}^{b}}\big{)}\lesssim d\exp\left(-\frac{c\sigma_{\gamma}^{4}C_{Y}^{2}u^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{4}d^{3/2}(\sqrt{d}+uC_{f}C_{Y})}\right)$
      $+\exp\left(-\frac{cu^{4}{n_{l,h}^{b}}}{1+u^{2}}\right)+d\exp\left(-\frac{c\sigma_{\gamma}^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)+n^{-\tau}\,.$
    Proof of part (a): By Proposition 7 part (a), we know that conditioned on $Y_{i}\in{R_{l,h}^{b}}$ and $\mathcal{B}_{i}$, with probability higher than $1-2n^{-\tau}$, we have $\left|\langle{v_{l,h}^{b}},X_{i}\rangle-\mathbb{E}\left[\langle{v_{l,h}^{b}},X\rangle\mid Y\in{R_{l,h}^{b}},\mathcal{B}\right]\right|\lesssim C_{f}C_{Y}R_{0}\sqrt{\tau\log n}l^{-1}$. Also Proposition 7 part (b) implies that $\mathrm{Var}\left[\langle{v_{l,h}^{b}},X\rangle\mid Y\in{R_{l,h}^{b}},\mathcal{B}\right]\lesssim C_{f}^{2}C_{Y}^{2}R_{0}^{2}l^{-2}$. Therefore, we have the following Bernstein-type inequality for any $\epsilon>0$:
    $\mathbb{P}\left(\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|>\frac{\sigma_{\gamma}^{2}\epsilon}{R_{0}\sqrt{d}}\ \Big{|}\ {n_{l,h}^{b}}\right)\lesssim d\exp\left(-\frac{c\sigma_{\gamma}^{4}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{C_{f}^{2}C_{Y}^{2}R_{0}^{4}d(l^{-2}+l^{-1}\epsilon)}\right)+n^{-\tau}$
    Proof of part (b) The main tool is the following Davis-Kahan type inequality. The Davis–Kahan Theorem (Bhatia, 1997, Theorem VII.3.1), together with (Stewart and guang Sun, 1990, Ch. 1, Sec. 5.3, Theorem 5.5) and Corollary 8, gives
    $\left\|{\widehat{v}_{l,h}^{b}}-{v_{l,h}^{b}}\right\|\leqslant\frac{\left\|({\widehat{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|}{\lambda_{d-1}({\widehat{\Sigma}_{l,h}^{b}})-\lambda_{d}({\Sigma_{l,h}^{b}})}$ (6)
    Step1: We want the denominator of right hand side of (6) to be large. By Corollary 8 and Weyl’s inequality (Vershynin, 2018, Theorem 4.5.3) we obtain that for $l\gtrsim C_{f}C_{Y}R_{0}/\sigma_{\gamma}$,
    $\left|\lambda_{d-1}({\widehat{\Sigma}_{l,h}^{b}})-\lambda_{d}({\Sigma_{l,h}^{b}})\right|$ $\geqslant\left|\lambda_{d-1}({\Sigma_{l,h}^{b}})-\lambda_{d}({\Sigma_{l,h}^{b}})\right|-\left|\lambda_{d-1}({\widehat{\Sigma}_{l,h}^{b}})-\lambda_{d}({\Sigma_{l,h}^{b}})\right|\gtrsim\sigma_{\gamma}^{2}-\left\|{\widehat{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|\,.$
    To bound ${\widehat{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}$, we split it as ${\widehat{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}={\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}+({\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}})({\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}})^{\intercal}$, where we introduced the intermediate-term
    ${\widetilde{\Sigma}_{l,h}^{b}}=\frac{1}{{n_{l,h}^{b}}}\sum_{i}(X_{i}-{\mu_{l,h}^{b}})(X_{i}-{\mu_{l,h}^{b}})^{\intercal}\mathbbm{1}{\{Y_{i}\in{R_{l,h}^{b}}\}\cap\mathcal{B}_{i}}\,.$
    We use Lemma 16 to obtain a high probability bound on $\left\|{\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|$ and $\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|^{2}$: for a fixed constant $\beta<\min(\frac{1}{2},\alpha d)=\frac{1}{2}$,
    $\mathbb{P}\left(\max\left(\left\|{\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|,\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|^{2}\right)\gtrsim\beta\sigma_{\gamma}^{2}\right)\lesssim d\exp\left(-\frac{c\beta^{2}\sigma_{\gamma}^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)\,.$
    This will show that the denominator of the right hand side of (6) is $\left|\lambda_{d-1}({\widehat{\Sigma}_{l,h}^{b}})-\lambda_{d}({\Sigma_{l,h}^{b}})\right|\gtrsim\sigma_{\gamma}^{2}$.
    Step 2: We are going to apply Bernstein inequality to upper bound the numerator $\left\|({\widehat{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|$ in the right-hand side of (6). Consider the following decomposition:
    $\left\|({\widehat{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|$ $\leqslant\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|+\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|$
    $\lesssim$ $\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|+R_{0}\sqrt{d}\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|$
    Recall that part (a) already gives desired bound for $\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|$, so we only need to bound $\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|$. Observe that $\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|$ has a priori upper-bound by $|{v_{l,h}^{b}}^{\intercal}(X_{i}-{\mu_{l,h}^{b}})|\left\|X_{i}-{\mu_{l,h}^{b}}\right\|$, Moreover, by Proposition 7 part (a), conditioned on $Y_{i}\in S_{l,h}$ and $\mathcal{B}_{i}$, we have, with probability no lower than $1-2n^{-\tau}$, $|{v_{l,h}^{b}}^{\intercal}(X_{i}-{\mu_{l,h}^{b}})|\left\|X_{i}-{\mu_{l,h}^{b}}\right\|\lesssim C_{f}C_{Y}R_{0}^{2}\sqrt{d\tau\log n}l^{-1}$. This serves as an $\ell^{\infty}$ bound on $\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|$.
    Next, we consider the $\ell^{2}$ bound (i.e., variance). Considering the following decomposition:
    $\mathbb{E}\left[\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|^{2}\mid{n_{l,h}^{b}}\right]=\mathbb{E}\left[\left\|{\widetilde{\Sigma}_{l,h}^{b}}{v_{l,h}^{b}}\right\|^{2}\mid{n_{l,h}^{b}}\right]-\left\|{\Sigma_{l,h}^{b}}{v_{l,h}^{b}}\right\|^{2}\,,$
    where
    $\mathbb{E}\left[\left\|{\widetilde{\Sigma}_{l,h}^{b}}{v_{l,h}^{b}}\right\|^{2}\mid{n_{l,h}^{b}}\right]$ $=\frac{1}{({n_{l,h}^{b}})^{2}}{v_{l,h}^{b}}^{\intercal}\mathbb{E}\left[\left(\sum_{i}(X_{i}-{\mu_{l,h}^{b}})(X_{i}-{\mu_{l,h}^{b}})^{\intercal}\right)^{2}\mid{n_{l,h}^{b}}\right]{v_{l,h}^{b}}$
    $\leqslant$ $\frac{R_{0}^{2}d}{{n_{l,h}^{b}}}\mathrm{Var}\left[\langle X_{i}-{\mu_{l,h}^{b}},{v_{l,h}^{b}}\rangle\mid{n_{l,h}^{b}}\right]+\left\|{\Sigma_{l,h}^{b}}{v_{l,h}^{b}}\right\|^{2}\,.$
    The above inequality, together with Proposition 7 part (b), gives us the following $\ell^{2}$ bound:
    $\mathbb{E}\left[\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|^{2}\mid{n_{l,h}^{b}}\right]\lesssim\frac{C_{f}^{2}C_{Y}^{2}R_{0}^{4}d}{{n_{l,h}^{b}}l^{2}}\,.$
    We use the $\ell^{\infty}$ and $\ell^{2}$ bounds above and apply Bernstein inequality 15 to obtain: for any $\epsilon>0$,
    $\mathbb{P}\left(\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|>\sigma_{\gamma}^{2}\epsilon\ \Big{|}\ {n_{l,h}^{b}}\right)\lesssim d\exp\left(-\frac{c\sigma_{\gamma}^{4}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{C_{f}^{2}C_{Y}^{2}R_{0}^{4}d(\epsilon l^{-1}+l^{-2})}\right)+n^{-\tau}\,.$
    Combining part (a) and the estimates in Step1-2 finishes the proof.
    Proof of part (c) Let $V_{i}=\langle{v_{l,h}^{b}},X_{i}-{\mu_{l,h}^{b}}\rangle^{2}$ then $\mathbb{E}[V_{i}\mid Y_{i}\in{R_{l,h}^{b}},\mathcal{B}_{i}]={\lambda_{d}(\Sigma_{l,h}^{b})}\lesssim C_{f}^{2}C_{Y}^{2}R_{0}^{2}l^{-2}$. Moreover, we can follow the same argument as in Proposition 6 and Proposition 7 to show that $\mathbb{E}[V_{i}^{2}\mid Y_{i}\in{R_{l,h}^{b}},\mathcal{B}_{i}]\lesssim C_{f}^{4}C_{Y}^{4}R_{0}^{4}l^{-4}$. Denote $\widehat{V}_{l,h}^{b}=\frac{1}{{n_{l,h}^{b}}}\sum_{i}\langle{v_{l,h}^{b}},X_{i}-{\mu_{l,h}^{b}}\rangle^{2}\mathbbm{1}\left\{Y_{i}\in{R_{l,h}^{b}}\right\}\cap\mathcal{B}_{i}$. Thus, we use Bernstein’s inequality to show that
    $\mathbb{P}\left(|\widehat{V}_{l,h}^{b}-{\lambda_{d}(\Sigma_{l,h}^{b})}|>\beta\sigma_{\gamma}^{2}\ \Big{|}\ {n_{l,h}^{b}}\right)\lesssim\exp\left(-\frac{c\beta^{2}\sigma_{\gamma}^{2}{n_{l,h}^{b}}}{\alpha C_{f}^{2}C_{Y}^{2}R_{0}^{2}l^{-4}(\beta l^{2}+C_{f}^{2}C_{Y}^{2}R_{0}^{2}\sigma_{\gamma}^{-2})}\right)\,.$
    We use $\bm{\mathfrak{S}}_{h}(\epsilon,\beta)$ to denote the event that
    $\bm{\mathfrak{S}}_{h}=\left\{\begin{split}&\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|<\frac{R_{0}\sqrt{d}}{2},\left|\langle{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}},{v_{l,h}^{b}}\rangle\right|\lesssim\frac{\epsilon\sigma_{\gamma}^{2}}{R_{0}\sqrt{d}},\left\|{v_{l,h}^{b}}-{\widehat{v}_{l,h}^{b}}\right\|\lesssim\epsilon,\\ &\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|<\epsilon\sigma_{\gamma}^{2},\left|\widehat{V}_{l,h}^{b}-{\lambda_{d}(\Sigma_{l,h}^{b})}\right|\lesssim\beta\sigma_{\gamma}^{2}\end{split}\right\}$
    We know from part (a), part (b), and Lemma 16 that, $\bm{\mathfrak{S}}_{h}(\epsilon,\beta)$ satisfies, for any $\epsilon,\beta>0$,
    $\mathbb{P}\left(\bm{\mathfrak{S}}_{h}(\epsilon,\beta)^{c}\right)\lesssim$ $d\exp\left(-\frac{c\sigma_{\gamma}^{4}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{C_{f}^{2}C_{Y}^{2}R_{0}^{4}d(l^{-2}+\epsilon l^{-1})}\right)$
    $+\exp\left(-\frac{c\sigma_{\gamma}^{2}\beta^{2}{n_{l,h}^{b}}}{C_{f}^{2}C_{Y}^{2}R_{0}^{2}l^{-4}(\beta l^{2}+C_{f}^{2}C_{Y}^{2}R_{0}^{2}\sigma_{\gamma}^{-2})}\right)+d\exp\left(-\frac{c\sigma_{\gamma}^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)+n^{-\tau}$
    Conditioned on the event $\bm{\mathfrak{S}}_{h}(\eta,\epsilon,\beta)$, we have the following estimate:
    $\left|{\lambda_{d}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{d}(\Sigma_{l,h}^{b})}\right|=$ $\left|\left|\langle{\widehat{v}_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|^{2}+\langle{\widehat{v}_{l,h}^{b}},({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){\widehat{v}_{l,h}^{b}}\rangle+\langle{\widehat{v}_{l,h}^{b}},{\Sigma_{l,h}^{b}}{\widehat{v}_{l,h}^{b}}\rangle-\langle{v_{l,h}^{b}},{\Sigma_{l,h}^{b}}{v_{l,h}^{b}}\rangle\right|$
    $\lesssim$ $R_{0}^{2}d\epsilon^{2}+\frac{\epsilon^{2}\sigma_{\gamma}^{4}}{R_{0}^{2}d}+\beta\sigma_{\gamma}^{2}+\epsilon^{2}\sigma_{\gamma}^{2}+R_{0}^{2}d\epsilon^{2}+{\lambda_{d}(\Sigma_{l,h}^{b})}\epsilon^{2}\,,$
    which is bounded by $\frac{u^{2}C_{f}^{2}C_{Y}^{2}R_{0}^{2}}{l^{2}}$ if one takes $\epsilon^{\prime}=c\frac{uC_{f}C_{Y}}{l\sqrt{d}}$ and $\beta^{\prime}=\frac{u^{2}C_{f}^{2}C_{Y}^{2}R_{0}^{2}}{\sigma_{\gamma}^{2}l^{2}}$, for $l>C_{f}C_{Y}R_{0}/\sigma_{\gamma}$. This means that we have for any $u>0$,
    $\mathbb{P}\left(\left|{\lambda_{d}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{d}(\Sigma_{l,h}^{b})}\right|\gtrsim\frac{u^{2}C_{f}^{2}C_{Y}^{2}R_{0}^{2}}{l^{2}}\ \bigg{|}\ {n_{l,h}^{b}}\right)\leqslant\mathbb{P}\left(\bm{\mathfrak{S}}_{h}(\epsilon^{\prime},\beta^{\prime})^{c}\right)$
    $\lesssim$ $d\exp\left(-\frac{cu^{2}C_{Y}^{2}\sigma_{\gamma}^{4}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{4}d^{3/2}(\sqrt{d}+uC_{f}C_{Y})}\right)+\exp\left(-\frac{c{n_{l,h}^{b}}u^{4}}{1+u^{2}}\right)+d\exp\left(-\frac{c\sigma_{\gamma}^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)+n^{-\tau}\,.$
    A.3.1 Proof of Corollary 9
    Proof  We simply take $\epsilon=C_{f}C_{Y}R_{0}^{2}\sigma_{\gamma}^{-2}\left(d(t+\log l+\log d)\frac{\sqrt{\tau\log n}}{nl}\right)^{\frac{1}{2}}$. The role of $\log d$ is to cancel the constant $d$ before the exponential term. Also, since now we consider all slices $h\in\mathcal{H}_{l}^{b}$, there will be an extra constant $l$ before the exponential term, and thus, we include $\log l$ to cancel this extra coefficient $l$ before the exponential term. We want $\epsilon<\frac{1}{l}$, and this gives us the requirement that $\frac{n}{\sqrt{\tau\log n}}\gtrsim C_{f}^{2}C_{Y}^{2}R_{0}^{4}\sigma_{\gamma}^{-4}dl(t+\log d+\log l)$. The expectation estimate can be derived by taking $e^{-t}\lesssim(C_{f}C_{Y}R_{0}^{2}\sigma_{\gamma}^{-2})^{2p}(d\log d)^{p}\left(\frac{\log n\log l}{nl}\right)^{p}$ and use conditional expectation formula.
    A.4 Proof of Proposition 10
    Proof  We use $C(c_{1})$ to denote some positive constant that increases with $c_{1}$ and $C(c_{1})\to\infty$ as $c_{1}\to\infty$. The value of $C(c_{1})$ may change from line to line and depend on other constants.
    Fix $y_{0}\in R,t_{0}\in f^{-1}(y_{0})$ and fix $x_{0}\in F^{-1}(y_{0})$ then we have $\Pi_{\gamma}x_{0}=t_{0}$. Let ${h^{\prime}_{x}}$ be the unique index $h\in\{1,\dots,l\}$ such that $y_{0}\in{R_{l,h}}$. Without loss of generality, suppose that $|y_{0}-\min R_{l,{h^{\prime}_{x}}}|\leqslant|y_{0}-\max R_{l,{h^{\prime}_{x}}}|$. Then the standing assumption $|{R_{l,h}}|\geqslant 2c_{1}\max(\sigma_{\zeta},\omega_{f})$ implies that for any $h\not\in\{{h^{\prime}_{x}},{h^{\prime}_{x}}-1\}$, $\min\{|y_{0}-y|:y\in{R_{l,h}}\}\geqslant c_{1}\max(\sigma_{\zeta},\omega_{f})$. Suppose that either ${h^{\prime}_{x}}$ or ${h^{\prime}_{x}}-1\in\mathcal{H}_{l}$.
    For each $h\in\mathcal{H}_{l}$, each slice ${S_{l,h}}$ is a conditional distribution $X|Y\in{R_{l,h}}$. For each $h\in\mathcal{H}_{l}$, we consider the push-forward conditional distribution $\Pi_{\gamma}X|Y\in{R_{l,h}}$ and denote its density function $\rho_{t|Y\in{R_{l,h}}}(\cdot)$. We further use $\rho_{t}(\cdot)$ to denote the density function of the push-forward distribution $\Pi_{\gamma}X$. By Bayes’ rule, for each $h\in\mathcal{H}_{l}$
    $\rho_{t|Y\in{R_{l,h}}}(t_{0})=\frac{\mathbb{P}(\zeta\in{R_{l,h}}-y_{0})}{\mathbb{P}(Y\in{R_{l,h}})}\rho_{t}(t_{0})\,.$
    We claim that either ${h^{\prime}_{x}}$ or ${h^{\prime}_{x}}-1$ is the maximizer of the left-hand side over $h\in\mathcal{H}_{l}$. Indeed, if $h\not\in\{{h^{\prime}_{x}},{h^{\prime}_{x}}-1\}$, then the property $\min\{|y_{0}-y|:y\in{R_{l,h}}\}\geqslant c_{1}\max(\sigma_{\zeta},\omega_{f})$ implies that the numerator $\mathbb{P}(\zeta\in{R_{l,h}}-y_{0})\leqslant 2\exp(-Cc_{1}^{2})$ for some absolute constant $C$. On the other hand, $\max_{h\in\{{h^{\prime}_{x}},{h^{\prime}_{x}}-1\}}\mathbb{P}(\zeta\in{R_{l,h}}-y_{0})\geqslant\frac{1}{2}\mathbb{P}(\zeta\in R_{l,{h^{\prime}_{x}}}\cup R_{{h^{\prime}_{x}}-1}-y_{0})\geqslant\frac{1}{2}(1-2\exp(-Cc_{1}^{2}))$. Therefore, $\frac{\max_{h\in\{{h^{\prime}_{x}},{h^{\prime}_{x}}-1\}}\mathbb{P}(\zeta\in R_{l,{h^{\prime}_{x}}}-y_{0})}{\max_{h\not\in\{{h^{\prime}_{x}},{h^{\prime}_{x}}-1\}}\mathbb{P}(\zeta\in R_{l,{h^{\prime}_{x}}}-y_{0})}\geqslant C(c_{1})$. Moreover, the term $\rho_{t}(t_{0})$ is independent of $h\in\mathcal{H}_{l}$. The a priori probability $\mathbb{P}(Y\in{R_{l,h}})$ should be comparable among $h\in\mathcal{H}_{l}$, that is, $\sup_{h\in\mathcal{H}_{l}}\mathbb{P}(Y\in{R_{l,h}})\leqslant c\mathbb{P}(Y\in{R_{l,h}})$ for some universal constant $c>0$. This is because the construction of $\mathcal{H}_{l}$ has discarded slices with little probability (or data, in the empirical version) and only consider slices with sufficient probability mass in Algorithm 1. As a consequence, $\frac{\max_{h\in\{{h^{\prime}_{x}},{h^{\prime}_{x}}-1\}}\rho_{t|Y\in{R_{l,h}}}(t_{0})}{\max_{h\not\in\{{h^{\prime}_{x}},{h^{\prime}_{x}}-1\}}\rho_{t|Y\in{R_{l,h}}}(t_{0})}\geqslant C(c_{1})$ for some constant $C(c_{1})$ which increases with $c_{1}$.
    Now we introduce a term that is an integral of the density $\rho_{t|Y\in{R_{l,h}}}$:
    $Q_{h}(t_{0}):=\min\left(\mathbb{P}(t\in(0,t_{0})|Y\in{R_{l,h}})\,,\,\mathbb{P}(t\in(t_{0},\mathrm{len}_{\gamma})|Y\in{R_{l,h}})\right)$
    The same argument shows that $\frac{\max_{h\in\{{h^{\prime}_{x}},{h^{\prime}_{x}}-1\}}Q_{h}(t_{0})}{\max_{h\not\in\{{h^{\prime}_{x}},{h^{\prime}_{x}}-1\}}Q_{h}(t_{0})}\geqslant C(c_{1})$. Notice that the term $Q_{h}(t_{0})$ takes the smaller conditional probability by comparing two tails $(0,t_{0})$ and $(t_{0},\mathrm{len}_{\gamma})$ for the conditional distribution $\Pi_{\gamma}X|Y\in{R_{l,h}}$. If we center the random variable $\Pi_{\gamma}X|Y\in{R_{l,h}}$ and consider $W_{h}=(\Pi_{\gamma}X-\mathbb{E}(\Pi_{\gamma}X|Y\in{R_{l,h}}))|Y\in{R_{l,h}}$ instead, we can show that $Q_{h}(t_{0})$ equals
    $Q_{h}(t_{0})=\mathbb{P}(|W_{h}|>|t_{0}-\mathbb{E}(\Pi_{\gamma}X|Y\in{R_{l,h}})|)\,.$
    Recall that the proof of Proposition 6 also shows that the variances for $W_{h}$ are comparable among $h\in\mathcal{H}_{l}$. That is, $C_{f}^{\prime 2}(|{R_{l,h}}|^{2}+\sigma_{\zeta}^{2})\lesssim\mathrm{Var}(W_{h})\lesssim C_{f}^{2}(|{R_{l,h}}|^{2}+\sigma_{\zeta}^{2})$. It follows similarly that $\frac{\min_{h\not\in\{{h^{\prime}_{x}},{h^{\prime}_{x}}-1\}}|t_{0}-\mathbb{E}(\Pi_{\gamma}X|Y\in{R_{l,h}})|^{2}}{\min_{h\in\{{h^{\prime}_{x}},{h^{\prime}_{x}}-1\}}|t_{0}-\mathbb{E}(\Pi_{\gamma}X|Y\in{R_{l,h}})|^{2}}\geqslant C(c_{1})$.
    It is readily that $|t_{0}-\mathbb{E}(\Pi_{\gamma}X|Y\in{R_{l,h_{x}}})|^{2}\lesssim C_{f}^{2}(|{R_{l,h}}|^{2}+\sigma_{\zeta}^{2})$ and $\min_{h\neq{h_{x}}}d(x,h)\gtrsim C_{f}^{\prime 2}(|{R_{l,h}}|^{2}+\sigma_{\zeta}^{2})$. Moreover, for any $h\neq{h_{x}}$, we have $|t_{0}-\mathbb{E}(\Pi_{\gamma}X|Y\in{R_{l,h_{x}}})|^{2}\gtrsim d(x,h)$. Now, properly choosing constant $c_{1}$, we can show that either ${h_{x}}={h^{\prime}_{x}}$ or ${h^{\prime}_{x}}-1$. Suppose for a moment that ${h_{x}}\not\in\{{h^{\prime}_{x}},{h^{\prime}_{x}}-1\}$, then we have
    $C_{f}^{2}(|{R_{l,h}}|^{2}+\sigma_{\zeta}^{2})\gtrsim|t_{0}-\mathbb{E}(\Pi_{\gamma}X|Y\in{R_{l,h_{x}}})|^{2}\gtrsim C(c_{1})\min_{h={h^{\prime}_{x}},{h^{\prime}_{x}}-1}|t_{0}-\mathbb{E}(\Pi_{\gamma}X|Y\in{R_{l,h_{x}}})|^{2}$
    $\gtrsim$ $C(c_{1})\min_{h={h^{\prime}_{x}},{h^{\prime}_{x}}-1}d(x,h)\gtrsim C(c_{1})C_{f}^{\prime 2}(|{R_{l,h}}|^{2}+\sigma_{\zeta}^{2})$
    which will not hold as long as we properly choose $c_{1}$ such that $C(c_{1})$ is sufficiently large.
    A.5 Proof of Proposition 11
    Proof  we know that in small neighborhoods, the curve can be viewed as slightly curved, that is, there exists $K_{0}>0$ depending only on the curvature of $\gamma$ such that for $2\leqslant|k|\leqslant K_{0}l$, we have the following inequality for nearby slices
    $\sqrt{{\mathrm{dist}}(x,{h_{x}}+k)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim\sum_{k^{\prime}=1}^{|k|}|\Delta\gamma_{{h_{x}}+k^{\prime}}|\cong|k|\mathrm{len}_{\gamma}/l$
    Note that we do not have such inequality for $|k|=1$ because points near the boundary of one slice may share a very similar distance to the adjacent slice, thus hand to distinct true slice index ${h_{x}}$ from adjacent one ${h_{x}}\pm 1$. This is why we only prove misclassification by at least two slices.
    On the other hand, for far-away slices, we can bound the difference in distance function by the reach of the curve. Given $l\gtrsim C_{f}C_{Y}R_{0}$, we deduce that for $|k|\geqslant K_{0}l$, we have the following inequality for far-away slices
    $\sqrt{{\mathrm{dist}}(x,{h_{x}}+k)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim K_{0}\sqrt{\frac{{\lambda_{d}(\Sigma_{l,h}^{b})}}{{\lambda_{1}(\Sigma_{l,h}^{b})}}}\mathrm{reach}_{\gamma}\gtrsim\frac{\mathrm{len}_{\gamma}}{l}\frac{\mathrm{reach}_{\gamma}}{\sigma_{\gamma}}\gtrsim\mathrm{len}_{\gamma}/l$
    As a consequence, we take all $|k|\geqslant 2$ and have the following inequality:
    $\text{For any }h\text{ such that }|h-{h_{x}}|\geqslant 2,\ \sqrt{{\mathrm{dist}}(x,h)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim\mathrm{len}_{\gamma}/l$
    In order to obtain correct classification, we want the estimation error of the distance function to be small, such that for all $|h^{\prime}-{h_{x}}|\geqslant 2$,
    $\left|{\widehat{\mathrm{dist}}}(x,h^{\prime})-{\mathrm{dist}}(x,h^{\prime})\right|+\left|{\widehat{\mathrm{dist}}}(x,{h_{x}})-{\mathrm{dist}}(x,{h_{x}})\right|<\left|{\mathrm{dist}}(x,h^{\prime})-{\mathrm{dist}}(x,{h_{x}})\right|$ (7)
    Indeed, this will imply that ${\widehat{h}_{x}}=\operatorname*{{arg\,min}}_{h^{\prime}\in{\mathcal{H}_{l}^{b}}}{\widehat{\mathrm{dist}}}(x,h^{\prime})$ is the correct or adjacent classification, i.e. $|{h_{x}}-{\widehat{h}_{x}}|\leqslant 1$.
    Consider the event $\bm{\mathfrak{S}}$ that we have a small estimation error for information in all slices:
    $\bm{\mathfrak{S}}(\epsilon,\delta,\beta,u)=\left\{\begin{split}&\text{ For all}h\in{\mathcal{H}_{l}^{b}},\left|\langle{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}},{v_{l,h}^{b}}\rangle\right|\lesssim\frac{\epsilon\sigma_{\gamma}^{2}}{R_{0}\sqrt{d}},\ \left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|\lesssim\sigma_{\gamma}\sqrt{\delta},\\ &\left\|{v_{l,h}^{b}}-{\widehat{v}_{l,h}^{b}}\right\|\lesssim\epsilon,\left|{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{1}(\Sigma_{l,h}^{b})}\right|\lesssim\beta\sigma_{\gamma}^{2},\left|{\lambda_{d}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{d}(\Sigma_{l,h}^{b})}\right|\lesssim\frac{u^{2}C_{f}^{2}C_{Y}^{2}R_{0}^{2}}{l^{2}}\end{split}\right\}$
    Notice that Proposition 17 and Lemma 16 state that, with $l>C_{f}C_{Y}R_{0}/\sigma_{\gamma}$, for any $\delta,\beta<R_{0}^{2}d\sigma_{\gamma}^{-2}$ and any $\epsilon,u>0$, event $\bm{\mathfrak{S}}$ has the following high probability bound:
    $\mathbb{P}\left(\bm{\mathfrak{S}}^{c}\right)\lesssim$ $ld\exp\left(-\frac{c\sigma_{\gamma}^{4}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{C_{f}^{2}C_{Y}^{2}R_{0}^{4}d(l^{-2}+l^{-1}\epsilon)}\right)+ld\exp\left(-c{n_{l,h}^{b}}\min\left(\frac{\sigma_{\gamma}^{2}\delta}{R_{0}^{2}d},\frac{\sigma_{\gamma}^{4}\beta^{2}}{R_{0}^{4}d^{2}},\frac{\sigma_{\gamma}^{4}}{R_{0}^{4}d^{2}}\right)\right)+ln^{-\tau}$
    $+ld\exp\left(-\frac{c\sigma_{\gamma}^{4}C_{Y}^{2}u^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{4}d^{3/2}(\sqrt{d}+uC_{f}C_{Y})}\right)+l\exp\left(-\frac{c{n_{l,h}^{b}}u^{4}}{1+u^{2}}\right)$
    We now investigate how small these parameters should be. Conditioned on event $\bm{\mathfrak{S}}(\epsilon,\delta,\beta,u)$, we can expand the estimation error in the distance function and estimate its upper-bound by the following calculation:
    $\left|{\widehat{\mathrm{dist}}}(x,h)-{\mathrm{dist}}(x,h)\right|$
    $\lesssim$ $\left(R_{0}\sqrt{d}\epsilon+\sigma_{\gamma}\sqrt{\delta}\epsilon+\frac{\sigma_{\gamma}^{2}\epsilon}{R_{0}\sqrt{d}}\right)^{2}+\left(R_{0}\sqrt{d}\epsilon+\sigma_{\gamma}\sqrt{\delta}\epsilon+\frac{\sigma_{\gamma}^{2}\epsilon}{R_{0}\sqrt{d}}\right)\sqrt{{\mathrm{dist}}(x,h)}$
    $+\frac{{\lambda_{d}(\Sigma_{l,h}^{b})}}{{\lambda_{1}(\Sigma_{l,h}^{b})}}\left(\sigma_{\gamma}^{2}\delta+R_{0}\sigma_{\gamma}\sqrt{d}\sqrt{\delta}\right)+\frac{u^{2}\mathrm{len}_{\gamma}^{2}}{l^{2}}\frac{R_{0}^{2}d}{{\lambda_{1}(\Sigma_{l,h}^{b})}}+\beta\sigma_{\gamma}^{2}\frac{{\lambda_{d}(\Sigma_{l,h}^{b})}R_{0}^{2}d}{{\lambda_{1}(\Sigma_{l,h}^{b})}^{2}}$
    Thus in above inequality, we want the coefficient before $\sqrt{{\mathrm{dist}}(x,h)}$ to be smaller than $c\frac{\mathrm{len}_{\gamma}}{l}$, and all other terms to be smaller than $c\frac{\mathrm{len}_{\gamma}^{2}}{l^{2}}$, so that (7) can be guaranteed. To achieve this, we let small scales $\epsilon,\delta,\beta,u$ to be the following: with $l\gtrsim C_{f}C_{Y}R_{0}/\sigma_{\gamma}$,
    $\epsilon^{\prime}=c\frac{\mathrm{len}_{\gamma}}{lR_{0}\sqrt{d}},\ \ \ \delta^{\prime}=c\frac{\sigma_{\gamma}^{2}}{R_{0}^{2}d},\ \ \ \beta^{\prime}=c\frac{\sigma_{\gamma}^{2}C_{f}^{\prime 2}}{R_{0}^{2}C_{f}^{2}d},\\ \ u^{\prime}=c\frac{\sigma_{\gamma}}{R_{0}\sqrt{d}}\,.$
    Here $\delta^{\prime},\beta^{\prime}<R_{0}^{2}d\sigma_{\gamma}^{-2}$ automatically holds because of $\sigma_{\gamma}\leqslant R_{0}\leqslant R_{0}\sqrt{d}$ given by Assumption (LCV). Therefore,
    $\mathbb{P}\left(\bm{\mathfrak{S}}(\epsilon^{\prime},\delta^{\prime},\beta^{\prime},u^{\prime})^{c}\right)\lesssim$ $ld\exp\left(-c\frac{nl^{-1}}{\sqrt{\tau\log n}}\min\left(\frac{C_{f}^{\prime}\sigma_{\gamma}^{4}}{C_{f}^{2}R_{0}^{3}d^{3/2}\mathrm{len}_{\gamma}},\frac{\sigma_{\gamma}^{8}C_{f}^{\prime 4}}{R_{0}^{8}C_{f}^{4}d^{4}}\right)\right)+ln^{-\tau}$
    where we have used the fact that $\frac{R_{0}}{\sqrt{\alpha}}<\sigma_{\gamma}\lesssim R_{0}$ and $C_{f}^{\prime}C_{Y}R_{0}\leqslant\mathrm{len}_{\gamma}\leqslant C_{f}C_{Y}R_{0}$.
    A.6 Proof of Proposition 12
    Proof  Let $s\wedge 1={s\wedge 1}\in[\frac{1}{2},1]$. Then we have $s\wedge 1\geqslant\frac{2s}{2s+1}$ for all $s\geqslant\frac{1}{2}$. First, we decompose $X=\gamma(\Pi_{\gamma}X)+M_{\gamma^{\prime}(\Pi_{\gamma}X)}(Z_{d-1},0)^{\intercal}$. We are going to control the curve approximation error
    $\operatorname{\mathrm{MSE}}_{(NCA)}:=\mathbb{E}\left[\left|f(\Pi_{\gamma}X)-\widetilde{f}(\langle{v_{l,h}^{b}},X\rangle)\right|^{2}\mathbbm{1}_{I^{(l,h)}}(\langle{v_{l,h}^{b}},X\rangle)\ \big{|}\ X\in{S_{l,h}}\right]$
    Denote $t_{1}\in\Pi_{\gamma}({S_{l,h}})$ to be such that $\gamma^{\prime}(t_{1})={v_{l,h}^{b}}$. Recall that $\widetilde{f}(s)=\mathbb{E}\left[F(X)\ \big{|}\ X\in{S_{l,h}},\ \langle{v_{l,h}^{b}},X\rangle=s\right]$. As a consequence, we can compute that $\widetilde{f}(\langle{v_{l,h}^{b}},\gamma(t_{1})\rangle)=F(\gamma(t_{1}))=f(t_{1})=f(\langle{v_{l,h}^{b}},\gamma(t_{1})\rangle+c(h))$ where alignment constant is defined as $c(h)=t_{1}-\langle\gamma(t_{1}),{v_{l,h}^{b}}\rangle$.
    Meanwhile, because $\widetilde{f}$ minimizes the population loss
    $\mathbb{E}\left[\left|Y-p(\langle{v_{l,h}^{b}},X\rangle)\right|^{2}\mathbbm{1}_{I^{(l,h)}}(\langle{v_{l,h}^{b}},X\rangle)\ \big{|}\ X\in{S_{l,h}}\right]\,,$
    we obtain an upper-bound upon replacing $\widetilde{f}(\cdot)$ by $f(\cdot+c(h))$ By Hölder continuity, the triangle inequality, and Taylor expansion, we have
    $\begin{split}\operatorname{\mathrm{MSE}}_{(NCA)}\leqslant&\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\mathbb{E}\left[\left|\Pi_{\gamma}X-\langle X,{v_{l,h}^{b}}\rangle-c(h)\right|^{2(s\wedge 1)}\ \big{|}\ X\in{S_{l,h}}\right]\\ \leqslant&C(s)\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\mathbb{E}\left[\left|\langle\gamma(\Pi_{\gamma}X)-\gamma(t_{1}),{v_{l,h}^{b}}\rangle-(\Pi_{\gamma}X-t_{1})\right|^{2(s\wedge 1)}\ \big{|}\ X\in{S_{l,h}}\right]\\ &+C(s)\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\mathbb{E}\left[\left|(Z_{d-1},0)^{\intercal}M_{\gamma^{\prime}(t)}^{\intercal}({v_{l,h}^{b}}-\gamma^{\prime}(t))\right|^{2(s\wedge 1)}\ \big{|}\ X\in{S_{l,h}}\right]\\ \lesssim&\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\sigma_{\gamma}^{2(s\wedge 1)}\left\|\gamma^{\prime\prime}\right\|^{2(s\wedge 1)}C_{f}^{2(s\wedge 1)}\max\left(\sigma_{\zeta},\omega_{f},\frac{\mathrm{len}_{\gamma}}{C_{f}^{\prime}l}\right)^{2(s\wedge 1)}\end{split}\,.$ (NL)
    We use Hölder continuity property, Cauchy-Schwarz inequality, and Corollary 9 to control the following error term
    $\begin{split}\operatorname{\mathrm{MSE}}_{(\Phi)}:=&\mathbb{E}\left[\left|\widetilde{f}(\langle{v_{l,h}^{b}},X\rangle)-\widetilde{f}(\langle{\widehat{v}_{l,h}^{b}},X\rangle)\right|^{2}\mathbbm{1}_{I^{(l,h)}}(\langle{v_{l,h}^{b}},X\rangle)\ \big{|}\ X\in{S_{l,h}}\right]\\ \leqslant&(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}(R_{0}\sqrt{d})^{2(s\wedge 1)}\mathbb{E}\left[\left\|{\widehat{v}_{l,h}^{b}}-{v_{l,h}^{b}}\right\|^{2(s\wedge 1)}\right]\\ \lesssim&(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}(\mathrm{len}_{\gamma}R_{0}^{2}\sigma_{\gamma}^{-2})^{2(s\wedge 1)}(d^{2}\log d)^{s\wedge 1}\left(\frac{\log n\log l}{nl}\right)^{2(s\wedge 1)}\,.\end{split}$ ($\Phi$)
    We exploit the Hölder continuity of $\widetilde{f}$ (see Liao et al., 2022, Appendix A, Example 1) to control the bias error term
    $\begin{split}\operatorname{\mathrm{MSE}}_{(B)}&:=\mathbb{E}\left[\left|\widetilde{f}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{f_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|^{2}\mathbbm{1}_{I^{(l,h)}}(\langle{v_{l,h}^{b}},X\rangle)\ \big{|}\ X\in{S_{l,h}}\right]\\ &\lesssim\left[\widetilde{f}\right]_{\mathcal{C}^{s}}^{2}\left(\frac{\left|I^{(l,h)}\right|}{j}\right)^{2s}\lesssim(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})^{2}C_{f}^{2s}\max\left(\sigma_{\zeta},\omega_{f},\frac{\mathrm{len}_{\gamma}}{C_{f}^{\prime}l}\right)^{2s}j^{-2s}\,.\end{split}$ (B)
    The variance term $\left|{f_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{\widehat{f}_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|$ can be concentrated with known calculations, see (Liao et al., 2022, Proposition 2 and Lemma 5):
    $\mathbb{E}\left[\left|{f_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{\widehat{f}_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|^{2}\right]\lesssim\sigma_{\zeta}^{2}\frac{j\log j}{n/l}\,.$ (V)
    $\begin{split}\operatorname{\mathrm{MSE}}_{(\Psi)}&:=\mathbb{E}\left[\left|{\widehat{f}_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{\widehat{f}_{j|{\widehat{v}_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|^{2}\right]\end{split}$ ($\Psi$)
    To control the projection error term $\left|{\widehat{f}_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{\widehat{f}_{j|{\widehat{v}_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|$, we use Lemma 23. For each $h\in\mathcal{H}_{l}$, we take $p_{1}\geqslant 1$, $s\wedge 1$, $g=\widetilde{f}$ , $r=\sigma_{\gamma}$, and apply Lemma 23 to each sample slice $\widehat{S}_{l,h}^{b}$ with data $\{(\langle X_{i},{\widehat{v}_{l,h}^{b}}\rangle,Y_{i}):X_{i}\in\bigcup_{|h^{\prime}-h|\leqslant 1}\widehat{S}_{l,h^{\prime}}^{b}\}$ with $v={v_{l,h}^{b}},{\widehat{v}}={\widehat{v}_{l,h}^{b}},n^{\prime}\asymp n/l$.
    A.7 Proof of Theorem 4
    Proof  Because $\sigma_{\zeta}=0$, we take largest $l$ which satisfies (5), that is,
    $l^{*}=\frac{1}{C_{\gamma,f}}\frac{n}{\log^{3/2}n},\ \ C_{\gamma,f}:=\frac{R_{0}^{3}C_{f}^{2}}{C_{f}^{\prime}}\max\left(\frac{d^{3/2}\mathrm{len}_{\gamma}}{\sigma_{\gamma}^{4}},\frac{R_{0}^{5}C_{f}^{2}d^{4}}{C_{f}^{\prime 3}\sigma_{\gamma}^{8}}\right)\,.$
    In Lemma 23, we take $p_{1}=1,{s\wedge 1},r=\frac{C_{f}C_{Y}R_{0}}{l}$, and
    $\epsilon_{0}^{2}=(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}(C_{f}C_{Y}R_{0})^{2(s\wedge 1)}C_{\gamma,f}^{s\wedge 1}(\log n)^{3(s\wedge 1)}n^{-2(s\wedge 1)}\,.$
    Then we have zero variance $\operatorname{\mathrm{MSE}}_{(V)}=0$ and can control the following errors:
    $\operatorname{\mathrm{MSE}}_{(NCA)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}\left(\frac{\sigma_{\gamma}^{2}C_{f}^{2}\mathrm{len}_{\gamma}^{2}C_{\gamma,f}^{2}}{\mathrm{reach}_{\gamma}^{2}C_{f}^{\prime 2}}\frac{\log^{3}n}{n^{2}}\right)^{s\wedge 1}$
    $\operatorname{\mathrm{MSE}}_{(\Phi)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}\left(\frac{\mathrm{len}_{\gamma}R_{0}^{2}}{\sigma_{\gamma}^{2}}\right)^{2(s\wedge 1)}(d^{2}\log d)^{s\wedge 1}C_{\gamma,f}^{2(s\wedge 1)}\frac{(\log n)^{7(s\wedge 1)}}{n^{4(s\wedge 1)}}$
    $\operatorname{\mathrm{MSE}}_{(B)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}(C_{f}C_{Y}R_{0})^{2s}C_{\gamma,f}^{2s}\frac{(\log n)^{3s}}{n^{2s}}$
    $\operatorname{\mathrm{MSE}}_{(\Psi)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}(C_{f}C_{Y}R_{0})^{2(s\wedge 1)}C_{\gamma,f}^{s\wedge 1}(\log n)^{3(s\wedge 1)}n^{-2(s\wedge 1)}$
    $\ \ +(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}(C_{f}C_{Y}R_{0})^{\frac{2}{2-s\wedge 1}}\left(R_{0}\right){\sigma_{\gamma}}^{\frac{4}{2-s\wedge 1}}\left(d\log dC_{\gamma,f}\frac{\log^{7/2}n}{n^{2}}\right)^{\frac{1}{1-s\wedge 1}}\,.$
    To disregard those high-order terms, we only need $\frac{n}{(\log n)^{2}}\gtrsim C_{\gamma,f}^{\frac{1}{2}}R_{0}^{2}\sigma_{\gamma}^{2}d\log^{1/2}d$, which is a consequence of the initial assumption $\frac{n}{\log^{3/2}n}\gtrsim\frac{C_{f}\mathrm{len}_{\gamma}}{C_{f}^{\prime}\sigma_{\gamma}}$.
    A.8 Proof of Theorem 3
    Proof  Because $\sigma_{\zeta}>0$, we consider bias-variance trade-off between the bias term $\operatorname{\mathrm{MSE}}_{(B)}$ and the variance term $\operatorname{\mathrm{MSE}}_{(V)}$. It follows that the optimal way is to let product $l^{*}j^{*}$ grow proportional to
    $n^{\frac{1}{2s+1}}M^{*}\text{ where }M^{*}=\left(\sigma_{\zeta}^{-1}(C_{f}C_{Y}R_{0})^{s}(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})\right)^{\frac{2}{2s+1}}$
    In practice, when the number of samples is not quite sufficient, we take $j^{*}=C$ and let $l^{*}$ grow with $n$. On the other hand, when $l^{*}$ already has the magnitude of $l_{upper}=\frac{C_{Y}R_{0}}{\max(\sigma_{\zeta},\omega_{f})}$, we fix $l^{*}=l_{upper}$ and let $j^{*}$ grows with $n$. Notice that $\operatorname{\mathrm{MSE}}_{(NCA)}$ converges to the limit $\frac{\left[f\right]_{\mathcal{C}^{s}}^{2}\sigma_{\gamma}^{2s}C_{f}^{2s}}{\mathrm{reach}_{\gamma}^{2s}}\max(\sigma_{\zeta},\omega_{f})^{2s}$. Also, the optimal $\operatorname{\mathrm{MSE}}_{(B)}+\operatorname{\mathrm{MSE}}_{(V)}$ is
    $\frac{C_{f}^{\frac{2s}{2s+1}}}{C_{f}^{\prime\frac{2s}{2s+1}}}\mathrm{len}_{\gamma}^{\frac{2s}{2s+1}}\sigma_{\zeta}^{{}^{\frac{4s}{2s+1}}}(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})^{\frac{2}{2s+1}}n^{-\frac{2s}{2s+1}}{\log n}$ (8)
    Therefore, we let $\epsilon_{0}^{2}$ to be the maximum of above two errors, that is,
    $\epsilon_{0}^{2}=\max\left(\frac{\left[f\right]_{\mathcal{C}^{s}}^{2}\sigma_{\gamma}^{2s}C_{f}^{2s}\max(\sigma_{\zeta},\omega_{f})^{2s}}{\mathrm{reach}_{\gamma}^{2s}},\frac{C_{f}^{\frac{2s}{2s+1}}\mathrm{len}_{\gamma}^{\frac{2s}{2s+1}}\sigma_{\zeta}^{{}^{\frac{4s}{2s+1}}}(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})^{\frac{2}{2s+1}}\log n}{C_{f}^{\prime\frac{2s}{2s+1}}n^{\frac{2s}{2s+1}}}\right)$
    Moreover, when $n$ is sufficiently large the error denoted by $\epsilon_{0}^{2}$ dominates all other errors: the residual error in the convergence of $\operatorname{\mathrm{MSE}}_{(NCA)}$, as well as the terms $\operatorname{\mathrm{MSE}}_{(\Phi)}$ and $\operatorname{\mathrm{MSE}}_{(\Psi)}$, become negligible high order terms compared with $\epsilon_{0}^{2}$, as soon as the standing assumption $\frac{n}{\log^{{3/2}}n}\gtrsim C_{\gamma,f}\frac{C_{f}\mathrm{len}_{\gamma}}{C_{f}^{\prime}\sigma_{\gamma}}$ is satisfied.
    A.9 Proof of Proposition 13
    Proof  Fix any interval $T\subseteq[0,\mathrm{len}_{\gamma}]$ with size $|T|=C_{f}\max(\sigma_{\zeta},\omega_{f})$, and consider the conditional mean $\mu_{T}:=\mathbb{E}[X\mid t\in T]=\mathbb{E}_{t}[\gamma(t)\mid t\in T]$, and its projection onto the underlying curve $\gamma$ at location $t_{1}=\Pi_{\gamma}\mu_{T}$. Clearly, $\gamma^{\prime}(t_{1})$ is perpendicular to both $\gamma^{\prime\prime}(t_{1})$ and $\mu_{T}-\gamma(t_{1})$. Recall that $\left\|\gamma^{\prime\prime}\right\|_{\infty}\leqslant\mathrm{reach}_{\gamma}^{-1}$ and hence (SC) implies that $C_{f}\max(\sigma_{\zeta},\omega_{f})\left\|\gamma^{\prime\prime}\right\|_{\infty}\leqslant c_{2}$. For any $t,t_{1}\in[0,\mathrm{len}_{\gamma}]$ with $|t-t_{1}|\leqslant C_{f}\max(\sigma_{\zeta},\omega_{f})$, we have
    $\left\|\gamma(t)-\gamma(t_{1})-\gamma^{\prime}(t_{1})(t-t_{1})\right\|\leqslant\frac{1}{2}\left\|\gamma^{\prime\prime}\right\|_{\infty}|t-t_{1}|^{2}\leqslant\frac{1}{2}c_{2}|t-t_{1}|\,.$ (9)
    This implies that the curve is well-approximated by a straight line on the whole interval $T$, implies that $t_{1}\in T$, and, combined with the minimizing property of $\Pi_{\gamma}$, yields
    $\left\|\gamma(t_{1})-\mu_{T}\right\|\leqslant\left\|\gamma(\mathbb{E}[t|t\in T])-\mu_{T}\right\|\leqslant\frac{1}{2}c_{2}|T|\,.$
    We have
    $\mathbb{E}[(X-\mu_{T})(X-\mu_{T})^{\intercal}\mid t\in T]$
    $=$ $(\gamma(t_{1})-\mu_{T})(\gamma(t_{1})-\mu_{T})^{\intercal}+\gamma^{\prime}(t_{1})\gamma^{\prime}(t_{1})^{\intercal}\mathbb{E}[|t-t_{1}|^{2}|t\in T]+(\gamma(t_{1})-\mu_{T})\gamma^{\prime}(t_{1})^{\intercal}\mathbb{E}[t-t_{1}|t\in T]+U\,,$
    with $\left\|U\right\|\leqslant\left\|\gamma(t_{1})-\mu_{T}\right\|\left\|\gamma^{\prime\prime}\right\|\mathbb{E}[(t-t_{1})^{2}|t\in T]+\left\|\gamma^{\prime\prime}\right\|^{2}\mathbb{E}[(t-t_{1})^{4}|t\in T]\leqslant 2c_{2}^{2}|T|^{2}$. It follows that the conditional covariance along $\gamma^{\prime}(t_{1})$ has a lower bound
    $\gamma^{\prime}(t_{1})^{\intercal}\Sigma_{T}\gamma^{\prime}(t_{1})=\mathbb{E}[|t-t_{1}|^{2}|t\in T]+\gamma^{\prime}(t_{1})^{\intercal}U\gamma^{\prime}(t_{1})\gtrsim(\frac{1}{4}-2c_{2}^{2})|T|^{2}\,,$
    which is positive for $c_{2}$ small enough, and the conditional covariance along any direction $\nu$ that is orthogonal to $\gamma^{\prime}(t_{1})$ has an upper bound
    $\nu^{\intercal}\Sigma_{T}\nu\leqslant\left\|\gamma(t_{1})-\mu_{T}\right\|^{2}+2c_{2}^{2}|T|^{2}+\sigma_{\gamma}^{2}\leqslant(3c_{2}^{2}+c_{1}^{2})|T|^{2}\,.$
    Therefore the largest eigenvalue of $\Sigma_{T}$ is significantly larger than others, as long as constants $c_{1},c_{2}$ are sufficiently small, i.e. we are in the “wide” slice scenario. Moreover, the largest principal component of $\Sigma_{T}$ is roughly tangential to the curve.
    A.10 Proof of Proposition 14 and Theorem 5
    Proof  Similar to Proposition 17, we have the following high probability bound on the estimation error of parameters such as $\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle$, $\left\|{\widehat{v}_{l,h}^{b}}-{v_{l,h}^{b}}\right\|$, $\left|{\lambda_{1}(\Sigma_{l,h}^{b})}-{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}\right|$. The argument is the same, so the proof is omitted.
    Proposition 18 (local NVM for “wide” slice)
    Suppose (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (SC), and ($\bm{\omega_{f}}$) hold true. Let ${\mu_{l,h}^{b}}$ be the mean of $h$-th slice and ${v_{l,h}^{b}}$ be the significant vector of $h$-th slice. Then, for every $l$ such that $|{R_{l,h}^{b}}|\asymp\max(\sigma_{\zeta},\omega_{f})$ for all $h\in\mathcal{H}_{l}^{b}$, for every $\epsilon\in(0,1)$ and $\tau\geqslant 1$, on each slice
    1. (a)
      For any $h\in\mathcal{H}_{l}$ and any $\epsilon>0$, the estimation error of the slice mean along the tangential direction can be bounded as
      $\mathbb{P}\left\{\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|>\frac{C_{f}^{2}|{R_{l,h}^{b}}|^{2}\epsilon}{R_{0}\sqrt{d}}\ \bigg{|}\ {n_{l,h}^{b}}\right\}\lesssim d\exp\left(-\frac{cC_{f}^{2}|{R_{l,h}^{b}}|^{2}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{2}d+\epsilon C_{f}|{R_{l,h}^{b}}|R_{0}\sqrt{d}}\right)+n^{-\tau}$
    2. (b)
      For any $h\in\mathcal{H}_{l}$, the estimation error of the significant vector can be bounded as,
      $\mathbb{P}\left\{\left\|{{\widehat{v}_{l,h}^{b}}}-{{v_{l,h}^{b}}}\right\|>\epsilon\ \big{|}\ {n_{l,h}^{b}}\right\}\lesssim d\exp\left(-\frac{cC_{f}^{2}|{R_{l,h}^{b}}|^{2}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{2}d+\epsilon C_{f}|{R_{l,h}^{b}}|R_{0}\sqrt{d}}\right)+d\exp\left(-\frac{cC_{f}^{4}|{R_{l,h}}|^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)+n^{-\tau}$
    3. (c)
      For any $h\in\mathcal{H}_{l}$ and any $10<u<\frac{R_{0}\sqrt{d}}{C_{f}|{R_{l,h}^{b}}|}<$ 1, the estimation error of the first principal value of the slice can be bounded as,
      $\mathbb{P}\left(\left|{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{1}(\Sigma_{l,h}^{b})}\right|>u^{2}C_{f}^{2}|{R_{l,h}^{b}}|^{2}\ \bigg{|}\ {n_{l,h}^{b}}\right)$
      $\qquad\qquad\lesssim d\exp\left(-c\frac{u^{2}C_{f}^{4}|{R_{l,h}^{b}}|^{4}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{4}d^{2}+uC_{f}^{2}|{R_{l,h}^{b}}|^{2}R_{0}^{2}d}\right)+d\exp\left(-\frac{cC_{f}^{4}|{R_{l,h}}|^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)+n^{-\tau}$
    Moreover, we have the following probability bound For fixed constant $\beta<\frac{R_{0}^{2}d}{C_{f}^{2}|{R_{l,h}}|^{2}}$,
    $\mathbb{P}\left(\max\left(\left\|{\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|,\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|^{2}\right)\gtrsim\beta C_{f}^{2}|{R_{l,h}^{b}}|^{2}\right)\lesssim d\exp\left(-\frac{c\beta^{2}C_{f}^{4}|{R_{l,h}^{b}}|^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)$
    Consider the event $\bm{\mathfrak{S}}$ that we have a small estimation error for information in all slices:
    $\bm{\mathfrak{S}}(\epsilon,\beta,u)=\left\{\begin{split}&\text{ For all }h\in{\mathcal{H}_{l}^{b}},\left|\langle{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}},{v_{l,h}^{b}}\rangle\right|\lesssim\frac{\epsilon|{R_{l,h}^{b}}|^{2}\epsilon}{R_{0}\sqrt{d}},\ \left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|\lesssim\frac{R_{0}\sqrt{d}}{2},\left\|{v_{l,h}^{b}}-{\widehat{v}_{l,h}^{b}}\right\|\lesssim\epsilon\\ &\left|{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{1}(\Sigma_{l,h}^{b})}\right|\lesssim u^{2}C_{f}^{2}|{R_{l,h}^{b}}|^{2},\max\left(\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|^{2},\left\|{\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|\right)\lesssim\beta^{2}C_{f}^{2}|{R_{l,h}^{b}}|^{2}\end{split}\right\}$
    We claim that the conditions $\epsilon^{\prime}\asymp\beta^{\prime}\asymp\frac{C_{f}^{2}|{R_{l,h}^{b}}|^{2}}{R_{0}^{2}d}$ and $u\asymp\frac{C_{f}|{R_{l,h}^{b}}|}{R_{0}\sqrt{d}}$ are sufficient to perform almost correct classification.
    Recall that similar to the “thin” slice scenario, in order to obtain correct classification, we want the estimation error of the distance function to be small, such that for all $|h^{\prime}-{h_{x}}|\geqslant 2$,
    $\left|{\widehat{\mathrm{dist}}}(x,h^{\prime})-{\mathrm{dist}}(x,h^{\prime})\right|+\left|{\widehat{\mathrm{dist}}}(x,{h_{x}})-{\mathrm{dist}}(x,{h_{x}})\right|<\left|{\mathrm{dist}}(x,h^{\prime})-{\mathrm{dist}}(x,{h_{x}})\right|$
    Indeed, this will imply that ${\widehat{h}_{x}}=\operatorname*{{arg\,min}}_{h^{\prime}\in{\mathcal{H}_{l}^{b}}}{\widehat{\mathrm{dist}}}(x,h^{\prime})$ is the correct or adjacent classification, i.e. $|{h_{x}}-{\widehat{h}_{x}}|\leqslant 1$.
    Now, we are going to analyze the distance function. Notice that in the “wide” slice scenario, the difference in distance can be bounded as follows:
    $\sqrt{{\mathrm{dist}}(x,{h_{x}}+k)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim C_{f}|{R_{l,h}^{b}}|\ \ \text{ for any }|k|\geqslant 2$
    (Recall that for large $k$, we can bound this by $\mathrm{reach}_{\gamma}\gtrsim C_{f}|{R_{l,h}^{b}}|$ by assumption (SC).) Thus this gives a lower bound for the right-hand side. We are going to control the estimation error on the left-hand side. We use the same argument in the proof of Proposition 11. Given event $\bm{\mathfrak{S}}(\epsilon,\beta,u)$,
    $\left|{\widehat{\mathrm{dist}}}(x,h)-{\mathrm{dist}}(x,h)\right|$
    $\lesssim$ $C_{f}^{2}|{R_{l,h}^{b}}|^{2}\beta^{2}+C_{f}|{R_{l,h}^{b}}|\beta\sqrt{{\mathrm{dist}}(x,h)}+u^{2}R_{0}^{2}d+\epsilon^{2}R_{0}^{2}d+C_{f}^{2}|{R_{l,h}^{b}}|^{2}\beta^{2}\epsilon^{2}+\frac{C_{f}^{4}|{R_{l,h}^{b}}|^{4}}{R_{0}^{2}d}\epsilon^{2}\lesssim C_{f}^{2}|{R_{l,h}^{b}}|^{2}$
    when $\epsilon^{\prime}\asymp\beta^{\prime}\asymp\frac{C_{f}^{2}|{R_{l,h}^{b}}|^{2}}{R_{0}^{2}d}$ and $u\asymp\frac{C_{f}|{R_{l,h}^{b}}|}{R_{0}\sqrt{d}}$. Therefore, we derive the conclusion that the event of misclassification by at least two slices has a small probability:
    $\mathbb{P}\left(\left|{\widehat{h}_{x}}-{h_{x}}\right|\geqslant 2\right)\lesssim ld\exp\left(-c\frac{C_{f}^{6}\max(\sigma_{\zeta},\omega_{f})^{7}n}{C_{Y}R_{0}^{7}d^{3}\sqrt{\log n}}\right)+ln^{-\tau}$
    This finishes the proof of Proposition 14. Theorem 5 follows easily because we mainly utilize the high accuracy of classification without performing one-dimensional regression.
    Appendix B Technical Results
    Lemma 19
    Let $X$ be a random variable, and let $X_{1},\dots,X_{n}$ be independent copies of $X$. Given a measurable set $E$, define $\rho(E)=\mathbb{P}\left\{X\in E\right\}$, and $\hat{\rho}(E)=n^{-1}\sum_{i}\mathbbm{1}\left\{X_{i}\in E\right\}$. Then
    $\mathbb{P}\left\{|\hat{\rho}(E)-\rho(E)|>t\right\}\leqslant 2\exp\left\{-\frac{nt^{2}/2}{\rho(E)+t/3}\right\}$
    In particular, for $t=\rho(E)/2$, we have
    $\mathbb{P}\left\{\hat{\rho}(E)\not\in\left[\frac{1}{2}\rho(E),\frac{3}{2}\rho(E)\right]\right\}\leqslant\mathbb{P}\left\{|\hat{\rho}(E)-\rho(E)|>\frac{1}{2}\rho(E)\right\}\leqslant 2\exp(-\frac{3}{28}n\rho(E))$
    Lemma 20
    Let $X\in\mathbb{R}^{d}$ be a sub-Gaussian vector with variance proxy $R_{0}^{2}$. Then for any $t>0$, we have $\mathbb{P}\left\{\left\|X\right\|>t\right\}\leqslant 2\exp\left(-\frac{t^{2}}{2dR_{0}^{2}}\right)$.
    Lemma 21
    Let $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$ be independent copies of a sub-Gaussian pair $(X,Y)\in\mathbb{R}^{d+1}$ with variance proxy $R_{0}^{2}$. Then, for every $\gamma\in(0,1)$ and every $a_{X},a_{Y}>0$, we have
    {#$\mathbb{P}\left\{\#\{(X_{i},Y_{i})\in B(0,a_{X}\sqrt{d}R_{0})\times[-a_{Y}R_{0},a_{Y}R_{0}]\}<\delta_{X,Y}\gamma n\right\}\leqslant 2\exp\left(-\frac{\delta_{X,Y}(1-\gamma)^{2}/2}{1+(1-\gamma)/3}n\right)\,,$
    where $\delta_{X,Y}=\delta_{X}\delta_{Y|X}$, $\delta_{X}=1-2\exp\left(-a_{X}^{2}/2\right)$, and $\delta_{Y|X}=1-2\exp\left(-a_{Y}^{2}\delta_{X}/2\right)$.
    Proof  See (Lanteri et al., 2022, Lemma B.3).
    Lemma 22
    Let $\rho$ be a probability distribution in $\mathbb{R}^{d}$, $v,w\in\mathbb{S}^{d-1}$, $\operatorname{\mathrm{Proj}}_{u}x=\langle u,x\rangle$ and $I\subseteq\mathbb{R}$ an interval with $\rho(\operatorname{\mathrm{Proj}}_{u}^{-1}I)>0$ for $u\in\{v,w\}$. Suppose $\rho$ has an upper bounded density and $\rho(\operatorname{\mathrm{Proj}}_{v}^{-1}I)\gtrsim|I|$, and let $\rho_{v}$ denote the pushforward measure under the map $\operatorname{\mathrm{Proj}}_{v}$. Then
    $W_{1}\left(\rho_{v}(x|\operatorname{\mathrm{Proj}}_{v}x\in I),\rho_{v}(x|\operatorname{\mathrm{Proj}}_{w}x\in I)\right)\lesssim\sin(\angle(v,w))\operatorname{\mathrm{diam}}(\operatorname{\mathrm{supp}}\rho)\,.$
    Lemma 23 ((Lanteri et al., 2022, Proposition 4 ))
    Assume (X subG), ($\bm{\zeta}$ subG), and (P). Suppose function $\widetilde{f}\in\mathcal{C}^{s}$ with $s\in[\frac{1}{2},1]$ is defined on interval $I$ with diameter $r\geqslant 1$. Let integer $m=\lfloor s\rfloor$ be the largest integer smaller than or equal to $s$, so $m=0$ if $s\in[\frac{1}{2},1)$ and $m=1$ if $s=1$. Given $j\geqslant 1$, construct $\{I_{j,k}\}_{k=1}^{j}$ the uniform partition of $I$ into $j$ intervals. Let ${\widehat{v}}\in\mathbb{S}^{d-1}$ be an estimator of $v\in\mathbb{S}^{d-1}$ such that $\langle{\widehat{v}},v\rangle\geqslant 0$. Given data $\{(X_{i}^{\prime},Y_{i}^{\prime})\}_{i=1}^{n^{\prime}}$. For $u\in\{v,{\widehat{v}}\}$, denote $n_{j,k|u}=\#\{X_{i}^{\prime}:\langle u,X_{i}^{\prime}\rangle\in I_{j,k}\}$. For each $k\in\mathcal{K}_{j|u}:=\{k:n_{j,k|u}\geqslant n^{\prime}/j\}$, compute
    ${\widehat{f}_{j,k|u}}=\operatorname*{{arg\,min}}_{\deg(p)\leqslant m}\sum_{i\leqslant n^{\prime}}\left|Y_{i}-p(\langle u,X_{i}^{\prime}\rangle)\right|^{2}\mathbbm{1}_{I_{j,k}}(\langle u,X_{i}^{\prime}\rangle)$
    and compute the piecewise constant $(s<1)$ or linear $(s=1)$ estimator of $\widetilde{f}$ at scale $j$ conditioned on $u$ and truncated at $\left|\widetilde{f}\right|_{L^{\infty}}$
    ${\widehat{f}_{j|u}}(r)=\sum_{k\in\mathcal{K}_{j}}{\widehat{f}_{j,k|u}}(r)\mathbbm{1}_{I_{j,k}}(r)$
    Conditioned on $\left\|{\widehat{v}}-v\right\|\leqslant t/j$ for some $t\geqslant 1$. For every $\epsilon_{0}>0$, we have
    $\mathbb{E}_{X}\left[\left|{\widehat{f}_{j,k|v}}(\langle v,X\rangle)-{\widehat{f}_{j,k|{\widehat{v}}}}(\langle{\widehat{v}},X\rangle)\right|^{2}\mathbbm{1}_{B(0,r)}(X)\right]\lesssim\left[\widetilde{f}\right]_{\mathcal{C}^{s}}^{2}t^{2}r^{\frac{2}{2-s}}\left\|{\widehat{v}}-v\right\|^{\frac{2}{2-s}}+\epsilon_{0}^{2}$
    with probability higher than $1-C\#\mathcal{K}_{j}\exp\left(-\frac{cn^{\prime}\epsilon_{0}^{2}}{\#\mathcal{K}_{j}\left[\widetilde{f}\right]_{\mathcal{C}^{s}}^{2}t^{2}r^{2s}}\right)$
    Appendix C Case Analysis: Meyer helix
    C.0.1 Background: standard & modified Meyer’s staircase
    We consider the standard Y. Meyer’s staircase. Fix constant $\delta\geqslant 1$. Consider the unit interval $I=[0,1]$ and the set of Gaussians $\mathcal{N}(t;\mu,\delta^{2})$ where the mean $\mu$ takes values in $I$, and the density function is truncated to accept arguments $t\in I$ only. Varying $\mu\in I$ in this manner induces a smooth embedding of the interval $I$ into the infinite-dimensional Hilbert space $L^{2}(I)$, i.e. a curve. Explicitly, we take the square root of the Gaussian density centered at $\mu\in I$ and truncate it to $t\in I$:
    $\begin{split}I\to L^{2}(I)\quad:\quad\mu\mapsto g_{\mu}(t):=\frac{1}{\sqrt[4]{2\pi}\sqrt{\delta}}\exp\left(-\frac{|t-\mu|^{2}}{4\delta^{2}}\right)\end{split}$ (10)
    By discretizing $I$, we may sample this manifold and project it into a finite-dimensional space. In particular, for any $d\in\mathbb{N}$, a grid $\Gamma_{d}\subseteq I$ of $d$ points may be generated. It is obtained by subdividing $I$ in $d$ equal parts and thus $\Gamma_{d}(k)=k/d$ for $k=1,\dots,d$. Explicitly, the evaluation function is
    $\begin{split}L^{2}(I)\to\mathbb{R}^{d}\quad:\quad g_{\mu}(t)\mapsto(g_{\mu}(1/d),\dots,g_{\mu}(1))^{\intercal}\end{split}$ (11)
    Thus, combining the above two maps (10) and (11) together produces an embedding of interval $I$ into $\mathbb{R}^{d}$ (which is equivalent to a curve in $\mathbb{R}^{d}$). We write it explicitly as $x(t)=(x_{1}(t),\dots,x_{d}(t))^{\intercal}$ where $t\in I$ and for each $k=1,\dots,d$. For the standard Meyer’s staircase, it follows that the expression for $x_{k}(t)$ is
    $x_{k}(t)=\frac{1}{\sqrt[4]{2\pi}\sqrt{\delta}}\exp\left(-\frac{|k/d-t|^{2}}{4\delta^{2}}\right)$ (12)
    Note that this expression differs from Definition 1 because it does not use the unit-speed parameterization. However, this expression has two advantages: first, it is uniform over dimension $d$, and there is no need to worry about different lengths; second, it strengthens the fact that these curves are finite-dimensional approximations of the function $g_{\mu}(t)\in L^{2}(I)$. In numerical simulations, we can clearly convert it to the unit-speed parameterization.
    Notice that the map (10) describes a curve in $L^{2}(I)$. Here $\mu\in I$ parameterizes this curve while $t\in I$ is merely the argument for the function $g_{\mu}$. On the other hand, equation (12) describes a curve in $\mathbb{R}^{d}$. Here $t\in I$ parameterizes the curve while $\mu$ is replaced by a discrete grid $\Gamma_{d}$.
    The standard Meyer’s staircase is an interesting example because it allows us to construct a curve in high dimensional Euclidean space $\mathbb{R}^{d}$. However, because of the construction, as dimension $d\to\infty$, the standard Meyer’s staircase defined in equation (12) will converge to a limit that corresponds to the function $g_{\mu}\in L^{2}(I)$. This implies that the complexity of the curve is bounded as $d\to\infty$. Because we are focusing on the regression problem for general curve classes, we need to consider various curves with different complexity and different ambient dimensions. Our strategy is to consider analogies of (12).
    One direct modification of the standard Meyer’s staircase is to let $\delta=\frac{1}{d}$ in equation 12. This modified Meyer-staircase is an example of a collection of curves whose complexity grows with dimension $d$. Besides parameters such as length, diameter, curvature, and reach, we also consider the effective linear dimension. One way to measure the effective linear dimension is to study the singular values of the curve. Suppose we perform Singular Value Decomposition on a curve in $\mathbb{R}^{d}$ and obtain its singular value $\lambda_{\gamma}(k),k=1,\dots,d$ in descending order. Then, we can consider the sum of the singular values divided by the largest singular value, $\left\|\lambda_{\gamma}\right\|_{1}:=\sum_{k=1}^{d}\lambda_{\gamma}(k)/\lambda_{\gamma}(1)$, or count the number of singular values that are greater than $0.05$ times the largest singular value, $\left\|\lambda_{\gamma}\right\|_{0}:=\#\{\lambda_{\gamma}(k):\lambda_{\gamma}(k)>0.05\lambda_{\gamma}(1)\}$. Both $\left\|\lambda_{\gamma}\right\|_{1}$ and $\left\|\lambda_{\gamma}\right\|_{0}$ are scaling invariant and measure the minimal number of linear dimensions needed to capture (in the mean squared sense) the underlying curve $\gamma$ up to a given relative error. This quantity is commonly used as a stable version of rank for a matrix (sometimes called numerical, or stable, rank).
    Figure 9 illustrates the relationship between the modified Meyer-staircase parameters and the dimension $d$. One can observe that the length is roughly proportional to $d^{1.5}$, diameter is roughly $d^{0.5}$, curvature is roughly proportional to $d^{-0.5}$, and reach is roughly $d^{0.5}$. It turns out that both $\left\|\lambda_{\gamma}\right\|_{1}$ and $\left\|\lambda_{\gamma}\right\|_{0}$ are roughly proportional to $d^{1}$. In this sense, the standard Meyer-staircase has its complexity growing with $d$.
    Figure 9: Behavior of geometric features of the modified Meyer staircase in $\mathbb{R}^{d}$ as a function of $d$.
    However, the modified Meyer-staircase is still special in the following two respects: (i) It approximately stays on the sphere $\sqrt{d}\mathbb{S}^{d-1}$: For $t\in(0,1)$,
[…]
    (ii) the local reach of modified Meyer’s staircase is almost the reciprocal of the magnitude of local curvature. The above two aspects indicate that the curve traverses the space with weak self-entanglement. Hence, it suggests the possibility of finding a linear projection $P:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ with $d^{\prime}$ much smaller than $d$ such that the projected image $P\gamma$ is a much simpler curve. For example, suppose we perform the linear projection of modified Meyer’s staircase onto its first few principal components: the projected image is a simple curve, and the learning problem can be significantly simplified if we study the regression problem on the projected curve. Hence, to test the performance algorithm for the regression problem, we aim to test curves that are so complex that there is no trivial dimension reduction, e.g. via standard techniques such as Principal Component Analysis.
    C.0.2 Meyer helix
    To summarize, we want to perform numerical tests of learning problems on curves that are complex enough. Here are some characteristics of the complexity of the curve: (i) parameters such that length, diameter, reach, and effective linear dimension $\left\|\lambda\right\|_{1},\left\|\lambda\right\|_{0}$ grow with the dimension $d$; (ii) the curve $\gamma$ has no trivial dimension reduction. In particular, consider linear projection $P_{d^{\prime}}:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ such as projection onto the first $d^{\prime}\leqslant d$ principal components. Define the “regression complexity”
    $C_{\gamma}:=\frac{\mathrm{len}_{\gamma}}{\mathrm{reach}_{\gamma}}$ (13)
    of a curve as its length divided by its reach. We consider curves complex enough such that whenever the projected curve $P_{d^{\prime}}\gamma$ has regression complexity $C_{P_{d^{\prime}}\gamma}\lesssim C_{\gamma}$ then it implies that the dimension $d^{\prime}$ cannot be small, for example $d^{\prime}\gtrsim d$.
    Because the regression problem should be scaling invariant, we can freely rescale the curve, and we choose the normalization such that the reach equals $\sqrt{d}$. This is consistent with Assumption ($\bm{\gamma_{1}}$) and allows us to take $\sigma_{\gamma}$, the deviation of data $X$ away from the curve, with order 1. In particular, we do not let curvature grow with $d$ here.
    We introduce the following curve, called Meyer helix, as an analogy of the standard and modified Meyer’s staircase:
    $x_{k}(t)=\frac{1}{\sqrt[4]{2\pi}\sqrt{\delta_{d}}}\cos\left(ak+\frac{t-k/d}{\delta_{d}^{\prime}}\right)G\left(\frac{\left|k/d-t\right|}{\delta_{d}}\right)$ (14)
    where $\delta_{d}=(1+0.3\cos(ak))/d\ \text{ and }\ \delta_{d}^{\prime}=(1+0.3\sin(ak))/d$, with constant $a=10$ and function $G$ is taken to be Bernstein-type decay $G(z)=\exp\left(-\frac{z^{2}}{1+z}\right)$. Compared with the standard and modified Meyer-staircase, this curve has an extra factor of cosine term. The effect of this cosine term is to facilitate the traversing of point $x(t)$ around the space $\mathbb{R}^{d}$ and introduce more self-entanglement. Moreover, the following values vary from one axis to another: the frequency $1/\delta_{d}$ in function $G$, the frequency $1/\delta^{\prime}$ in cosine term, and the phase $ak$ in cosine term. That variation makes the curve less special while keeping the desired complexity.
    In Figure 10, we plot the parameters of Meyer helix after scaling the reach to be $\sqrt{d}$. Similar to the modified Meyer’s staircase, we see that the length is roughly proportional to $d^{1.5}$, the diameter is roughly proportional to $d^{0.5}$, the curvature is roughly proportional to $d^{-0.5}$, the reach is roughly proportional to $d^{0.5}$, and effective linear dimension $\left\|\lambda_{\gamma}\right\|_{1}$, $\left\|\lambda_{\gamma}\right\|_{0}$ is roughly proportional to $d$. Moreover, we can adopt the regression complexity (13) to measure the effective linear dimension: define $d_{SVD}$ to be the smallest $d^{\prime}$ such that $C_{P_{d^{\prime}}\gamma}\leqslant 1.2C_{\gamma}$, then this effective linear dimension is roughly proportional to $d^{1}$.
    Figure 10: Behavior of geometric features of the Meyer helix in $\mathbb{R}^{d}$ as a function of $d$.
    Moreover, the above properties of complexity are pretty robust: we can also choose Gaussian-type to decay $G(z)=\exp(-z^{2})$ or choose another constant in the variants (e.g.,, $a,\delta_{d}$, and $\delta_{d}^{\prime}$) and the above properties persist. This indicates that this collection of Meyer helix curves indeed has its complexity growing with ambient dimension $d$. Numerical tests suggest that this collection of curves does not have simple dimension reduction via random projections that preserves geometric properties such that length and reach. Recall the following form of Johnson-Lindenstrauss random projection lemma for manifolds:
    Lemma 24 ((Baraniuk and Wakin, 2009, Theorem 3.1) )
    Let $\mathcal{M}$ be a compact $K$-dimensional submanifold of $\mathbb{R}^{N}$ having condition number $1/\tau$, volume $V$, and geodesic covering regularity $R$. Fix $10<\epsilon<$ 1 and $10<\rho<$ 1. Let $\Phi$ be a random orthoprojector from $\mathbb{R}^{N}$ to $\mathbb{R}^{M}$ with $M\gtrsim\epsilon^{-2}K\log(NVR\tau^{-1}\epsilon^{-1})\log(1/\rho)$. If $M\leqslant N$, then with probability at least $1-\rho$, the following statement holds: for every pair of points $x,y\in\mathcal{M}$,
    $(1-\epsilon)\sqrt{\frac{M}{N}}\leqslant\frac{\left\|\Phi x-\Phi y\right\|_{2}}{\left\|x-y\right\|_{2}}\leqslant(1+\epsilon)\sqrt{\frac{M}{N}}$
    If we apply this Johnson-Lindenstrauss lemma to a curve $\gamma$ on $\mathbb{R}^{d}$ with length $\mathrm{len}_{\gamma}$ and reach $\mathrm{reach}_{\gamma}$, then we can take ambient dimension $N=d$, the condition number $1/\tau=1/\mathrm{reach}_{\gamma}$, volume $V=\mathrm{len}_{\gamma}$, and geodesic covering regularity $R=\mathcal{O}(1)$, and thus $M=\mathcal{O}\left(\epsilon^{-2}\log(1/\rho)\log\left(\frac{d}{\epsilon}\frac{\mathrm{len}_{\gamma}}{\mathrm{reach}_{\gamma}}\right)\right)$, which suggests the possibility of dimension reduction for the curve via linear projection. However, in our context what matters more is whether the complexity of the curve is simplified, for example $\mathrm{len}_{\gamma}/\mathrm{reach}_{\gamma}$; furthermore, our samples are not distributed on the curve, but in a tube around the curve of radius as large as a fraction of the reach. Here are numerical tests in the same setup as Figure 1: for the Meyer helix in $d=36$ dimensions, we consider the projection onto a 12-dimensional subspace, obtained by PCA or by random projection, and compute the length and the reach of the image. We consider 10 independent repetitions and, for comparison, we also include the original curve as well as the projection onto the first 12 principal components. To be consistent with the scaling in the Johnson-Lindenstrauss lemma, for the PCA and random projection, we rescale points on the projected image of curve by the factor $\sqrt{36/12}=\sqrt{3}$.
    projection $P$ original $\gamma$ PCA 1 2 3 4 5 6 7 8 9 10
    $\mathrm{len}_{P\gamma}$ 731 1136 703 720 709 717 639 748 766 709 737 706
    $\mathrm{reach}_{P\gamma}$ 6.0 5.0 2.2 4.4 2.4 3.9 2.9 4.4 2.0 4.3 2.5 2.3
    $\frac{\mathrm{len}_{P\gamma}}{\mathrm{reach}_{P\gamma}}$ 122 226 322 164 295 187 223 171 383 166 291 313
    These numerical results support our argument that the Meyer helix cannot be easily embedded in lower dimensional space without significantly affecting its complexity, which involves pointwise curvature/reach that are beyond the scope of random projections.

### v2

@@
 Abstract
-Several statistical models for regression of a function $F$ on $\mathbb{R}^{d}$ without the statistical and computational curse of dimensionality exist, for example by imposing and exploiting geometric assumptions on the distribution of the data (e.g.
-that its support is low-dimensional), or strong smoothness assumptions on $F$, or a special structure $F$.
-Among the latter, compositional models assume $F=f\circ g$ with $g$ mapping to $\mathbb{R}^{r}$ with $r\ll d$, have been studied, and include classical single- and multi-index models and recent works on neural networks.
-While the case where $g$ is linear is rather well-understood, much less is known when $g$ is nonlinear, and in particular for which $g$’s the curse of dimensionality in estimating $F$, or both $f$ and $g$, may be circumvented.
-In this paper, we consider a model $F(X):=f(\Pi_{\gamma}X)$ where $\Pi_{\gamma}:\mathbb{R}^{d}\to[0,\textrm{len}_{\gamma}]$ is the closest-point projection onto the parameter of a regular curve $\gamma:[0,\textrm{len}_{\gamma}]\to\mathbb{R}^{d}$ and $f:[0,\textrm{len}_{\gamma}]\to\mathbb{R}^{1}$.
-The input data $X$ is not low-dimensional, far from $\gamma$, conditioned on $\Pi_{\gamma}(X)$ being well-defined.
-The distribution of the data, $\gamma$ and $f$ are unknown.
-This model is a natural nonlinear generalization of the single-index model, which corresponds to $\gamma$ being a line.
-We propose a nonparametric estimator, based on conditional regression, and show that under suitable assumptions, the strongest of which being that $f$ is coarsely monotone, it can achieve the one-dimensional optimal min-max rate for non-parametric regression, up to the level of noise in the observations, and be constructed in time $\mathcal{O}(d^{2}n\log n)$.
+Regressing a function $F$ on $\mathbb{R}^{d}$ without the statistical and computational curse of dimensionality requires special statistical models, for example that impose geometric assumptions on the distribution of the data (e.g., that its support is low-dimensional), or strong smoothness assumptions on $F$, or a special structure $F$.
+Among the latter, compositional models $F=f\circ g$ with $g$ mapping to $\mathbb{R}^{r}$ with $r\ll d$ include classical single- and multi-index models, as well as neural networks.
+While the case where $g$ is linear is well-understood, less is known when $g$ is nonlinear, and in particular for which $g$’s the curse of dimensionality in estimating $F$, or both $f$ and $g$, may be circumvented.
+Here we consider a model $F(X):=f(\Pi_{\gamma}X)$ where $\Pi_{\gamma}:\mathbb{R}^{d}\to[0,\textrm{len}_{\gamma}]$ is the closest-point projection onto the parameter of a regular curve $\gamma:[0,\textrm{len}_{\gamma}]\to\mathbb{R}^{d}$, and $f:[0,\textrm{len}_{\gamma}]\to\mathbb{R}^{1}$.
+The input data $X$ is not low-dimensional: it can be as far from $\gamma$ as the condition that $\Pi_{\gamma}(X)$ is well-defined allows.
+The distribution $X$, the curve $\gamma$ and the function $f$ are all unknown.
+This model is a natural nonlinear generalization of the single-index model, corresponding to $\gamma$ being a line.
+We propose a nonparametric estimator, based on conditional regression, that under suitable assumptions, the strongest of which being that $f$ is coarsely monotone, achieves, up to log factors, the one-dimensional optimal min-max rate for non-parametric regression, up to the level of noise in the observations, and be constructed in time $\mathcal{O}(d^{2}n\log n)$.
 All the constants in the learning bounds, in the minimal number of samples required for our bounds to hold, and in the computational complexity are at most low-order polynomials in $d$.
-Keywords: Nonparametric regression; compositional models; single-index model.
+Keywords: High-dimensional regression; nonparametric regression; compositional models; single-index model.
 1 Introduction
@@
 $Y_{i}=F(X_{i})+\zeta_{i}\,.$
-In the general nonparametric and distribution-free setup where we only know that $F\in\mathcal{C}^{s}(\mathbb{R}^{d})$ is Hölder continuous with exponent $s>0$ and, say, compacted supported, the min-max nonparametric rate for estimating $F$ in $L^{2}(\rho)$ is $n^{-\frac{s}{2s+d}}$, see (Binev et al., 2023; Györfi et al., 2002) and references therein.
+In the general nonparametric and distribution-free setup where we only know that $F\in\mathcal{C}^{s}(\mathbb{R}^{d})$ is Hölder continuous with exponent $s>0$ and, say, compactly supported, the min-max nonparametric rate for estimating $F$ in $L^{2}(\rho_{X})$ is $n^{-\frac{s}{2s+d}}$, see (Binev et al., 2023; Györfi et al., 2002) and references therein.
 Kernel estimators attain this learning rate with properly chosen bandwidth and kernel, as so do a variety of other well-understood estimators based on Fourier or multiscale decompositions, see (Györfi et al., 2002; Binev et al., 2005, 2007) and the numerous references therein.
@@
 As no estimator can achieve a faster learning rate, in the min-max sense, for all such functions, and yet in many applications $d$ is very large, it is of interest to consider special classes of regression problems where the curse of dimensionality may be avoided.
+In this work, we introduce and construct efficient estimators for the model where $F(x)=f(\Pi_{\gamma}x)$, where $\gamma$ is a curve in $\mathbb{R}^{d}$, and $\Pi_{\gamma}$ maps $x\in\mathbb{R}^{d}$ to the closest point to $x$ that belongs to $\gamma$.
+The input data $x$ is sampled from a distribution supported around $\gamma$, and it is therefore not low-dimensional.
+With the curve $\gamma$, the distribution of the input data, and the function $f$ all unknown, we construct an estimator for $F$, using conditional inverse regression techniques, that converges at the near min-max one-dimensional rate for regression (up to log factors), and that can constructed in $\mathcal{O}(d^{2}n\log n)$, where $n$ is the number of samples.
+It therefore avoids both the statistical and the computational curse of dimensionality.
+The model is detailed in the next section, together with an informal statement of the main theorem.
+Before we delve into details, we provide some context and motivation for this work, and in particular discuss statistical models for regression that have been introduced and studied in order to avoid the curse of dimensionality.
 Intrinsically low-dimensional models.
 In one direction, one may make geometric assumptions about the distribution $\rho_{X}$ of the input data $X_{i}$, for example that $\rho_{X}$ is supported on a low-dimensional manifold $\mathcal{M}$, say of dimension $r\ll d$, while $F$ is, say, a generic $s$-Hölder function on $\mathcal{M}$.
-In these settings, estimators exist that converge at an optimal rate concerning the intrinsic dimension $r$, see (Bickel and Li, 2007; Kpotufe, 2011; Kpotufe and Garg, 2013; Liao et al., 2016, 2022; Liu et al., 2024a), with at least some of them with associated fast algorithms, for example the estimators in (Liao et al., 2016, 2022) can be constructed in time $\mathcal{O}(C^{r}d^{s}n\log n)$ and they achieve, adaptively, min-max optimal rate on large function spaces with unknown regularity possibly varying across locations and scales.
+In these settings, estimators exist that converge at an optimal rate with respect to the intrinsic dimension $r$, see (Bickel and Li, 2007; Kpotufe, 2011; Kpotufe and Garg, 2013; Liao et al., 2016, 2022; Liu et al., 2024a), with at least some of them with associated fast algorithms, for example the estimators in (Liao et al., 2016, 2022) can be constructed in time $\mathcal{O}(C^{r}d^{s}n\log n)$ and they achieve, adaptively, min-max optimal (up to log factors) rate on a large family of function spaces with unknown regularity, possibly varying across locations and scales.
 Functions with high degree of regularity.
 In a different direction, one may consider function classes with smaller complexity (e.g., as measured by metric entropy) than $s$-Hölder functions on $\mathbb{R}^{d}$.
-A straightforward but highly limiting assumption is that the Hölder exponent $s$ is proportional to $d$ so that the min-max rate above is independent of $d$.
-Larger function classes are obtained by imposing integrability of the Fourier transform (Barron, 1993), or strong mixed-smoothness (Strömberg, 1998), for example by using highly anisotropic Besov spaces.
-These choices however still require a degree of smoothness increasing with $d$ in order to avoid the curse of dimensionality.
-For example in these anisotropic Besov spaces, functions need to have $\mathcal{O}(1)$ axis-oriented directions in which they have “regular, $\mathcal{O}(1)$ smoothness”, and have “$\mathcal{O}(d)$ smoothness” in all other directions.
-In the context of deep neural networks, Suzuki and Nitanda (2024) demonstrate that functions in this class can be well-approximated well, with suitable architectures, in a way that avoids the curse of dimensionality if the degree of anisotropic smoothness is sufficiently large, while leaving open the aspects of learning and optimization.
+One straightforward but highly limiting assumption is that the Hölder exponent $s$ is proportional to $d$ so that the min-max rate above is independent of $d$.
+Larger function classes are obtained by either imposing strong mixed-smoothness (Strömberg, 1998) such as highly anisotropic Besov spaces (where functions need to have $\mathcal{O}(1)$ axis-oriented directions in which they have “regular, $\mathcal{O}(1)$ smoothness”, and have “$\mathcal{O}(d)$ smoothness” in all other directions), or imposing integrability of the Fourier transform (Barron, 1993).
+In the context of deep neural networks, Suzuki and Nitanda (2024) demonstrate that functions in this class can be approximated well, with suitable architectures, in a way that avoids the curse of dimensionality if the degree of anisotropic smoothness is sufficiently large (“$\mathcal{O}(d)$”), while leaving open the aspects of learning and optimization.
+The condition of belonging to a Barron’s space requires sufficient fast decaying rate for the Fourier transform of $\nabla F$, which is a condition that gets stronger with the dimension, except for very special cases, among which, notably the single- and multi-index models that we discuss momentarily, where the Fourier transform of $\nabla F$ is a singular measure.
+This requirement excludes intrinsically low-dimensional models from the Barron’s space since, as it is well-known in harmonic analysis, the Fourier transform of a measure supported on a general nonlinear curve decays slowly (Arkhipov et al., 1987; Brandolini et al., 2007).
+Similar arguments can be generalized to our main model, where the distribution is supported in a neighborhood of some nonlinear curve; our main model here is therefore typically not included in the Barron’s space, though both the Barron space and our main model include the single-index model as a special case.
 Compositional models. Yet other functional classes are obtained by imposing structural assumptions, which are often in the form of compositional models. For example assume that
@@
 Another direction is to design different types of neural networks that exploit nonlinear compositions, see (Lai and Shen, 2021; Liu et al., 2024b).
-Most of these results only address whether a function can well-approximated by a neural network, and do not cover the learning problem nor address computational aspects, in particular whether some optimization algorithm can find (efficiently?) the parameters of the desired estimator.
+Most of these results only address whether a function can be well-approximated by a neural network, and do not cover the learning problem nor address computational aspects, in particular whether some optimization algorithm can find efficiently the parameters of the desired estimator.
 Another aspect that is often left unaddressed is the dependency of constants on the ambient dimension, which is unfortunately often exponential: see Shamir (2020) for a discussion of some of these aspects, and Shen et al.
@@
 The particular case $r=1$ is called the single-index model, where the function $F$ has the structure $F(x)=f(\langle v,x\rangle)$ for some unknown index vector $v\in\mathbb{S}^{d-1}$ and unknown link function $f\in\mathcal{C}^{s}(\mathbb{R}^{1})$.
-These models have been intensively studied: Stone (1982) conjectured that the min-max rate for regression of single-index models is $n^{-\frac{s}{2s+1}}$ (resp.
+These models have been intensively studied: Stone (1982) conjectured that the min-max rate for regression for the single-index model is $n^{-\frac{s}{2s+1}}$ (resp.
 $n^{-\frac{s}{2s+r}}$ for multi-index models), coinciding with the one-dimensional min-max nonparametric rate, thereby escaping the curse of dimensionality.
@@
 (2024), and references therein, showing that deep neural networks appear to exploit low-dimensional linear subspaces for classification and prediction.
-Several methods for estimating the index vector $v$ (or the matrix $G$) and $f$ were developed over the years, with varying tradeoffs between the assumptions needed to produce optimal estimators and the computational cost for constructing them.
-One category includes semiparametric methods based on maximum likelihood estimation (Ichimura, 1993; Hardle et al., 1993; Delecroix et al., 1997; Delecroix and Hristache, 1999; Delecroix et al., 2006; R.
+Several methods for jointly estimating the index vector $v$ (or the matrix $G$) and $f$ were developed over the years, with varying tradeoffs between the assumptions needed to produce optimal or near-optimal estimators and the computational cost for constructing them, see Lanteri et al.
+(2022) for an extended discussion.
+One category includes semiparametric methods based on maximum likelihood estimation (Ichimura, 1993; Hardle et al., 1993; Delecroix et al., 2003; Delecroix and Hristache, 1999; Delecroix et al., 2006; R.
 J.
@@
 A category of techniques particularly relevant to the present work includes conditional, or inverse regression methods, which derive their estimators from statistics of the conditional distribution of the explanatory variable $X$ given observations of the (noisy) response variable $Y$.
-Prominent examples include sliced inverse regression (Duan and Li, 1991; Li, 1991), sliced average variance estimation (Cook, 2000), simple contour regression (Li et al., 2005; Coudret et al., 2013), with its analysis (in the multi-index case) in the work (Li and Wang, 2007), which yields an estimator with near-optimal rates for the multi-index model.
+Prominent examples include sliced inverse regression (Duan and Li, 1991; Li, 1991), sliced average variance estimation (Cook, 2000), simple contour regression (Li et al., 2005; Coudret et al., 2014), with its analysis (in the multi-index case) in the work (Li and Wang, 2007), which yields an estimator with near-optimal rates for the multi-index model.
 Conditional methods partition the range of the response variable $Y$ into small intervals and consider their pre-images that are, when $f$ is monotone, slices with the thinnest side along the index direction $v$.
 They are often straightforward in implementation, consisting of the computation of conditional empirical moments, or other statistics related to the level sets of $f$, and can have only one parameter to tune: the width of slices.
-Unfortunately, several of the works above, including (Hristache et al., 2001; Li and Wang, 2007), while obtaining asymptotic learning rates that avoid the curse of dimensionality and are in some cases min-max optimal, require a minimal number of samples that is exponential in the ambient dimension $d$ (either explicitly, or through constants exponential in $d$ in the bounds).
+Unfortunately, several of the works above, including (Hristache et al., 2001; Li and Wang, 2007), while obtaining asymptotic learning rates that avoid the curse of dimensionality and are in some cases min-max optimal or near-optimal, require a minimal number of samples that is exponential in the ambient dimension $d$ (either explicitly, or through constants exponential in $d$ in the bounds).
 Besides the statistical view that the single- and multi-index model do not entail a statistical cost cursed by the ambient dimension, another critical problem is to find algorithms with reasonable computational cost to implement a statistically optimal estimator for these models.
@@
 Lanteri et al.
-(2022) discuss in detail these techniques for the single-index model, and introduce of a variation, called Smallest Vector Regression, which enjoys optimal statistical guarantees up to log factors, provides a theoretically optimal choice for selecting of the slice width parameter (a crucial parameter whose choice is often not discussed (Coudret et al., 2013, p.
-75)), and is amenable to a computationally-efficient implementation with cost $\mathcal{O}(d^{2}n\log n)$, all without the curse of dimensionality in the exponents and in the constants, for the rate, the minimum sample requirements and the computational cost.
-In conclusion, while the situation is rather well-understood for compositional models $F=f\circ g$ with $g$ linear, much is open when $g$ is nonlinear, in which case it is not clear when the compositional structure allows one to circumvent the curse of dimensionality, except in cases where $g$ has high regularity (global or anisotropic), increasing with $d$, or belong to certain families of polynomials.
-We propose a simple model that, perhaps surprisingly, without such strong regularity assumptions, still is amenable to nonparametric estimators that avoid the curse of dimensionality, both statistically and computationally.
+(2022) discuss in detail these techniques for the single-index model, and introduce of a variation, called Smallest Vector Regression, which enjoys optimal (possibly up to log terms) statistical guarantees up to log factors, provides a theoretically optimal choice for selecting of the slice width parameter (a crucial parameter whose choice is often not discussed (Coudret et al., 2014, p.
+75), and is amenable to a computationally-efficient implementation with cost $\mathcal{O}(d^{2}n\log n)$, all without the curse of dimensionality in the exponents and in the constants, for the rate, the minimum sample requirements and the computational cost.
+In conclusion, while the situation is rather well-understood for compositional models $F=f\circ g$ with $g$ linear, much is open when $g$ is nonlinear, in which case it is not clear when the compositional structure allows one to circumvent the curse of dimensionality, except in cases where $g$ has high regularity (global or anisotropic), increasing with $d$, or (in particular) belongs to certain families of polynomials.
+In fact, we are not aware of models where $g$ is nonlinear, with regularity not infinite nor growing with $D$, for which $F$ may be estimated without the curse of dimensionality.
+We propose a simple model that, perhaps surprisingly, without strong regularity assumptions, still is amenable to nonparametric estimators that avoid the curse of dimensionality, both statistically and computationally.
 1.1 The Nonlinear Single-Variable Model
@@
 $\mathbb{E}[Y|X]=F(X)=f(\Pi_{\gamma}X)$
-where the underlying curve $\gamma:[0,\mathrm{len}_{\gamma}]\to\mathbb{R}^{d}$ is parametrized by arc-length and has length $\mathrm{len}_{\gamma}$ , and the link function $f\,:\,[0,\mathrm{len}_{\gamma}]\to\mathbb{R}^{1}$ depends only on one variable.
+where the underlying curve $\gamma:[0,\mathrm{len}_{\gamma}]\to\mathbb{R}^{d}$ is twice continuously differentiable, is (without loss of generality) parametrized by arc-length and has length $\mathrm{len}_{\gamma}$ , and the link function $f\,:\,[0,\mathrm{len}_{\gamma}]\to\mathbb{R}^{1}$ depends only on one variable.
 The random vector $X$ is supported in some domain $\Omega_{\gamma}\subseteq\mathbb{R}^{d}$ containing the image of $\gamma$, such that the closest-point projection $\Pi_{\gamma}:\Omega_{\gamma}\to[0,\mathrm{len}_{\gamma}]$, with
@@
 with $t=\Pi_{\gamma}X$, $t$ a random variable on $[0,\mathrm{len}_{\gamma}]$ and $Z_{d-1}$ a centered random vector in $\mathbb{R}^{d-1}$.
-For each unit vector $v\in\mathbb{S}^{d-1}$, $M_{v}\in\mathbb{O}(d)$ is an orthogonal matrix that maps the $d$-th canonical basis vector $\hat{e}_{d}$ to the vector $v$.
-Therefore, conditioned on $t=t_{0}\in[0,\mathrm{len}_{\gamma}]$, the random vector $M_{\gamma^{\prime}(t_{0})}\begin{pmatrix}Z_{d-1}\\ 0\in\end{pmatrix}\in\operatorname{span}\{\gamma^{\prime}(t_{0})\}^{\perp}$ is the displacement of $X$ away from $\gamma$.
-The regression problem for the Nonlinear Single-Variable Model: Given pairs $(X_{i},Y_{i})_{i=1}^{n}$ with $X_{i}$’s independent copies of $X$ as above, and $Y_{i}=F(X_{i})+\zeta_{i}$, with $F$ as in Definition 1 and observational noise $\zeta_{i}$, we goal is to estimate the regression function $F$ on the support of $X$.
+For each unit vector $v\in\mathbb{S}^{d-1}$, we let $M_{v}\in\mathbb{O}(d)\subseteq\mathrm{GL}(d,\mathbb{R})$ denote an orthogonal matrix that maps the $d$-th canonical basis vector $\hat{e}_{d}$ to the vector $v$.
+Therefore, conditioned on $t=t_{0}\in[0,\mathrm{len}_{\gamma}]$, the random vector $M_{\gamma^{\prime}(t_{0})}\begin{pmatrix}Z_{d-1}\\ 0\end{pmatrix}\in\operatorname{span}\{\gamma^{\prime}(t_{0})\}^{\perp}$ is the displacement of $X$ away from $\gamma$.
+The regression problem for the Nonlinear Single-Variable Model: Given pairs $(X_{i},Y_{i})_{i=1}^{n}$ with $X_{i}$’s independent copies of $X$ as above, and $Y_{i}=F(X_{i})+\zeta_{i}$, with $F$ as in Definition 1 and observational noise $\zeta_{i}$, the goal is to estimate the regression function $F$ on the support of $X$.
 The $\zeta_{i}$’s are assumed sub-Gaussian, independent among themselves and from the $X_{i}$’s.
@@
 Figure 1: One example of a Nonlinear Single-Variable Model (1): the underlying curve $\gamma$, plotted in black, is a Meyer helix in $\mathbb{R}^{36}$ (details in Appendix C) with $\sigma_{\gamma}=0.5$ and the link function $f\in\mathcal{C}^{0.7}(\mathbb{R}^{1})$ is strictly monotone, and $\zeta\equiv 0$.
-We generate $n=5000$ samples $X_{i}$ scattered the curve in a tube of radius 6, colored by $Y_{i}=F(X_{i})=f(\Pi_{\gamma}X_{i})$.
+We generate $n=5000$ samples $X_{i}$ scattered near the curve in a tube of radius 6, colored by $Y_{i}=F(X_{i})=f(\Pi_{\gamma}X_{i})$.
 Left: Random projection of the data onto $\mathbb{R}^{3}$.
@@
 The function space $\mathcal{C}^{s}(\mathbb{R}^{m})$ consists of functions $f\in L^{\infty}(\mathbb{R}^{m})$ with finite semi-norm $\left[f\right]_{\mathcal{C}^{s}}$.
-We ready for an informal version of the main Theorems 3 and 4:
+We are ready for an informal version of the main Theorems 3 and 4:
 Theorem 2 (Informal)
@@
 Each empirical slice $\widehat{S}_{l,h}$ is the empirical pre-image of the interval $R_{l,h}$ in the output variable $Y$.
+For each $l,h$, let ${S_{l,h}}$ denote the conditional distribution $X|Y\in{R_{l,h}}$.
+Then the set $\widehat{S}_{l,h}$ of sample points is the empirical version of ${S_{l,h}}$, with moments that should approximate well those of ${S_{l,h}}$ provided the number of samples is large enough.
 We perform the Principal Component Analysis (PCA) of each $\widehat{S}_{l,h}$ to obtain its mean ${\widehat{\mu}_{l,h}}$ and its “significant vector” ${\widehat{v}_{l,h}}$: ${\widehat{\mu}_{l,h}}$ will be approximately on $\gamma$, and ${\widehat{v}_{l,h}}$ will be approximately tangent to $\gamma$ at ${\widehat{\mu}_{l,h}}$, yielding a local first-order approximation to $\gamma$.
-Step 2: design a distance function ${\mathrm{dist}}(x,h)$ (and its empirical version ${\widehat{\mathrm{dist}}}(x,h)$), based on the geometric shape of ${S_{l,h}}$ (and $\widehat{S}_{l,h}$ respectively), that measure how far the point $x$ is away from the slice ${S_{l,h}}$ ($\widehat{S}_{l,h}$, respectively).
+Step 2: design a distance function ${\mathrm{dist}}(x,h)$ (and its empirical version ${\widehat{\mathrm{dist}}}(x,h)$), based on the geometric shape of ${S_{l,h}}$ (and $\widehat{S}_{l,h}$ respectively), that measures how far the point $x$ is away from the slice ${S_{l,h}}$ ($\widehat{S}_{l,h}$, respectively).
 By assigning each $x\in\Omega_{\gamma}$ to the “nearest” slice according to this distance function, we partition the domain $\Omega_{\gamma}$ into several local neighborhoods and we use the significant vector ${\widehat{v}_{l,h}}$ to project the points $X_{i}\in\widehat{S}_{l,h}$ onto a one-dimensional interval $I^{(l,h)}$ in each local neighborhood.
@@
 For $x\in\Omega_{\gamma}$ we define the signed projected distance from $x$ to $\gamma$ to be $\tilde{d}(x,\gamma):=\langle x-\gamma(\Pi_{\gamma}x),\hat{n}(\Pi_{\gamma}x)\rangle_{\mathbb{R}^{d}}$, which is zero if $x$ is on the curve $\gamma$, positive if $x$ is inside the circle of curvature, and negative otherwise.
-A direct calculation shows that the compositional structure $F=f\circ\Pi_{\gamma}$ in Definition 1 implies that for $x\in\Omega_{\gamma}$
+A direct calculation shows that the compositional structure $F=f\circ\Pi_{\gamma}$ in Definition 1 implies that, for curve $\gamma$ parameterized by arc-length and for $x\in\Omega_{\gamma}$
 ${\nabla F}(x)=\frac{f^{\prime}(\Pi_{\gamma}x)}{1-\left\|\gamma^{\prime\prime}(\Pi_{\gamma}x)\right\|\tilde{d}(x,\gamma)}{\gamma^{\prime}}(\Pi_{\gamma}x)\,.$ (2)
-Therefore, the gradient vectors of points in each level set $\Pi_{\gamma}^{-1}(t_{0})=\{x:\Pi_{\gamma}x=t_{0}\}$ are all parallel to ${\gamma^{\prime}}(t_{0})$, albeit with magnitude depending on the relative position of $x$ and $\gamma$.
+Therefore, the gradient vectors of points in each level set $\Pi_{\gamma}^{-1}(t_{0})=\{x:\Pi_{\gamma}x=t_{0}\}$ are all parallel to the unit-vector ${\gamma^{\prime}}(t_{0})$, albeit with magnitude depending on the relative position of $x$ and $\gamma$.
 As a consequence, each $\Pi_{\gamma}^{-1}(t_{0})$ is contained in a hyperplane.
@@
 Instead, we can only rely on local information to estimate the nearest-point projection onto $\gamma$, which is now a nonlinear function.
-Each $\widehat{S}_{l,h}$ still approximates some level set $\Pi_{\gamma}^{-1}(t_{0})$ under the Assumption (LCV) below: roughly speaking, when ${R_{l,h}}$ has a sufficiently small diameter, the $\widehat{S}_{l,h}$ is “thin” along the tangential direction $\gamma^{\prime}(t_{0})$ and “wide” along directions orthogonal to $\gamma^{\prime}(t_{0})$.
+Each $\widehat{S}_{l,h}$ still approximates some level set $\Pi_{\gamma}^{-1}(t_{0})$ under the Assumption (LCV) in Section 3 below: roughly speaking, when ${R_{l,h}}$ has a sufficiently small diameter, the $\widehat{S}_{l,h}$ is “thin” along the tangential direction $\gamma^{\prime}(t_{0})$ and “wide” along directions orthogonal to $\gamma^{\prime}(t_{0})$.
 Consequently, PCA on each empirical slice $\widehat{S}_{l,h}$ locally will yield:
@@
 We use the sample values $\{Y_{i}:X_{i}\in\bigcup_{|h-{\widehat{h}_{x}}|\leqslant 1}\widehat{S}_{l,h}\}$ to construct a local polynomial ${\widehat{f}_{j,k|{\widehat{v}_{l,\widehat{h}_{x}}}}}$ on each interval $I^{(l,{\widehat{h}_{x}})}_{j,k}$ of the partition by solving a least squares fitting problem, obtaining a piecewise polynomial estimator ${\widehat{f}_{j|{\widehat{v}_{l,\widehat{h}_{x}}}}}$ of the link function.
-The degree of the local polynomials needed to obtain optimal estimation rates depends on the regularity of the function.
+The degree of the local polynomials needed to obtain optimal (up to log factors) estimation rates depends on the regularity of the function.
 A proper partition (or scale) is then chosen to minimize the expected mean squared error (MSE) using classical bias-variance trade-off arguments.
@@
     Compute interval $R$ of range of samples $\{Y_{i}\}_{i=1}^{n}$.
-Construct $\{{R_{l,h}}\}_{h=1}^{l}$, the uniform partition of $R$ into $l$ intervals whose preimages are $\{\widehat{S}_{l,h}\}_{h=1}^{l}$ where $\widehat{S}_{l,h}=\{X_{i}:Y_{i}\in{R_{l,h}}\}$;
+Construct $\{{R_{l,h}}\}_{h=1}^{l}$, the uniform partition of $R$ into $l$ intervals whose preimages are $\{\widehat{S}_{l,h}\}_{h=1}^{l}$ where $\widehat{S}_{l,h}=\{X_{i}:Y_{i}\in{R_{l,h}}\}$.
   2. 1.b
     Denote ${n_{l,h}}=\#\{Y_{i}\in{R_{l,h}}\}$. For each $h\in\mathcal{H}_{l}=\{h:{n_{l,h}}\geq n/l\}$, compute:
-${\widehat{\mu}_{l,h}}=\frac{1}{{n_{l,h}}}\sum_{i}X_{i}\mathbbm{1}\{X_{i}\in\widehat{S}_{l,h}\}$, the empirical mean
-${\widehat{\Sigma}_{l,h}}=\frac{1}{{n_{l,h}}}\sum_{i}(X_{i}-{\widehat{\mu}_{l,h}})(X_{i}-{\widehat{\mu}_{l,h}})^{\intercal}\mathbbm{1}\{X_{i}\in\widehat{S}_{l,h}\}$, the empirical covariance matrix
-${\widehat{\lambda}_{l,h,m}}$, the $m$-th eigenvalue of ${\widehat{\Sigma}_{l,h}}$, in descending order
-${\widehat{v}}_{l,h,m}$, the $m$-th eigenvector of ${\widehat{\Sigma}_{l,h}}$ corresponding to eigenvalue ${\widehat{\lambda}_{l,h,m}}$
-$H_{l,h}=\log\left(\frac{({\widehat{\lambda}_{l,h,2}}\times\cdots\times{\widehat{\lambda}_{l,h,d-1}})^{{2}/{(d-2)}}}{{\widehat{\lambda}_{l,h,1}}{\widehat{\lambda}_{l,h,d}}}\right)$
-the empirical significant vector ${\widehat{v}_{l,h}}$, equals ${\widehat{v}}_{l,h,d}$ if $H_{l,h}>0$ and to ${\widehat{v}}_{l,h,1}$ otherwise
+${\widehat{\mu}_{l,h}}=\frac{1}{{n_{l,h}}}\sum_{i}X_{i}\mathbbm{1}\{X_{i}\in\widehat{S}_{l,h}\}$, the empirical mean;
+${\widehat{\Sigma}_{l,h}}=\frac{1}{{n_{l,h}}}\sum_{i}(X_{i}-{\widehat{\mu}_{l,h}})(X_{i}-{\widehat{\mu}_{l,h}})^{\intercal}\mathbbm{1}\{X_{i}\in\widehat{S}_{l,h}\}$, the empirical covariance matrix;
+${\widehat{\lambda}_{l,h,m}}$, the $m$-th eigenvalue of ${\widehat{\Sigma}_{l,h}}$, in descending order;
+${\widehat{v}}_{l,h,m}$, the $m$-th eigenvector of ${\widehat{\Sigma}_{l,h}}$ corresponding to eigenvalue ${\widehat{\lambda}_{l,h,m}}$;
+$\widehat{H}_{l,h}=\log\left(\frac{({\widehat{\lambda}_{l,h,2}}\times\cdots\times{\widehat{\lambda}_{l,h,d-1}})^{{2}/{(d-2)}}}{{\widehat{\lambda}_{l,h,1}}{\widehat{\lambda}_{l,h,d}}}\right)$  .
+Let the empirical significant vector ${\widehat{v}_{l,h}}$ equal ${\widehat{v}}_{l,h,d}$ if $\widehat{H}_{l,h}>0$ and equal ${\widehat{v}}_{l,h,1}$ otherwise.
   • 2.a
-    Given $x\in\mathbb{R}^{d}$, for each $h\in\mathcal{H}_{l}$ compute the estimated distance between $x$ and $\widehat{S}_{l,h}$
-    ${\widehat{\mathrm{dist}}}(x,h)=\begin{cases}|\langle x-{\widehat{\mu}_{l,h}},{\widehat{v}_{l,h}}\rangle|^{2}+\frac{{\lambda_{d}(\widehat{\Sigma}_{l,h})}}{{\lambda_{1}(\widehat{\Sigma}_{l,h})}}\left\|x-{\widehat{\mu}_{l,h}}\right\|^{2}\ \text{ if }\widehat{H}_{l,h}>0\\ \left\|x-{\widehat{\mu}_{l,h}}\right\|^{2}+\frac{{\lambda_{d}(\widehat{\Sigma}_{l,h})}}{{\lambda_{1}(\widehat{\Sigma}_{l,h})}}|\langle x-{\widehat{\mu}_{l,h}},{\widehat{v}_{l,h}}\rangle|^{2}\ \ \text{ if }\widehat{H}_{l,h}<0\end{cases}$
+    Given $x\in\mathbb{R}^{d}$, for each $h\in\mathcal{H}_{l}$, compute the estimated distance between $x$ and $\widehat{S}_{l,h}$:
+    ${\widehat{\mathrm{dist}}}(x,h)=\begin{cases}|\langle x-{\widehat{\mu}_{l,h}},{\widehat{v}_{l,h}}\rangle|^{2}+\frac{{\lambda_{d}(\widehat{\Sigma}_{l,h})}}{{\lambda_{1}(\widehat{\Sigma}_{l,h})}}\left\|x-{\widehat{\mu}_{l,h}}\right\|^{2}\ \text{ if }\widehat{H}_{l,h}>0\\ \left\|x-{\widehat{\mu}_{l,h}}\right\|^{2}+\frac{{\lambda_{d}(\widehat{\Sigma}_{l,h})}}{{\lambda_{1}(\widehat{\Sigma}_{l,h})}}|\langle x-{\widehat{\mu}_{l,h}},{\widehat{v}_{l,h}}\rangle|^{2}\ \ \text{ if }\widehat{H}_{l,h}<0\end{cases}\,.$
   • 2.b
-    compute the estimated nearest slice index ${\widehat{h}_{x}}=\operatorname*{{arg\,min}}_{h\in\mathcal{H}_{l}}{\widehat{\mathrm{dist}}}(x,h)$
+    Compute the estimated nearest slice index ${\widehat{h}_{x}}=\operatorname*{{arg\,min}}_{h\in\mathcal{H}_{l}}{\widehat{\mathrm{dist}}}(x,h)$.
   • 3.a
@@
     For each $h\in\mathcal{H}_{l}$ and each $k\in\mathcal{K}_{j}=\{k:n^{(l,h)}_{j,k}\geqslant{n_{l,h}}/j\}$ compute
-    ${\widehat{f}_{j,k|{\widehat{v}_{l,h}}}}=\operatorname*{{arg\,min}}_{\deg(p)\leqslant m}\sum_{X_{i}\in\bigcup_{|h-{\widehat{h}_{x}}|\leqslant 1}\widehat{S}_{l,h}}\left|Y_{i}-p(\langle{\widehat{v}_{l,h}},X_{i}\rangle)\right|^{2}\mathbbm{1}_{{I^{(l,h)}_{j,k}}}(\langle{\widehat{v}_{l,h}},X_{i}\rangle)$
+    ${\widehat{f}_{j,k|{\widehat{v}_{l,h}}}}=\operatorname*{{arg\,min}}_{\deg(p)\leqslant m}\sum_{X_{i}\in\bigcup_{|h-{\widehat{h}_{x}}|\leqslant 1}\widehat{S}_{l,h}}\left|Y_{i}-p(\langle{\widehat{v}_{l,h}},X_{i}\rangle)\right|^{2}\mathbbm{1}_{{I^{(l,h)}_{j,k}}}(\langle{\widehat{v}_{l,h}},X_{i}\rangle)\ \,.$
   • 3.c
     For each $h\in\mathcal{H}_{l}$, compute ${\widehat{f}_{j|{\widehat{v}_{l,h}}}}(r)=\sum_{k\in\mathcal{K}_{j}}{\widehat{f}_{j,k|{\widehat{v}_{l,h}}}}(r)\mathbbm{1}_{{I^{(l,h)}_{j,k}}}(r)$ and return the estimator
-    ${\widehat{F}}(x)={\widehat{f}_{j|{\widehat{v}_{l,\widehat{h}_{x}}}}}(\langle{\widehat{v}_{l,\widehat{h}_{x}}},x\rangle)$
+    ${\widehat{F}}(x)={\widehat{f}_{j|{\widehat{v}_{l,\widehat{h}_{x}}}}}(\langle{\widehat{v}_{l,\widehat{h}_{x}}},x\rangle)\ \,.$
   • Algorithm 1 Significant Vector Regression
@@
     Algorithm 1 summarizes the construction of the proposed estimator of the regression function $F$.
-We prove that the mean squared error of our estimator is the sum of the estimation error and the curve approximation error.
-The estimation error decays the one-dimensional min-max optimal nonparametric learning rate, up to a saturation level dependent on quantities expected to be small, or even zero, as we discuss in detail below.
+We prove, under assumptions that are detailed in Section 3, that the mean squared error of our estimator is the sum of the estimation error and the curve approximation error.
+The estimation error decays at the one-dimensional min-max optimal nonparametric learning rate, up to log factors and, more importantly, up to a saturation level dependent on quantities expected to be small, or even zero, as we discuss in detail below.
     Theorem 3 (MSE of the Estimator constructed by Algorithm 1)
-    Assume that (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), ($\bm{\omega_{f}}$), and (P) hold true, and that $f\in\mathcal{C}^{s}(\mathbb{R}^{1})$ with $s\in\left[\frac{1}{2},2\right]$.
-Let $C_{\gamma,f},M^{*},l_{\max}$ be the constants specified in (4) below, and $C$ a universal constant.
-Then, for $n$ such that $\frac{n}{\log^{{3/2}}n}\gtrsim C_{\gamma,f}\frac{C_{f}\mathrm{len}_{\gamma}}{C_{f}^{\prime}\sigma_{\gamma}}$, if we choose $l^{*},j^{*}$ as
+[…]
     $(l^{*},j^{*}):=\begin{cases}\left(C_{\gamma,f}^{-1}n\log^{-2}n,\,C\right)&\text{ if }\frac{n^{\frac{2s}{2s+1}}}{\log^{2}n}\lesssim C_{\gamma,f}M^{*}\\ \left(l_{\max},\,C\frac{M^{*}}{l_{\max}}n^{\frac{1}{2s+1}}\right)&\text{ if }n^{\frac{1}{2s+1}}\gtrsim\frac{CC_{Y}R_{0}}{M^{*}\max(\sigma_{\zeta},\omega_{f})}\\ \left(n^{\frac{1}{2s+1}}M^{*},\,C\right)&\text{ otherwise}\end{cases}\ \,,$
     the estimator constructed by Algorithm 1, in running time $\mathcal{O}(d^{2}n\log n)$, satisfies
-    $\mathbb{E}\left[\left|{\widehat{F}}(X)-F(X)\right|^{2}\right]\lesssim C_{1}(f,\gamma,\rho_{X},\sigma_{\zeta},d)n^{-\frac{2s}{2s+1}}{\log n}+C_{2}(f,\gamma,\rho_{X},\sigma_{\zeta},d)$
+    $\mathbb{E}\left[\left|{\widehat{F}}(X)-F(X)\right|^{2}\right]\lesssim C_{1}(f,\gamma,\rho_{X},\sigma_{\zeta},d)n^{-\frac{2s}{2s+1}}{\log n}+C_{2}(f,\gamma,\rho_{X},\sigma_{\zeta},d)\ \,.$
     We specify the constants in the Theorem above:
@@
     $M^{*}$ $:=\left(\sigma_{\zeta}^{-1}(C_{f}C_{Y}R_{0})^{s}(\left[f\right]_{\mathcal{C}^{s}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s}})\right)^{\frac{2}{2s+1}}\quad,\quad l_{\max}:=\frac{CC_{Y}R_{0}}{\max(\sigma_{\zeta},\omega_{f})}\,.$
-    The above theorem states that our estimator achieves the min-max optimal rate for the one-dimensional nonparametric regression, with an additional approximation error of magnitude $\mathcal{O}(\sigma_{\zeta}^{2})$ for functions that are both Lipschitz and monotone, therefore defeating the curse of dimensionality by exploiting the compositional structure of $F$, even if the inner function is nonlinear (unlike the single- and multi-index model) and not particularly smooth (and its regularity does not scale with the ambient dimension).
+    The above theorem states that our estimator achieves the min-max optimal rate, (up to log factors), for the one-dimensional nonparametric regression, with an additional approximation error of magnitude $\mathcal{O}(\sigma_{\zeta}^{2})$ for functions that are both Lipschitz and monotone, therefore defeating the curse of dimensionality by exploiting the compositional structure of $F$, even if the inner function is nonlinear (unlike the single- and multi-index model) and not particularly smooth (and its regularity does not scale with the ambient dimension).
 It is worth mentioning that the result of Theorem 3 is scaling invariant in $X$ and $Y$.
@@
       the minimal number of samples $n$ required is only a low-order polynomial in $d$, $\mathrm{len}_{\gamma}$, and $\left\|\gamma^{\prime\prime}\right\|$.
-This is not exponential in ambient dimension $d$ (unless in the extreme case where $\gamma$ has length or curvature growing exponentially in $d$, e.g.
-a space-filling curve).
+This is not exponential in ambient dimension $d$ (unless in the extreme case where $\gamma$ has length or curvature growing exponentially in $d$, e.g., a space-filling curve);
     4. (iv)
+      the regularity assumptions on both $f$ and $\gamma$ are independent of the dimension $d$;
+    5. (v)
       Algorithm 1 in fact also estimates $\gamma$, the nonlinear projection $\Pi_{\gamma}$, and $f$.
@@
       the assumption that $f$ is coarsely monotone.
-While similar conditions have appeared in other works using conditional regression (see Lanteri et al., 2022, for a discussion), and even in recent approaches based on gradient descent for the single- and multi-index models (see e.g., Damian et al., 2024; Arous et al., 2021; Bietti et al., 2023, and references therein) the “amount of oscillation” of $f$ appears prominently and imposes additional sampling requirements (e.g., scaling as $d^{\mathcal{O}(L)}$ where $L$ is the number of vanishing moments in the Hermite polynomial basis, albeit it is not clear to us that such assumptions and sampling requirements are sharp);
+While similar conditions have appeared in other works using conditional regression (see Lanteri et al., 2022, for a discussion), and even in recent approaches based on gradient descent for the single- and multi-index models (see e.g., Damian et al., 2024; Arous et al., 2021; Bietti et al., 2023, and references therein) the “amount of oscillation” of $f$ appears prominently and imposes additional sampling requirements (e.g., scaling as $d^{\mathcal{O}(L)}$ where $L$ is the number of vanishing moments in the Hermite polynomial basis, albeit it is not clear to us that such assumptions and sampling requirements are sharp).
+At this point we do not know if this limitation can be overcome by estimators similar to ours with a cost still polynomial in $d$.
+In similar and related models, albeit not applicable to the nonlinear variable model introduced here, the estimators in Bach (2017) and Chen and Meka (2020) do not require restrictive assumptions, so it may indeed the case that this restrictions are due to the specific family of estimators we introduce here.
     2. (ii)
@@
 In the limiting case where the underlying curve $\gamma$ is a straight line segment, we have no curve approximation error, recovering the results of (Lanteri et al., 2022).
-Otherwise, the nonzero approximation error exists because we are using first-order approximations in the estimation of the local directions of the curve, and with noisy observations our technique for constructing the estimator does not allow us to consider local pieces below the scale of the noise.
-While it seems possible to use approximations of higher order, they would require significant additional computational complexity and more refined statistical analysis, possibly with minimal impact in practical applications, and are left to future work.
+Otherwise, the nonzero approximation error appears here because we are using first-order approximations in the estimation of the local directions of the curve, and with noisy observations our technique for constructing the estimator does not allow us to consider local pieces below the scale of the noise, no matter what the sample size.
+While it seems possible to use approximations of higher order, increasing with $n$, would allow us to obtain a term vanishing in $n$, they would require significant additional computational complexity and more refined statistical analysis, possibly with minimal impact in practical applications, and are left to future work.
+At this point we do not know if this is an artifact of the use of inverse-regression techniques, or if any polynomial (in $n$ and $D$) time algorithm would necessarily require similar hypotheses.
     In the case of a strictly monotone Lipschitz function with zero external noise, we obtain the rate $\mathcal{O}(n^{-2})$.
@@
     Theorem 4 (MSE of Algorithm 1 in the noiseless case)
-    Assume that (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), ($\bm{\omega_{f}}$), and (P) hold true.
-Assume that there is no observational noise, i.e., $\zeta\equiv 0$ almost surely, that the link function $f$ is perfectly monotone, i.e., $\omega_{f}=0$, and that $f\in\mathcal{C}^{s}$ with $s\in\left[\frac{1}{2},2\right]$.
-With $C_{\gamma,f}$ in (4), when $\frac{n}{\log^{3/2}n}\gtrsim C_{\gamma,f}\frac{C_{f}\mathrm{len}_{\gamma}}{C_{f}^{\prime}\sigma_{\gamma}}$, if we choose $l^{*}=\frac{1}{C_{\gamma,f}}\frac{n}{\log^{3/2}n}$ and $j^{*}=C$, then the estimation error of Algorithm 1 satisfies
+[…]
     $\mathbb{E}\left[\left|{\widehat{F}}(X)-F(X)\right|^{2}\right]\lesssim(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}\left(C_{\gamma,f}(C_{f}C_{Y}R_{0})^{2}\frac{\log^{3}n}{n^{2}}\right)^{s\wedge 1}\,.$
@@
 (2017) proves that, in the case of the $L^{\infty}$-norm, the min-max rate of nonparametric regression for functions in $\mathcal{C}^{s}(\mathbb{R}^{d})$ with noiseless observations is $(\log n/n)^{s/d}$.
-The rate on the mean squared error in Theorem 4 agrees with this when $s\leqslant 1$ and $d=1$, as if $\gamma$ (and therefore $\Pi_{\gamma}$) were known.
+The rate on the mean squared error in Theorem 4 is consistent with this rate in $L^{\infty}$ when $s\leqslant 1$ and $d=1$, as if $\gamma$ (and therefore $\Pi_{\gamma}$) were known.
 Our rate is suboptimal for $s>1$, as we only perform a first-order approximation of the underlying curve and do not estimate higher-order parameters such as the curvature.
@@
     Theorem 5 (NVM without (LCV))
-    Assume that (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), ($\bm{\omega_{f}}$), and (SC) hold true.
-When $\frac{n}{\log^{1/2}n}\gtrsim\frac{C_{Y}R_{0}^{7}d^{3}}{C_{f}^{6}\max(\sigma_{\zeta},\omega_{f})^{7}}$, if we choose $l^{*}=\frac{C_{Y}R_{0}}{\max(\sigma_{\zeta},\omega_{f})}$ and $j^{*}=C$, the estimation error of Algorithm 1 satisfies
+[…]
     $\mathbb{E}\left[\left|{\widehat{F}}(X)-F(X)\right|^{2}\right]\lesssim\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\left(\frac{\sigma_{\gamma}C_{f}}{\mathrm{reach}_{\gamma}}\max(\sigma_{\zeta},\omega_{f})\right)^{2(s\wedge 1)}\,.$
@@
 We start by collecting a few conditions on the distribution of $X,Y$, and $\zeta$ that are fairly standard:
-    1. (X subG)
-      $X$ has sub-Gaussian distribution with variance proxy $R_{0}^{2}$, and has a density function $\rho_{X}$ with $\left[\rho_{X}\right]_{\mathcal{C}^{2}}<\infty$.
-    1. (Y subG)
+[…]
+      $X$ has sub-Gaussian distribution with variance proxy $R_{0}^{2}$, and has a density function $\rho_{X}$ which is $\mathcal{C}^{2}$ with $\left[\rho_{X}\right]_{\mathcal{C}^{2}}<\infty$ and has compact support.
+[…]
       $Y$ has sub-Gaussian distribution with variance proxy $C_{Y}^{2}R_{0}^{2}$.
-    1. ($\bm{\zeta}$ subG)
+    1. ($\bm{\zeta}\in\psi_{2}$)
       $\zeta$ is sub-Gaussian with variance proxy $\sigma_{\zeta}^{2}$.
@@
 For each $t_{0}\in[0,\mathrm{len}_{\gamma}]$, the conditional random vector $Z_{d-1}|t=t_{0}$ is mean zero, isotropic with variance $\sigma_{\gamma}(t_{0})^{2}\mathbb{I}_{d-1}$, and supported in an Euclidean ball $B(\gamma(t_{0}),c\ \mathrm{reach}_{\gamma}(t_{0}))\subseteq\mathbb{R}^{d-1}$ for some $c<1$.
+    This assumption is satisfied, for example, by a natural generative model where a point $\gamma(t)$ is sampled on the curve $\gamma$ and, conditional to that, a point $X$ is produced according to (1) with $Z_{d-1}$ a sub-Gaussian distribution as in ($\bm{\gamma_{1}}$).
+Note that this does not imply that, overall, the points in the normal directions to the curve are uniformly or isotropically distributed.
 Assumption ($\bm{\gamma_{1}}$) implies that for any $t_{0}\in[0,\mathrm{len}_{\gamma}]$, the conditional mean is on the curve
@@
     $\Sigma_{T}$ $:=\mathbb{E}[(X-\mu_{T})(X-\mu_{T})^{\intercal}\mid t\in T]$
-    $=\mathbb{E}\left[(\gamma(t)-\mu)(\gamma(t)-\mu)^{\intercal}\mid t\in T\right]+\mathbb{E}[\sigma_{\gamma}(t)^{2}\mid t\in T]\mathbb{I}_{d-1}-\mathbb{E}[\sigma_{\gamma}(t)^{2}\gamma^{\prime}(t)\gamma^{\prime}(t)^{\intercal}\mid t\in T]$
+    $=\mathbb{E}\left[(\gamma(t)-\mu)(\gamma(t)-\mu)^{\intercal}\mid t\in T\right]+\mathbb{E}[\sigma_{\gamma}(t)^{2}\mid t\in T]\mathbb{I}_{d}-\mathbb{E}[\sigma_{\gamma}(t)^{2}\gamma^{\prime}(t)\gamma^{\prime}(t)^{\intercal}\mid t\in T]\ \,.$
     The above identity illustrates that when slices are thin enough, Significant Vector Regression is approximately estimating the tangential vector $\gamma^{\prime}$: if $T$ is small enough such that the first term has a negligible spectral norm, compared with the second and the third term, because the second term is a multiple of identity matrix and the third term has negative sign, the smallest principal component of $\Sigma_{T}$ is roughly the largest principal component of $\mathbb{E}[\sigma_{\gamma}(t)^{2}\gamma^{\prime}(t)\gamma^{\prime}(t)^{\intercal}\mid t\in T]$, yielding an estimate of the direction of $\mathbb{E}[\sigma_{\gamma}(t)\gamma^{\prime}(t)\mid t\in T]$.
@@
     Finally, we consider the following property, which is satisfied by various regular densities on compact normal domains with no geometrical symmetry required:
-    1. (P)
-      The density $\rho_{X}$ has compact support.
-Let $\rho_{v}$ denote the push-forward measure of the distribution of $X$ along the map $x\mapsto\langle x,v\rangle$.
-Then there exists constant $c>0$ such that, for every $v\in\mathbb{S}^{d-1}$ and for every interval $I\subseteq\operatorname{\mathrm{supp}}\rho_{v}$, we have $\rho_{v}(I)\geqslant c|I|$ and $\mathrm{Var}[\langle v,X\rangle|\langle v,X\rangle\in I]\geqslant c|I|^{2}$.
-    Assumption ($\bm{\gamma_{1}}$) implies that for $t_{0}\in[0,\mathrm{len}_{\gamma}]$, $\sigma_{\gamma}(t_{0})<\min(R_{0},\mathrm{reach}_{\gamma}/\sqrt{d})$; (LCV) implies that $\sigma_{\gamma}\geqslant 2C_{f}\sigma_{\zeta}$; ($\bm{\omega_{f}}$) implies that ${C_{f}^{\prime}}C_{Y}R_{0}\leqslant\mathrm{len}_{\gamma}\leqslant C_{f}C_{Y}R_{0}$.
+    1. ($\bm{\text{P}_{\gamma}}$)
+      Let $\rho_{t}$ denote the push-forward measure of the distribution of $X$ along the map $x\mapsto\Pi_{\gamma}(x)\in[0,\mathrm{len}_{\gamma}]$.
+Then there exists constant $c>0$ such that for every interval $I\subseteq[0,\mathrm{len}_{\gamma}]$ with $|I|\leqslant\sigma_{\gamma}$, we have $\rho_{t}(I)\geqslant c|I|$ and $\mathrm{Var}[\Pi_{\gamma}(X)|\Pi_{\gamma}(X)\in I]\geqslant c|I|^{2}$.
+    Assumption ($\bm{\text{P}_{\gamma}}$) is rather weak: it states that if we project the random variable $X$ onto the underlying curve $\gamma$ then the corresponding distribution on the curve has a nontrivial lower-bound on its density.
+    It is worth noticing that the above assumptions put some restriction on the parameters of the curve $\gamma$: Assumption ($\bm{\gamma_{1}}$) implies that for $t_{0}\in[0,\mathrm{len}_{\gamma}]$, $\sigma_{\gamma}(t_{0})<\min(R_{0},\mathrm{reach}_{\gamma}/\sqrt{d})$; (LCV) implies that $\sigma_{\gamma}\geqslant 2C_{f}\sigma_{\zeta}$; ($\bm{\omega_{f}}$) implies that ${C_{f}^{\prime}}C_{Y}R_{0}\leqslant\mathrm{len}_{\gamma}\leqslant C_{f}C_{Y}R_{0}$.
     3.1 Estimation of slice parameters with Assumption (LCV)
@@
     for some $C_{X},C_{Y}\geqslant 1$ fixed constant from now on. We define the following bounded version of ${\mu_{l,h}}$ and $\Sigma_{l,h}$:
-    ${\mu_{l,h}^{b}}:=\mathbb{E}\left[X|Y\in{R_{l,h}},\mathcal{B}\right],\ \ \ {\Sigma_{l,h}^{b}}:=\operatorname{{Cov}}\left[X|Y\in{R_{l,h}},\mathcal{B}\right]$
+    ${\mu_{l,h}^{b}}:=\mathbb{E}\left[X|Y\in{R_{l,h}},\mathcal{B}\right],\ \ \ {\Sigma_{l,h}^{b}}:=\operatorname{{Cov}}\left[X|Y\in{R_{l,h}},\mathcal{B}\right]\ \,.$
     Given the event
-    $\mathcal{B}_{i}:=\{\left\|X_{i}\right\|\leqslant C_{X}\sqrt{d}R_{0},|Y_{i}|\leqslant C_{Y}R_{0}\}$
+    $\mathcal{B}_{i}:=\{\left\|X_{i}\right\|\leqslant C_{X}\sqrt{d}R_{0},|Y_{i}|\leqslant C_{Y}R_{0}\}\ \,,$
     we define
     $n^{b}:=\sum_{i}\mathbbm{1}_{\mathcal{B}_{i}}(X_{i}),\ \ \ {n_{l,h}^{b}}:=\sum_{i}\mathbbm{1}(\left\{Y_{i}\in{R_{l,h}}\right\}\cap\mathcal{B}_{i})\ \,.$
-    The random variable $n^{b}$, assuming (X subG) and (Y subG), is larger than a constant fraction of $n$ with high probability (see Lemma 21).
-The sample counterparts of ${\mu_{l,h}^{b}}$ and ${\Sigma_{l,h}^{b}}$ are
+[…]
     ${\widehat{\mu}_{l,h}^{b}}:=\frac{1}{{n_{l,h}^{b}}}\sum_{i:\left\{Y_{i}\in{R_{l,h}}\right\}\cap\mathcal{B}_{i}}X_{i}\,,\,{\widehat{\Sigma}_{l,h}^{b}}:=\frac{1}{{n_{l,h}^{b}}}\sum_{i:\left\{Y_{i}\in{R_{l,h}}\right\}\cap\mathcal{B}_{i}}(X_{i}-{\widehat{\mu}_{l,h}^{b}})(X_{i}-{\widehat{\mu}_{l,h}^{b}})^{\intercal}\,.$
@@
 Given $\mathcal{B}$, it is natural to pick the interval $R$ in Significant Vector Regression as
-    $R:=[-C_{Y}R_{0},C_{Y}R_{0}]$
+    $R:=[-C_{Y}R_{0},C_{Y}R_{0}]\ \,.$
     Taking slices $Y\in{R_{l,h}}$ is equivalent to conditioning on $Y\in{R_{l,h}}$.
@@
     Proposition 6
-    Suppose ($\bm{\zeta}$ subG) and ($\bm{\omega_{f}}$) hold true. Let $T\subseteq R$ be a bounded interval with $|T|\geqslant\max(\sigma_{\zeta},\omega_{f})$. Then:
+    Suppose ($\bm{\zeta}\in\psi_{2}$) and ($\bm{\omega_{f}}$) hold true. Let $T\subseteq R$ be a bounded interval with $|T|\geqslant\max(\sigma_{\zeta},\omega_{f})$. Then:
     1. (a)
@@
     Proposition 7
-    Suppose ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$) and ($\bm{\omega_{f}}$) hold true. Let $T\subseteq R$ be a bounded interval with $|T|\geqslant\max(\sigma_{\zeta},\omega_{f})$. Then:
+    Suppose ($\bm{\zeta}\in\psi_{2}$), ($\bm{\gamma_{1}}$) and ($\bm{\omega_{f}}$) hold true. Let $T\subseteq R$ be a bounded interval with $|T|\geqslant\max(\sigma_{\zeta},\omega_{f})$. Then:
     1. (a)
@@
     Corollary 8
-    Suppose (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), and ($\bm{\omega_{f}}$) hold true.
-Then, for every $l$ such that $l\gtrsim C_{f}C_{Y}R_{0}/\sigma_{\gamma}$, and $|{R_{l,h}^{b}}|\geqslant\max(\sigma_{\zeta},\omega_{f}),h\in\mathcal{H}_{l}^{b}$, we have
+[…]
     $\lambda_{d-1}({\Sigma_{l,h}^{b}})-\lambda_{d}({\Sigma_{l,h}^{b}})\gtrsim\sigma_{\gamma}^{2}\,.$
@@
     Corollary 9
-    Suppose (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), and ($\bm{\omega_{f}}$) hold true.
-Then, for every $l$ such that $l\gtrsim C_{f}C_{Y}R_{0}/\sigma_{\gamma}$, and $|{R_{l,h}^{b}}|\geqslant\max(\sigma_{\zeta},\omega_{f})$ for all $h\in\mathcal{H}_{l}^{b}$, for every $\epsilon>0$ and $\tau\geqslant 1$, if $n$ is sufficiently large so that $\frac{n}{\sqrt{\tau\log n}}\gtrsim\left(\frac{C_{f}C_{Y}R_{0}}{\sigma_{\gamma}}\right)^{2}dl(t+\log d+\log l)$, we have
-    $\mathbb{P}\left\{\exists h\in{\mathcal{H}_{l}^{b}}:\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|\gtrsim C_{Y}C_{f}R_{0}\sqrt{t+\log l+\log d}\sqrt{\frac{\sqrt{\tau\log n}}{nl}}\ \bigg{|}\ {n_{l,h}^{b}}\right\}\lesssim e^{-t}+n^{-\tau}$
-    $\mathbb{P}\left\{\exists h\in{\mathcal{H}_{l}^{b}}:\left\|{{\widehat{v}_{l,h}^{b}}}-{{v_{l,h}^{b}}}\right\|\gtrsim C_{Y}C_{f}R_{0}^{2}\sigma_{\gamma}^{-2}\sqrt{t+\log l+\log d}\sqrt{\frac{d\sqrt{\tau\log n}}{nl}}\ \bigg{|}\ {n_{l,h}^{b}}\right\}\lesssim e^{-t}+n^{-\tau}$
+[…]
+    $\mathbb{P}\left\{\exists h\in{\mathcal{H}_{l}^{b}}:\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|\gtrsim C_{Y}C_{f}R_{0}\sqrt{t+\log l+\log d}\sqrt{\frac{\sqrt{\tau\log n}}{nl}}\ \bigg{|}\ {n_{l,h}^{b}}\right\}\lesssim e^{-t}+n^{-\tau}\ \,;$
+    $\mathbb{P}\left\{\exists h\in{\mathcal{H}_{l}^{b}}:\left\|{{\widehat{v}_{l,h}^{b}}}-{{v_{l,h}^{b}}}\right\|\gtrsim C_{Y}C_{f}R_{0}^{2}\sigma_{\gamma}^{-2}\sqrt{t+\log l+\log d}\sqrt{\frac{d\sqrt{\tau\log n}}{nl}}\ \bigg{|}\ {n_{l,h}^{b}}\right\}\lesssim e^{-t}+n^{-\tau}\ \,.$
     Moreover, if $\frac{n}{\log^{3/2}n}\gtrsim C_{Y}^{2}C_{f}^{2}R_{0}^{4}\sigma_{\gamma}^{-4}dl(\log d+\log l)$, then for any $h\in{\mathcal{H}_{l}^{b}}$ and $p\geq 1$,
-    $\mathbb{E}\left[\left\|{\widehat{v}_{l,h}^{b}}-{v_{l,h}^{b}}\right\|^{2p}\right]\lesssim C(p)(C_{Y}C_{f}R_{0}^{2}\sigma_{\gamma}^{-2})^{2p}(d\log d)^{p}\left(\frac{\log n\log l}{nl}\right)^{p}$
+    $\mathbb{E}\left[\left\|{\widehat{v}_{l,h}^{b}}-{v_{l,h}^{b}}\right\|^{2p}\right]\lesssim C(p)(C_{Y}C_{f}R_{0}^{2}\sigma_{\gamma}^{-2})^{2p}(d\log d)^{p}\left(\frac{\log n\log l}{nl}\right)^{p}\ \,.$
     3.2 Estimation of the distance function and classification accuracy
-    Here we assume (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), and ($\bm{\omega_{f}}$): by Corollary 8 we are in “thin” slice scenario, i.e.
-$H_{l,h}>0$, and the distance function simplifies to ${\widehat{\mathrm{dist}}}(x,h)=|\langle x-{\widehat{\mu}_{l,h}^{b}},{\widehat{v}_{l,h}^{b}}\rangle|^{2}+\frac{{\lambda_{d}(\widehat{\Sigma}_{l,h}^{b})}}{{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}}\left\|x-{\widehat{\mu}_{l,h}^{b}}\right\|^{2}$.
-In Algorithm 1, we take the slice ${\widehat{h}_{x}}$, which has the smallest estimated distance, to $x$, as an estimator of the true correct ${h^{\prime}_{x}}$.
-The following proposition states that the population counterpart ${h_{x}}$ of the nearest index almost equals the correct index ${h^{\prime}_{x}}$.
+[…]
     Proposition 10 (nearest index is almost correct)
     Let ${h_{x}}=\operatorname*{{arg\,min}}_{h\in\mathcal{H}_{l}}{\mathrm{dist}}(x,h)$ be the nearest index and define the correct index ${h^{\prime}_{x}}$ be the unique $h\in\mathcal{H}_{l}$ such that $F(x)\in{R_{l,h}}$.
-Suppose ($\bm{\zeta}$ subG), ($\bm{\omega_{f}}$), and (LCV) hold true.
+Suppose ($\bm{\zeta}\in\psi_{2}$), ($\bm{\omega_{f}}$), and (LCV) hold true.
 Suppose that $|{R_{l,h}}|\geqslant\max(\sigma_{\zeta},\omega_{f})$ for all $h\in\mathcal{H}_{l}$.
@@
     Proposition 11 (classification accuracy)
-    Assume (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), and ($\bm{\omega_{f}}$) hold true.
-If $l\gtrsim\frac{C_{f}\mathrm{len}_{\gamma}}{C_{f}^{\prime}\sigma_{\gamma}}$, and $|{R_{l,h}}|\geqslant\max(\sigma,\omega_{f})$ for all $h\in\mathcal{H}_{l}^{b}$, then the probability of misclassification by at least two slices in part 2.b of Algorithm 1 satisfies
+[…]
     $\mathbb{P}\left(\left|{\widehat{h}_{x}}-{h_{x}}\right|\geqslant 2\right)\lesssim ld\exp\left(-c\frac{nl^{-1}}{\sqrt{\tau\log n}}\min\left(\frac{C_{f}^{\prime}\sigma_{\gamma}^{4}}{C_{f}^{2}R_{0}^{3}d^{3/2}\mathrm{len}_{\gamma}},\frac{\sigma_{\gamma}^{8}C_{f}^{\prime 4}}{R_{0}^{8}C_{f}^{4}d^{4}}\right)\right)+ln^{-\tau}\,.$
@@
     One can show that $\widetilde{f}$ has the following explicit formula
-    $\widetilde{f}(s)=\mathbb{E}\left[F(X)\ \big{|}\ X\in{S_{l,h}},\ \langle{v_{l,h}^{b}},X\rangle=s\right]$
+    $\widetilde{f}(s)=\mathbb{E}\left[F(X)\ \big{|}\ X\in{S_{l,h}},\ \langle{v_{l,h}^{b}},X\rangle=s\right]\ \,,$
     and hence $\widetilde{f}$ is Hölder continuous with semi-norm $\left[\widetilde{f}\right]_{\mathcal{C}^{s}}\leqslant\left[f\right]_{\mathcal{C}^{s}}+\frac{C_{Y}R_{0}}{l}|\rho_{X}|_{\mathcal{C}^{s}}$.
@@
     $+\underbrace{\left|\widetilde{f}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{f_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|}_{(B)}+\underbrace{\left|{f_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{\widehat{f}_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|}_{(V)}$
-    $+\underbrace{\left|{\widehat{f}_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{\widehat{f}_{j|{\widehat{v}_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|}_{(\Psi)}$
+    $+\underbrace{\left|{\widehat{f}_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{\widehat{f}_{j|{\widehat{v}_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|}_{(\Psi)}\ \,.$
     Proposition 12 (Mean Squared Error conditioned on almost correct classification)
-    Consider assumptions (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), ($\bm{\omega_{f}}$), and (P).
-Suppose $f\in\mathcal{C}^{s}$ with $s\in\left[\frac{1}{2},2\right]$.
-Conditioned on almost correct classification in part 2.
-b) in Algorithm 1, then we have the following estimates for $n$ is sufficiently large,
-    $\operatorname{\mathrm{MSE}}_{(NCA)}\lesssim$ $\left[f\right]_{\mathcal{C}^{{s\wedge 1}}}^{2}(\sigma_{\gamma}\left\|\gamma^{\prime\prime}\right\|C_{f})^{{2(s\wedge 1)}}\max\left(\sigma_{\zeta},\omega_{f},\frac{\mathrm{len}_{\gamma}}{C_{f}^{\prime}l}\right)^{{2(s\wedge 1)}}$
-    $\operatorname{\mathrm{MSE}}_{(\Phi)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{{s\wedge 1}}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{{s\wedge 1}}})^{2}(\mathrm{len}_{\gamma}R_{0}^{2}\sigma_{\gamma}^{-2})^{{2(s\wedge 1)}}(d^{2}\log d)^{{s\wedge 1}}\left(\frac{\log n\log l}{nl}\right)^{{2(s\wedge 1)}}$
-    $\operatorname{\mathrm{MSE}}_{(B)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})^{2}C_{f}^{2s}\max\left(\sigma_{\zeta},\omega_{f},\frac{\mathrm{len}_{\gamma}}{C_{f}^{\prime}l}\right)^{2s}j^{-2s}$
-    $\operatorname{\mathrm{MSE}}_{(V)}\lesssim$ $\sigma_{\zeta}^{2}\frac{lj\log j}{n}$
+[…]
+    $\operatorname{\mathrm{MSE}}_{(NCA)}\lesssim$ $\left[f\right]_{\mathcal{C}^{{s\wedge 1}}}^{2}(\sigma_{\gamma}\left\|\gamma^{\prime\prime}\right\|C_{f})^{{2(s\wedge 1)}}\max\left(\sigma_{\zeta},\omega_{f},\frac{\mathrm{len}_{\gamma}}{C_{f}^{\prime}l}\right)^{{2(s\wedge 1)}}\ \,;$
+    $\operatorname{\mathrm{MSE}}_{(\Phi)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{{s\wedge 1}}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{{s\wedge 1}}})^{2}(\mathrm{len}_{\gamma}R_{0}^{2}\sigma_{\gamma}^{-2})^{{2(s\wedge 1)}}(d^{2}\log d)^{{s\wedge 1}}\left(\frac{\log n\log l}{nl}\right)^{{2(s\wedge 1)}}\ \,;$
+    $\operatorname{\mathrm{MSE}}_{(B)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})^{2}C_{f}^{2s}\max\left(\sigma_{\zeta},\omega_{f},\frac{\mathrm{len}_{\gamma}}{C_{f}^{\prime}l}\right)^{2s}j^{-2s}\ \,;$
+    $\operatorname{\mathrm{MSE}}_{(V)}\lesssim$ $\sigma_{\zeta}^{2}\frac{lj\log j}{n}\ \,;$
     $\operatorname{\mathrm{MSE}}_{(\Psi)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}\left(\mathrm{reach}_{\gamma}^{2}\frac{C_{f}^{2}C_{Y}^{2}R_{0}^{4}}{\sigma_{\gamma}^{4}}d\log d\frac{\log n\log l}{nl}\right)^{\frac{1}{2-(s\wedge 1)}}$
@@
 In this way we aim at reducing the effect of any errors originating from the estimation of the parameters $({\widehat{\mu}_{l,h}},{\widehat{v}_{l,h}})_{h=1}^{l}$ and from the particular choice $l=l^{*}$ and $j=j^{*}$, thereby imputing the reported “MSE at $n=2\times 10^{6}$” mainly to the nonlinear curve approximation error $\operatorname{\mathrm{MSE}}_{(NCA)}$.
+    Remark: Throughout this section, we test the performance of Algorithm 1 using theoretically predicted parameters $(l^{*},j^{*})$, and we verify in many aspects that the performance in estimation is consistent with the main theorems.
+Nevertheless, it is noteworthy that the theoretical argument focuses mainly on the optimal choice $(l^{*},j^{*})$ which depends on many quantities, including parameters of the underlying curve $\gamma$, that are likely to be unknown in real world applications.
+In practice, the optimal choice for $l,j$ can be found using cross-validation.
+For example, one can apply the 10-fold cross-validation to the training data to extract the optimal value of parameters $l,j$.
+In experiments where we used cross-validation we obtained very similar results to those using theoretically optimal parameters.
     For all figures in this section, we use loglog plots to study how the following quantities decay with the number of samples $n$: the mean squared error $\mathbb{E}[|{\widehat{F}}_{n}(X)-F(X)|^{2}]$, the estimation error of the center along the tangential direction $\mathbb{E}[|\langle{\hat{\mu}}_{l^{*},h}-\mu_{l^{*},h},\gamma^{\prime}_{l^{*},h}\rangle|]$, and the difference between the significant vector and the tangential direction $\mathbb{E}[||{\widehat{v}}_{l^{*},h}-\gamma^{\prime}_{l^{*},h}||]$.
-On the interval $n\in[10^{5},10^{6}]$, we use least squares linear regression to estimate the decaying rate.
+On the interval $n\in[10^{5},10^{6}]$, we use the least squares linear regression to estimate the decaying rates; note that this estimation is not ideal both in the range of small $n$, where our theory may not apply, and for large $n$, due to the error saturation predicted in the additive term in the main theorem.
 We add a dashed vertical line $n=10^{5}$ for those figures to emphasize that the learning rate only corresponds to $n\in[10^{5},10^{6}]$.
@@
 The external noise follows the normal distribution $\mathcal{N}(0,\sigma_{\zeta}^{2})$ with $\sigma_{\zeta}=0.03$.
-Figure 4 supports the following conclusions, consistent with Theorem 3 and our analysis: (i) when $\sigma_{\zeta}>0$, then the mean squared error has a nonzero lower bound independent of $n$; (ii) the curve approximation error is proportional to $\left\|\gamma^{\prime\prime}\right\|^{{2(s\wedge 1)}}$; (iii) Corollary 9 states that the decaying rate is $\mathcal{O}(n^{-\frac{1}{2}})$ for $\mathbb{E}[|\langle{\hat{\mu}}_{l^{*},h}-\mu_{l^{*},h},\gamma^{\prime}_{l^{*},h}\rangle|]$ and $\mathbb{E}[||{\hat{\mu}}_{l^{*},h}-\gamma^{\prime}_{l^{*},h}||]$.
+Figure 4 supports the following conclusions, consistent with Theorem 3 and our analysis: (i) When we have nontrivial external noise $\sigma_{\zeta}>0$, then the mean squared error has a nonzero lower bound independent of $n$; Indeed, in the upper-left plot, we observe that as $n$ increases, the MSE decays at a slower rate and barely changes when $n\approx 10^{6}$; This verifies the existence of constant approximation error.
+Note that the average decaying rate for $n\in[10^{5},10^{6}]$ is estimated using least squares.
+(ii) The curve approximation error is proportional to $\left\|\gamma^{\prime\prime}\right\|^{{2(s\wedge 1)}}$; In fact, in the left plot, we observe that the value of MSE at $n=10^{6}$ is larger for curves with higher curvature and smaller for curves for smaller curvature.
+In the right plot, we further study their relation between the value of MSE when $n=10^{6}$ and the magnitude of curvature.
+We also use the least squares to estimate the growth rate and the estimated growth rate is consistent with the theoretical counterpart.
+(iii) Corollary 9 states that the decaying rate is $\mathcal{O}(n^{-\frac{1}{2}})$ for $\mathbb{E}[|\langle{\hat{\mu}}_{l^{*},h}-\mu_{l^{*},h},\gamma^{\prime}_{l^{*},h}\rangle|]$ and $\mathbb{E}[||{\hat{\mu}}_{l^{*},h}-\gamma^{\prime}_{l^{*},h}||]$.
     4.2 Meyer helix
@@
 Right: Mean Squared Error of ${\widehat{F}}$ at $n=2\times 10^{6}$ as a function of $d$.
-    In the upper-left plot of Figure 6, we observe that, for fixed $n=10^{5}$, the mean squared error is larger for Meyer helix curves with in ambient dimension $d$.
+    In the left plot of Figure 6, we observe that, for fixed $n=10^{5}$, the mean squared error is larger for Meyer helix curves with increasing ambient dimension $d$.
 This is because the coefficient before $n^{-\frac{2s}{2s+1}}$ in Theorem 3 is larger for curves with larger $\mathrm{len}_{\gamma}$.
-In upper right plot, we observe that, by contrast, the mean squared error at $n=2\times 10^{6}$ is smaller for Meyer helix curves in larger ambient dimension $d$.
+In the right plot, we observe that, in contrast, the mean squared error at $n=2\times 10^{6}$ is smaller for Meyer helix curves in larger ambient dimension $d$.
 This verifies the statement in Theorem 3 that the curve approximation error is proportional to $\mathrm{reach}_{\gamma}^{-2{s\wedge 1}}$.
@@
     Proposition 13 ((SC) implies “wide” slices)
-    Suppose (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), and ($\bm{\omega_{f}}$) hold true, together with (SC), with $c_{1},c_{2}$ smaller than a small-enough universal constant.
-Then, for every $l$ such that $|{R_{l,h}}|\asymp\max(\sigma_{\zeta},\omega_{f}),h\in\mathcal{H}_{l}$, we have
-    $\lambda_{1}({\Sigma_{l,h}^{b}})-\lambda_{2}({\Sigma_{l,h}^{b}})\gtrsim C_{f}^{2}|{R_{l,h}}|^{2}$
+[…]
+    $\lambda_{1}({\Sigma_{l,h}^{b}})-\lambda_{2}({\Sigma_{l,h}^{b}})\gtrsim C_{f}^{2}|{R_{l,h}}|^{2}\ \,.$
     Recall that we have the following distance function in this situation of “wide” slices, i.e.
@@
     Proposition 14 (classification accuracy without (LCV))
-    Assume that (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), ($\bm{\omega_{f}}$), and (SC) hold true.
-Fix $l=\frac{C_{Y}R_{0}}{\max(\sigma,\omega_{f})}$.
-Then, the probability of misclassification by at least two slices in part 2.
-b) of Algorithm 1 can be bounded by
-    $\mathbb{P}\left(\left|{\widehat{h}_{x}}-{h_{x}}\right|\geqslant 2\right)\lesssim ld\exp\left(-c\frac{C_{f}^{6}\max(\sigma_{\zeta},\omega_{f})^{7}n}{C_{Y}R_{0}^{7}d^{3}\sqrt{\log n}}\right)+ln^{-\tau}$
+[…]
+    $\mathbb{P}\left(\left|{\widehat{h}_{x}}-{h_{x}}\right|\geqslant 2\right)\lesssim ld\exp\left(-c\frac{C_{f}^{6}\max(\sigma_{\zeta},\omega_{f})^{7}n}{C_{Y}R_{0}^{7}d^{3}\sqrt{\log n}}\right)+ln^{-\tau}\ \,.$
     5.1 Example: Verifying Theorem 5
@@
 Then
-    $\mathbb{P}\left(\left\|\widehat{M}\right\|\gtrsim t\right)\lesssim(d_{1}+d_{2})\exp\left(-\frac{cnt^{2}}{Bt+\sigma^{2}}\right)$
+    $\mathbb{P}\left(\left\|\widehat{M}\right\|\gtrsim t\right)\lesssim(d_{1}+d_{2})\exp\left(-\frac{cnt^{2}}{Bt+\sigma^{2}}\right)\ \,.$
     Lemma 16 (Concentration Inequalities for means, covariances and eigenvalues)
@@
 Then,
-    $\mathbb{P}\left(\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|\gtrsim R_{0}\sqrt{\frac{\delta}{\alpha}}\right)\lesssim d\exp\left(-\frac{c{n_{l,h}^{b}}\delta}{\alpha d}\right),\ \ \text{ where }\delta<\alpha d$
-    $\mathbb{P}\left(\left\|{\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|\gtrsim\beta\frac{R_{0}^{2}}{\alpha}\right)\lesssim d\exp\left(-\frac{c\beta^{2}{n_{l,h}^{b}}}{\alpha^{2}d^{2}}\right),\ \ \text{ where }\beta<\alpha d$
-    $\mathbb{P}\left(\left|{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{1}(\Sigma_{l,h}^{b})}\right|\gtrsim\beta\frac{R_{0}^{2}}{\alpha}\right)\lesssim d\exp\left(-\frac{c\beta^{2}{n_{l,h}^{b}}}{\alpha^{2}d^{2}}\right),\ \ \text{ where }\beta<\alpha d$
+    $\mathbb{P}\left(\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|\gtrsim R_{0}\sqrt{\frac{\delta}{\alpha}}\right)\lesssim d\exp\left(-\frac{c{n_{l,h}^{b}}\delta}{\alpha d}\right),\ \ \text{ where }\delta<\alpha d\ \,;$
+    $\mathbb{P}\left(\left\|{\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|\gtrsim\beta\frac{R_{0}^{2}}{\alpha}\right)\lesssim d\exp\left(-\frac{c\beta^{2}{n_{l,h}^{b}}}{\alpha^{2}d^{2}}\right),\ \ \text{ where }\beta<\alpha d\ \,;$
+    $\mathbb{P}\left(\left|{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{1}(\Sigma_{l,h}^{b})}\right|\gtrsim\beta\frac{R_{0}^{2}}{\alpha}\right)\lesssim d\exp\left(-\frac{c\beta^{2}{n_{l,h}^{b}}}{\alpha^{2}d^{2}}\right),\ \ \text{ where }\beta<\alpha d\ \,.$
     Proof  The first two inequalities can be shown directly by Lemma 15.
@@
     Proof  Define intervals $I_{k}:=(-\sqrt{2(k+1)}\sigma_{\zeta},\sqrt{2k}\sigma_{\zeta}]\cup[\sqrt{2k}\sigma_{\zeta},\sqrt{2(k+1)}\sigma_{\zeta})$ for $k=0,1,2,\dots$.
-We first note that, thanks to ($\bm{\zeta}$ subG), we have $\zeta_{i}\in\bigcup_{k\leqslant\tau\log n}I_{k}$ with probability higher than $1-2n^{-\tau}$, for every $i=1,\dots,n$.
+We first note that, thanks to ($\bm{\zeta}\in\psi_{2}$), we have $\zeta_{i}\in\bigcup_{k\leqslant\tau\log n}I_{k}$ with probability higher than $1-2n^{-\tau}$, for every $i=1,\dots,n$.
 Conditioned on this event and on $Y_{i}\in T$, $\Pi_{\gamma}X_{i}\in f^{-1}(T+\bigcup_{k\leq\tau\log n}I_{k})$.
@@
     $\left|\Pi_{\gamma}X_{i}-\mathbb{E}\left[\Pi_{\gamma}X|Y\in T,\mathcal{B}\right]\right|\leqslant\sum_{k=0}^{\infty}\left|\Pi_{\gamma}X_{i}-\mathbb{E}\left[\Pi_{\gamma}X|Y\in T,\mathcal{B},\zeta\in I_{k}\right]\right|\mathbb{P}(\zeta\in I_{k})$
-    $\lesssim$ $C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta}+\sigma_{\zeta}\sum_{k>\tau\log n}\sqrt{k}e^{-k})\lesssim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})$
+    $\lesssim$ $C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta}+\sigma_{\zeta}\sum_{k>\tau\log n}\sqrt{k}e^{-k})\lesssim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})\ \,.$
     which finishes the proof of (a). For (b), we use the law of total expectation to write
@@
 Following the proof of Proposition 6, we have
-    $\mathbb{P}\left\{\left\|\gamma(\Pi_{\gamma}X_{i})-\mathbb{E}[\gamma(\Pi_{\gamma}X_{i})\mid Y\in T,\mathcal{B}]\right\|\gtrsim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})\mid Y_{i}\in T,\mathcal{B}_{i}\right\}\leqslant 2n^{-\tau}$
-    $\left\|\operatorname{{Cov}}[\gamma(\Pi_{\gamma}X)|Y\in T,\mathcal{B}]\right\|\lesssim C_{f}^{2}(|T|^{2}+\sigma_{\zeta}^{2})$
+    $\mathbb{P}\left\{\left\|\gamma(\Pi_{\gamma}X_{i})-\mathbb{E}[\gamma(\Pi_{\gamma}X_{i})\mid Y\in T,\mathcal{B}]\right\|\gtrsim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})\mid Y_{i}\in T,\mathcal{B}_{i}\right\}\leqslant 2n^{-\tau}\ \,;$
+    $\left\|\operatorname{{Cov}}[\gamma(\Pi_{\gamma}X)|Y\in T,\mathcal{B}]\right\|\lesssim C_{f}^{2}(|T|^{2}+\sigma_{\zeta}^{2})\ \,,$
     and as a consequence,
-    $\mathbb{P}\left\{\left|\langle{v_{l,h}^{b}},\gamma(\Pi_{\gamma}X_{i})\rangle-\mathbb{E}[\langle{v_{l,h}^{b}},\gamma(\Pi_{\gamma}X)\rangle\mid Y\in T,\mathcal{B}]\right|\gtrsim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})\mid Y_{i}\in T,\mathcal{B}_{i}\right\}\leqslant 2n^{-\tau}$
-    $\mathrm{Var}[\langle{v_{l,h}^{b}},\gamma(\Pi_{\gamma}X)\rangle|Y\in T,\mathcal{B}]\lesssim C_{f}^{2}(|T|^{2}+\sigma_{\zeta}^{2})$
+    $\mathbb{P}\left\{\left|\langle{v_{l,h}^{b}},\gamma(\Pi_{\gamma}X_{i})\rangle-\mathbb{E}[\langle{v_{l,h}^{b}},\gamma(\Pi_{\gamma}X)\rangle\mid Y\in T,\mathcal{B}]\right|\gtrsim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})\mid Y_{i}\in T,\mathcal{B}_{i}\right\}\leqslant 2n^{-\tau}\ \,;$
+    $\mathrm{Var}[\langle{v_{l,h}^{b}},\gamma(\Pi_{\gamma}X)\rangle|Y\in T,\mathcal{B}]\lesssim C_{f}^{2}(|T|^{2}+\sigma_{\zeta}^{2})\ \,.$
     By construction, $W_{i}^{\prime}\in\mathbb{R}^{d-1}$ are independent, identical, and centered distribution on $B(0,\mathrm{reach}_{\gamma})\subseteq\mathbb{R}^{d-1}$ (each $W_{i}^{\prime}$ may be dependent of $t_{i}=\Pi_{\gamma}X_{i}$).
@@
     Proposition 17 (local NVM)
-    Suppose (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), and ($\bm{\omega_{f}}$) hold true.
-Let ${\mu_{l,h}^{b}}$ be the mean of the $h$-th slice and ${v_{l,h}^{b}}$ be the significant vector of the $h$-th slice.
-Then, for every $l$ such that $l\gtrsim C_{f}C_{Y}R_{0}/\sigma_{\gamma}$, $|{R_{l,h}^{b}}|\geqslant\max\{\sigma_{\zeta},\omega_{f}\}$ for all $h\in\mathcal{H}_{l}^{b}$, for every $\epsilon\in(0,1)$ and $\tau\geqslant 1$, we have:
+[…]
     1. (a)
@@
 Therefore, we have the following Bernstein-type inequality for any $\epsilon>0$:
-    $\mathbb{P}\left(\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|>\frac{\sigma_{\gamma}^{2}\epsilon}{R_{0}\sqrt{d}}\ \Big{|}\ {n_{l,h}^{b}}\right)\lesssim d\exp\left(-\frac{c\sigma_{\gamma}^{4}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{C_{f}^{2}C_{Y}^{2}R_{0}^{4}d(l^{-2}+l^{-1}\epsilon)}\right)+n^{-\tau}$
+    $\mathbb{P}\left(\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|>\frac{\sigma_{\gamma}^{2}\epsilon}{R_{0}\sqrt{d}}\ \Big{|}\ {n_{l,h}^{b}}\right)\lesssim d\exp\left(-\frac{c\sigma_{\gamma}^{4}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{C_{f}^{2}C_{Y}^{2}R_{0}^{4}d(l^{-2}+l^{-1}\epsilon)}\right)+n^{-\tau}\ \,.$
     Proof of part (b) The main tool is the following Davis-Kahan type inequality.
@@
 5.3, Theorem 5.5) and Corollary 8, gives
-    $\left\|{\widehat{v}_{l,h}^{b}}-{v_{l,h}^{b}}\right\|\leqslant\frac{\left\|({\widehat{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|}{\lambda_{d-1}({\widehat{\Sigma}_{l,h}^{b}})-\lambda_{d}({\Sigma_{l,h}^{b}})}$ (6)
+    $\left\|{\widehat{v}_{l,h}^{b}}-{v_{l,h}^{b}}\right\|\leqslant\frac{\left\|({\widehat{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|}{\lambda_{d-1}({\widehat{\Sigma}_{l,h}^{b}})-\lambda_{d}({\Sigma_{l,h}^{b}})}\ \,.$ (6)
     Step1: We want the denominator of right hand side of (6) to be large.
@@
     $\left\|({\widehat{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|$ $\leqslant\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|+\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|$
-    $\lesssim$ $\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|+R_{0}\sqrt{d}\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|$
+    $\lesssim$ $\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|+R_{0}\sqrt{d}\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|\ \,.$
     Recall that part (a) already gives desired bound for $\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|$, so we only need to bound $\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|$.
@@
     We use $\bm{\mathfrak{S}}_{h}(\epsilon,\beta)$ to denote the event that
-    $\bm{\mathfrak{S}}_{h}=\left\{\begin{split}&\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|<\frac{R_{0}\sqrt{d}}{2},\left|\langle{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}},{v_{l,h}^{b}}\rangle\right|\lesssim\frac{\epsilon\sigma_{\gamma}^{2}}{R_{0}\sqrt{d}},\left\|{v_{l,h}^{b}}-{\widehat{v}_{l,h}^{b}}\right\|\lesssim\epsilon,\\ &\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|<\epsilon\sigma_{\gamma}^{2},\left|\widehat{V}_{l,h}^{b}-{\lambda_{d}(\Sigma_{l,h}^{b})}\right|\lesssim\beta\sigma_{\gamma}^{2}\end{split}\right\}$
+    $\bm{\mathfrak{S}}_{h}=\left\{\begin{split}&\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|<\frac{R_{0}\sqrt{d}}{2},\left|\langle{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}},{v_{l,h}^{b}}\rangle\right|\lesssim\frac{\epsilon\sigma_{\gamma}^{2}}{R_{0}\sqrt{d}},\left\|{v_{l,h}^{b}}-{\widehat{v}_{l,h}^{b}}\right\|\lesssim\epsilon,\\ &\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|<\epsilon\sigma_{\gamma}^{2},\left|\widehat{V}_{l,h}^{b}-{\lambda_{d}(\Sigma_{l,h}^{b})}\right|\lesssim\beta\sigma_{\gamma}^{2}\end{split}\right\}\ \,.$
     We know from part (a), part (b), and Lemma 16 that, $\bm{\mathfrak{S}}_{h}(\epsilon,\beta)$ satisfies, for any $\epsilon,\beta>0$,
     $\mathbb{P}\left(\bm{\mathfrak{S}}_{h}(\epsilon,\beta)^{c}\right)\lesssim$ $d\exp\left(-\frac{c\sigma_{\gamma}^{4}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{C_{f}^{2}C_{Y}^{2}R_{0}^{4}d(l^{-2}+\epsilon l^{-1})}\right)$
-    $+\exp\left(-\frac{c\sigma_{\gamma}^{2}\beta^{2}{n_{l,h}^{b}}}{C_{f}^{2}C_{Y}^{2}R_{0}^{2}l^{-4}(\beta l^{2}+C_{f}^{2}C_{Y}^{2}R_{0}^{2}\sigma_{\gamma}^{-2})}\right)+d\exp\left(-\frac{c\sigma_{\gamma}^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)+n^{-\tau}$
+    $+\exp\left(-\frac{c\sigma_{\gamma}^{2}\beta^{2}{n_{l,h}^{b}}}{C_{f}^{2}C_{Y}^{2}R_{0}^{2}l^{-4}(\beta l^{2}+C_{f}^{2}C_{Y}^{2}R_{0}^{2}\sigma_{\gamma}^{-2})}\right)+d\exp\left(-\frac{c\sigma_{\gamma}^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)+n^{-\tau}\ \,.$
     Conditioned on the event $\bm{\mathfrak{S}}_{h}(\eta,\epsilon,\beta)$, we have the following estimate:
@@
     Now we introduce a term that is an integral of the density $\rho_{t|Y\in{R_{l,h}}}$:
-    $Q_{h}(t_{0}):=\min\left(\mathbb{P}(t\in(0,t_{0})|Y\in{R_{l,h}})\,,\,\mathbb{P}(t\in(t_{0},\mathrm{len}_{\gamma})|Y\in{R_{l,h}})\right)$
+    $Q_{h}(t_{0}):=\min\left(\mathbb{P}(t\in(0,t_{0})|Y\in{R_{l,h}})\,,\,\mathbb{P}(t\in(t_{0},\mathrm{len}_{\gamma})|Y\in{R_{l,h}})\right)\ \,.$
     The same argument shows that $\frac{\max_{h\in\{{h^{\prime}_{x}},{h^{\prime}_{x}}-1\}}Q_{h}(t_{0})}{\max_{h\not\in\{{h^{\prime}_{x}},{h^{\prime}_{x}}-1\}}Q_{h}(t_{0})}\geqslant C(c_{1})$.
@@
     $C_{f}^{2}(|{R_{l,h}}|^{2}+\sigma_{\zeta}^{2})\gtrsim|t_{0}-\mathbb{E}(\Pi_{\gamma}X|Y\in{R_{l,h_{x}}})|^{2}\gtrsim C(c_{1})\min_{h={h^{\prime}_{x}},{h^{\prime}_{x}}-1}|t_{0}-\mathbb{E}(\Pi_{\gamma}X|Y\in{R_{l,h_{x}}})|^{2}$
-    $\gtrsim$ $C(c_{1})\min_{h={h^{\prime}_{x}},{h^{\prime}_{x}}-1}d(x,h)\gtrsim C(c_{1})C_{f}^{\prime 2}(|{R_{l,h}}|^{2}+\sigma_{\zeta}^{2})$
+    $\gtrsim$ $C(c_{1})\min_{h={h^{\prime}_{x}},{h^{\prime}_{x}}-1}d(x,h)\gtrsim C(c_{1})C_{f}^{\prime 2}(|{R_{l,h}}|^{2}+\sigma_{\zeta}^{2})\ \,.$
     which will not hold as long as we properly choose $c_{1}$ such that $C(c_{1})$ is sufficiently large.
@@
     Proof  we know that in small neighborhoods, the curve can be viewed as slightly curved, that is, there exists $K_{0}>0$ depending only on the curvature of $\gamma$ such that for $2\leqslant|k|\leqslant K_{0}l$, we have the following inequality for nearby slices
-    $\sqrt{{\mathrm{dist}}(x,{h_{x}}+k)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim\sum_{k^{\prime}=1}^{|k|}|\Delta\gamma_{{h_{x}}+k^{\prime}}|\cong|k|\mathrm{len}_{\gamma}/l$
+    $\sqrt{{\mathrm{dist}}(x,{h_{x}}+k)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim\sum_{k^{\prime}=1}^{|k|}|\Delta\gamma_{{h_{x}}+k^{\prime}}|\cong|k|\mathrm{len}_{\gamma}/l\ \,.$
     Note that we do not have such inequality for $|k|=1$ because points near the boundary of one slice may share a very similar distance to the adjacent slice, thus hand to distinct true slice index ${h_{x}}$ from adjacent one ${h_{x}}\pm 1$.
@@
 Given $l\gtrsim C_{f}C_{Y}R_{0}$, we deduce that for $|k|\geqslant K_{0}l$, we have the following inequality for far-away slices
-    $\sqrt{{\mathrm{dist}}(x,{h_{x}}+k)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim K_{0}\sqrt{\frac{{\lambda_{d}(\Sigma_{l,h}^{b})}}{{\lambda_{1}(\Sigma_{l,h}^{b})}}}\mathrm{reach}_{\gamma}\gtrsim\frac{\mathrm{len}_{\gamma}}{l}\frac{\mathrm{reach}_{\gamma}}{\sigma_{\gamma}}\gtrsim\mathrm{len}_{\gamma}/l$
+    $\sqrt{{\mathrm{dist}}(x,{h_{x}}+k)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim K_{0}\sqrt{\frac{{\lambda_{d}(\Sigma_{l,h}^{b})}}{{\lambda_{1}(\Sigma_{l,h}^{b})}}}\mathrm{reach}_{\gamma}\gtrsim\frac{\mathrm{len}_{\gamma}}{l}\frac{\mathrm{reach}_{\gamma}}{\sigma_{\gamma}}\gtrsim\mathrm{len}_{\gamma}/l\ \,.$
     As a consequence, we take all $|k|\geqslant 2$ and have the following inequality:
-    $\text{For any }h\text{ such that }|h-{h_{x}}|\geqslant 2,\ \sqrt{{\mathrm{dist}}(x,h)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim\mathrm{len}_{\gamma}/l$
+    $\text{For any }h\text{ such that }|h-{h_{x}}|\geqslant 2,\ \sqrt{{\mathrm{dist}}(x,h)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim\mathrm{len}_{\gamma}/l\ \,.$
     In order to obtain correct classification, we want the estimation error of the distance function to be small, such that for all $|h^{\prime}-{h_{x}}|\geqslant 2$,
-    $\left|{\widehat{\mathrm{dist}}}(x,h^{\prime})-{\mathrm{dist}}(x,h^{\prime})\right|+\left|{\widehat{\mathrm{dist}}}(x,{h_{x}})-{\mathrm{dist}}(x,{h_{x}})\right|<\left|{\mathrm{dist}}(x,h^{\prime})-{\mathrm{dist}}(x,{h_{x}})\right|$ (7)
+    $\left|{\widehat{\mathrm{dist}}}(x,h^{\prime})-{\mathrm{dist}}(x,h^{\prime})\right|+\left|{\widehat{\mathrm{dist}}}(x,{h_{x}})-{\mathrm{dist}}(x,{h_{x}})\right|<\left|{\mathrm{dist}}(x,h^{\prime})-{\mathrm{dist}}(x,{h_{x}})\right|\ \,.$ (7)
     Indeed, this will imply that ${\widehat{h}_{x}}=\operatorname*{{arg\,min}}_{h^{\prime}\in{\mathcal{H}_{l}^{b}}}{\widehat{\mathrm{dist}}}(x,h^{\prime})$ is the correct or adjacent classification, i.e.
@@
     Consider the event $\bm{\mathfrak{S}}$ that we have a small estimation error for information in all slices:
-    $\bm{\mathfrak{S}}(\epsilon,\delta,\beta,u)=\left\{\begin{split}&\text{ For all}h\in{\mathcal{H}_{l}^{b}},\left|\langle{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}},{v_{l,h}^{b}}\rangle\right|\lesssim\frac{\epsilon\sigma_{\gamma}^{2}}{R_{0}\sqrt{d}},\ \left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|\lesssim\sigma_{\gamma}\sqrt{\delta},\\ &\left\|{v_{l,h}^{b}}-{\widehat{v}_{l,h}^{b}}\right\|\lesssim\epsilon,\left|{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{1}(\Sigma_{l,h}^{b})}\right|\lesssim\beta\sigma_{\gamma}^{2},\left|{\lambda_{d}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{d}(\Sigma_{l,h}^{b})}\right|\lesssim\frac{u^{2}C_{f}^{2}C_{Y}^{2}R_{0}^{2}}{l^{2}}\end{split}\right\}$
+    $\bm{\mathfrak{S}}(\epsilon,\delta,\beta,u)=\left\{\begin{split}&\text{ For all}h\in{\mathcal{H}_{l}^{b}},\left|\langle{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}},{v_{l,h}^{b}}\rangle\right|\lesssim\frac{\epsilon\sigma_{\gamma}^{2}}{R_{0}\sqrt{d}},\ \left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|\lesssim\sigma_{\gamma}\sqrt{\delta},\\ &\left\|{v_{l,h}^{b}}-{\widehat{v}_{l,h}^{b}}\right\|\lesssim\epsilon,\left|{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{1}(\Sigma_{l,h}^{b})}\right|\lesssim\beta\sigma_{\gamma}^{2},\left|{\lambda_{d}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{d}(\Sigma_{l,h}^{b})}\right|\lesssim\frac{u^{2}C_{f}^{2}C_{Y}^{2}R_{0}^{2}}{l^{2}}\end{split}\right\}\ \,.$
     Notice that Proposition 17 and Lemma 16 state that, with $l>C_{f}C_{Y}R_{0}/\sigma_{\gamma}$, for any $\delta,\beta<R_{0}^{2}d\sigma_{\gamma}^{-2}$ and any $\epsilon,u>0$, event $\bm{\mathfrak{S}}$ has the following high probability bound:
     $\mathbb{P}\left(\bm{\mathfrak{S}}^{c}\right)\lesssim$ $ld\exp\left(-\frac{c\sigma_{\gamma}^{4}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{C_{f}^{2}C_{Y}^{2}R_{0}^{4}d(l^{-2}+l^{-1}\epsilon)}\right)+ld\exp\left(-c{n_{l,h}^{b}}\min\left(\frac{\sigma_{\gamma}^{2}\delta}{R_{0}^{2}d},\frac{\sigma_{\gamma}^{4}\beta^{2}}{R_{0}^{4}d^{2}},\frac{\sigma_{\gamma}^{4}}{R_{0}^{4}d^{2}}\right)\right)+ln^{-\tau}$
-    $+ld\exp\left(-\frac{c\sigma_{\gamma}^{4}C_{Y}^{2}u^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{4}d^{3/2}(\sqrt{d}+uC_{f}C_{Y})}\right)+l\exp\left(-\frac{c{n_{l,h}^{b}}u^{4}}{1+u^{2}}\right)$
+    $+ld\exp\left(-\frac{c\sigma_{\gamma}^{4}C_{Y}^{2}u^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{4}d^{3/2}(\sqrt{d}+uC_{f}C_{Y})}\right)+l\exp\left(-\frac{c{n_{l,h}^{b}}u^{4}}{1+u^{2}}\right)\ \,.$
     We now investigate how small these parameters should be.
@@
     $\lesssim$ $\left(R_{0}\sqrt{d}\epsilon+\sigma_{\gamma}\sqrt{\delta}\epsilon+\frac{\sigma_{\gamma}^{2}\epsilon}{R_{0}\sqrt{d}}\right)^{2}+\left(R_{0}\sqrt{d}\epsilon+\sigma_{\gamma}\sqrt{\delta}\epsilon+\frac{\sigma_{\gamma}^{2}\epsilon}{R_{0}\sqrt{d}}\right)\sqrt{{\mathrm{dist}}(x,h)}$
-    $+\frac{{\lambda_{d}(\Sigma_{l,h}^{b})}}{{\lambda_{1}(\Sigma_{l,h}^{b})}}\left(\sigma_{\gamma}^{2}\delta+R_{0}\sigma_{\gamma}\sqrt{d}\sqrt{\delta}\right)+\frac{u^{2}\mathrm{len}_{\gamma}^{2}}{l^{2}}\frac{R_{0}^{2}d}{{\lambda_{1}(\Sigma_{l,h}^{b})}}+\beta\sigma_{\gamma}^{2}\frac{{\lambda_{d}(\Sigma_{l,h}^{b})}R_{0}^{2}d}{{\lambda_{1}(\Sigma_{l,h}^{b})}^{2}}$
+    $+\frac{{\lambda_{d}(\Sigma_{l,h}^{b})}}{{\lambda_{1}(\Sigma_{l,h}^{b})}}\left(\sigma_{\gamma}^{2}\delta+R_{0}\sigma_{\gamma}\sqrt{d}\sqrt{\delta}\right)+\frac{u^{2}\mathrm{len}_{\gamma}^{2}}{l^{2}}\frac{R_{0}^{2}d}{{\lambda_{1}(\Sigma_{l,h}^{b})}}+\beta\sigma_{\gamma}^{2}\frac{{\lambda_{d}(\Sigma_{l,h}^{b})}R_{0}^{2}d}{{\lambda_{1}(\Sigma_{l,h}^{b})}^{2}}\ \,.$
     Thus in above inequality, we want the coefficient before $\sqrt{{\mathrm{dist}}(x,h)}$ to be smaller than $c\frac{\mathrm{len}_{\gamma}}{l}$, and all other terms to be smaller than $c\frac{\mathrm{len}_{\gamma}^{2}}{l^{2}}$, so that (7) can be guaranteed.
@@
 We are going to control the curve approximation error
-    $\operatorname{\mathrm{MSE}}_{(NCA)}:=\mathbb{E}\left[\left|f(\Pi_{\gamma}X)-\widetilde{f}(\langle{v_{l,h}^{b}},X\rangle)\right|^{2}\mathbbm{1}_{I^{(l,h)}}(\langle{v_{l,h}^{b}},X\rangle)\ \big{|}\ X\in{S_{l,h}}\right]$
+    $\operatorname{\mathrm{MSE}}_{(NCA)}:=\mathbb{E}\left[\left|f(\Pi_{\gamma}X)-\widetilde{f}(\langle{v_{l,h}^{b}},X\rangle)\right|^{2}\mathbbm{1}_{I^{(l,h)}}(\langle{v_{l,h}^{b}},X\rangle)\ \big{|}\ X\in{S_{l,h}}\right]\ \,.$
     Denote $t_{1}\in\Pi_{\gamma}({S_{l,h}})$ to be such that $\gamma^{\prime}(t_{1})={v_{l,h}^{b}}$.
@@
     we obtain an upper-bound upon replacing $\widetilde{f}(\cdot)$ by $f(\cdot+c(h))$ By Hölder continuity, the triangle inequality, and Taylor expansion, we have
-    $\begin{split}\operatorname{\mathrm{MSE}}_{(NCA)}\leqslant&\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\mathbb{E}\left[\left|\Pi_{\gamma}X-\langle X,{v_{l,h}^{b}}\rangle-c(h)\right|^{2(s\wedge 1)}\ \big{|}\ X\in{S_{l,h}}\right]\\ \leqslant&C(s)\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\mathbb{E}\left[\left|\langle\gamma(\Pi_{\gamma}X)-\gamma(t_{1}),{v_{l,h}^{b}}\rangle-(\Pi_{\gamma}X-t_{1})\right|^{2(s\wedge 1)}\ \big{|}\ X\in{S_{l,h}}\right]\\ &+C(s)\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\mathbb{E}\left[\left|(Z_{d-1},0)^{\intercal}M_{\gamma^{\prime}(t)}^{\intercal}({v_{l,h}^{b}}-\gamma^{\prime}(t))\right|^{2(s\wedge 1)}\ \big{|}\ X\in{S_{l,h}}\right]\\ \lesssim&\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\sigma_{\gamma}^{2(s\wedge 1)}\left\|\gamma^{\prime\prime}\right\|^{2(s\wedge 1)}C_{f}^{2(s\wedge 1)}\max\left(\sigma_{\zeta},\omega_{f},\frac{\mathrm{len}_{\gamma}}{C_{f}^{\prime}l}\right)^{2(s\wedge 1)}\end{split}\,.$ (NL)
+    $\begin{split}\operatorname{\mathrm{MSE}}_{(NCA)}\leqslant&\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\mathbb{E}\left[\left|\Pi_{\gamma}X-\langle X,{v_{l,h}^{b}}\rangle-c(h)\right|^{2(s\wedge 1)}\ \big{|}\ X\in{S_{l,h}}\right]\\ \leqslant&C(s)\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\mathbb{E}\left[\left|\langle\gamma(\Pi_{\gamma}X)-\gamma(t_{1}),{v_{l,h}^{b}}\rangle-(\Pi_{\gamma}X-t_{1})\right|^{2(s\wedge 1)}\ \big{|}\ X\in{S_{l,h}}\right]\\ &+C(s)\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\mathbb{E}\left[\left|(Z_{d-1},0)^{\intercal}M_{\gamma^{\prime}(t)}^{\intercal}({v_{l,h}^{b}}-\gamma^{\prime}(t))\right|^{2(s\wedge 1)}\ \big{|}\ X\in{S_{l,h}}\right]\\ \lesssim&\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\sigma_{\gamma}^{2(s\wedge 1)}\left\|\gamma^{\prime\prime}\right\|^{2(s\wedge 1)}C_{f}^{2(s\wedge 1)}\max\left(\sigma_{\zeta},\omega_{f},\frac{\mathrm{len}_{\gamma}}{C_{f}^{\prime}l}\right)^{2(s\wedge 1)}\ \,.\end{split}$ (NL)
     We use Hölder continuity property, Cauchy-Schwarz inequality, and Corollary 9 to control the following error term
@@
     $\mathbb{E}\left[\left|{f_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{\widehat{f}_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|^{2}\right]\lesssim\sigma_{\zeta}^{2}\frac{j\log j}{n/l}\,.$ (V)
-    $\begin{split}\operatorname{\mathrm{MSE}}_{(\Psi)}&:=\mathbb{E}\left[\left|{\widehat{f}_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{\widehat{f}_{j|{\widehat{v}_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|^{2}\right]\end{split}$ ($\Psi$)
+    $\begin{split}\operatorname{\mathrm{MSE}}_{(\Psi)}&:=\mathbb{E}\left[\left|{\widehat{f}_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{\widehat{f}_{j|{\widehat{v}_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|^{2}\right]\end{split}\ \,.$ ($\Psi$)
     To control the projection error term $\left|{\widehat{f}_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{\widehat{f}_{j|{\widehat{v}_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|$, we use Lemma 23.
@@
     Then we have zero variance $\operatorname{\mathrm{MSE}}_{(V)}=0$ and can control the following errors:
-    $\operatorname{\mathrm{MSE}}_{(NCA)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}\left(\frac{\sigma_{\gamma}^{2}C_{f}^{2}\mathrm{len}_{\gamma}^{2}C_{\gamma,f}^{2}}{\mathrm{reach}_{\gamma}^{2}C_{f}^{\prime 2}}\frac{\log^{3}n}{n^{2}}\right)^{s\wedge 1}$
-    $\operatorname{\mathrm{MSE}}_{(\Phi)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}\left(\frac{\mathrm{len}_{\gamma}R_{0}^{2}}{\sigma_{\gamma}^{2}}\right)^{2(s\wedge 1)}(d^{2}\log d)^{s\wedge 1}C_{\gamma,f}^{2(s\wedge 1)}\frac{(\log n)^{7(s\wedge 1)}}{n^{4(s\wedge 1)}}$
-    $\operatorname{\mathrm{MSE}}_{(B)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}(C_{f}C_{Y}R_{0})^{2s}C_{\gamma,f}^{2s}\frac{(\log n)^{3s}}{n^{2s}}$
+    $\operatorname{\mathrm{MSE}}_{(NCA)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}\left(\frac{\sigma_{\gamma}^{2}C_{f}^{2}\mathrm{len}_{\gamma}^{2}C_{\gamma,f}^{2}}{\mathrm{reach}_{\gamma}^{2}C_{f}^{\prime 2}}\frac{\log^{3}n}{n^{2}}\right)^{s\wedge 1}\ \,;$
+    $\operatorname{\mathrm{MSE}}_{(\Phi)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}\left(\frac{\mathrm{len}_{\gamma}R_{0}^{2}}{\sigma_{\gamma}^{2}}\right)^{2(s\wedge 1)}(d^{2}\log d)^{s\wedge 1}C_{\gamma,f}^{2(s\wedge 1)}\frac{(\log n)^{7(s\wedge 1)}}{n^{4(s\wedge 1)}}\ \,;$
+    $\operatorname{\mathrm{MSE}}_{(B)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}(C_{f}C_{Y}R_{0})^{2s}C_{\gamma,f}^{2s}\frac{(\log n)^{3s}}{n^{2s}}\ \,;$
     $\operatorname{\mathrm{MSE}}_{(\Psi)}\lesssim$ $(\left[f\right]_{\mathcal{C}^{s\wedge 1}}+\left|f\right|_{L^{\infty}}\left[\rho_{X}\right]_{\mathcal{C}^{s\wedge 1}})^{2}(C_{f}C_{Y}R_{0})^{2(s\wedge 1)}C_{\gamma,f}^{s\wedge 1}(\log n)^{3(s\wedge 1)}n^{-2(s\wedge 1)}$
@@
 It follows that the optimal way is to let product $l^{*}j^{*}$ grow proportional to
-    $n^{\frac{1}{2s+1}}M^{*}\text{ where }M^{*}=\left(\sigma_{\zeta}^{-1}(C_{f}C_{Y}R_{0})^{s}(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})\right)^{\frac{2}{2s+1}}$
+    $n^{\frac{1}{2s+1}}M^{*}\text{ where }M^{*}=\left(\sigma_{\zeta}^{-1}(C_{f}C_{Y}R_{0})^{s}(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})\right)^{\frac{2}{2s+1}}\ \,.$
     In practice, when the number of samples is not quite sufficient, we take $j^{*}=C$ and let $l^{*}$ grow with $n$.
@@
 Also, the optimal $\operatorname{\mathrm{MSE}}_{(B)}+\operatorname{\mathrm{MSE}}_{(V)}$ is
-    $\frac{C_{f}^{\frac{2s}{2s+1}}}{C_{f}^{\prime\frac{2s}{2s+1}}}\mathrm{len}_{\gamma}^{\frac{2s}{2s+1}}\sigma_{\zeta}^{{}^{\frac{4s}{2s+1}}}(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})^{\frac{2}{2s+1}}n^{-\frac{2s}{2s+1}}{\log n}$ (8)
+    $\frac{C_{f}^{\frac{2s}{2s+1}}}{C_{f}^{\prime\frac{2s}{2s+1}}}\mathrm{len}_{\gamma}^{\frac{2s}{2s+1}}\sigma_{\zeta}^{{}^{\frac{4s}{2s+1}}}(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})^{\frac{2}{2s+1}}n^{-\frac{2s}{2s+1}}{\log n}\ \,.$ (8)
     Therefore, we let $\epsilon_{0}^{2}$ to be the maximum of above two errors, that is,
-    $\epsilon_{0}^{2}=\max\left(\frac{\left[f\right]_{\mathcal{C}^{s}}^{2}\sigma_{\gamma}^{2s}C_{f}^{2s}\max(\sigma_{\zeta},\omega_{f})^{2s}}{\mathrm{reach}_{\gamma}^{2s}},\frac{C_{f}^{\frac{2s}{2s+1}}\mathrm{len}_{\gamma}^{\frac{2s}{2s+1}}\sigma_{\zeta}^{{}^{\frac{4s}{2s+1}}}(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})^{\frac{2}{2s+1}}\log n}{C_{f}^{\prime\frac{2s}{2s+1}}n^{\frac{2s}{2s+1}}}\right)$
+    $\epsilon_{0}^{2}=\max\left(\frac{\left[f\right]_{\mathcal{C}^{s}}^{2}\sigma_{\gamma}^{2s}C_{f}^{2s}\max(\sigma_{\zeta},\omega_{f})^{2s}}{\mathrm{reach}_{\gamma}^{2s}},\frac{C_{f}^{\frac{2s}{2s+1}}\mathrm{len}_{\gamma}^{\frac{2s}{2s+1}}\sigma_{\zeta}^{{}^{\frac{4s}{2s+1}}}(\left[f\right]_{\mathcal{C}^{s}}+C_{Y}R_{0}l^{-1}\left[\rho_{X}\right]_{\mathcal{C}^{s}})^{\frac{2}{2s+1}}\log n}{C_{f}^{\prime\frac{2s}{2s+1}}n^{\frac{2s}{2s+1}}}\right)\ \,.$
     Moreover, when $n$ is sufficiently large the error denoted by $\epsilon_{0}^{2}$ dominates all other errors: the residual error in the convergence of $\operatorname{\mathrm{MSE}}_{(NCA)}$, as well as the terms $\operatorname{\mathrm{MSE}}_{(\Phi)}$ and $\operatorname{\mathrm{MSE}}_{(\Psi)}$, become negligible high order terms compared with $\epsilon_{0}^{2}$, as soon as the standing assumption $\frac{n}{\log^{{3/2}}n}\gtrsim C_{\gamma,f}\frac{C_{f}\mathrm{len}_{\gamma}}{C_{f}^{\prime}\sigma_{\gamma}}$ is satisfied.
@@
     Proposition 18 (local NVM for “wide” slice)
-    Suppose (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (SC), and ($\bm{\omega_{f}}$) hold true.
-Let ${\mu_{l,h}^{b}}$ be the mean of $h$-th slice and ${v_{l,h}^{b}}$ be the significant vector of $h$-th slice.
-Then, for every $l$ such that $|{R_{l,h}^{b}}|\asymp\max(\sigma_{\zeta},\omega_{f})$ for all $h\in\mathcal{H}_{l}^{b}$, for every $\epsilon\in(0,1)$ and $\tau\geqslant 1$, on each slice
+[…]
     1. (a)
       For any $h\in\mathcal{H}_{l}$ and any $\epsilon>0$, the estimation error of the slice mean along the tangential direction can be bounded as
-      $\mathbb{P}\left\{\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|>\frac{C_{f}^{2}|{R_{l,h}^{b}}|^{2}\epsilon}{R_{0}\sqrt{d}}\ \bigg{|}\ {n_{l,h}^{b}}\right\}\lesssim d\exp\left(-\frac{cC_{f}^{2}|{R_{l,h}^{b}}|^{2}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{2}d+\epsilon C_{f}|{R_{l,h}^{b}}|R_{0}\sqrt{d}}\right)+n^{-\tau}$
+      $\mathbb{P}\left\{\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|>\frac{C_{f}^{2}|{R_{l,h}^{b}}|^{2}\epsilon}{R_{0}\sqrt{d}}\ \bigg{|}\ {n_{l,h}^{b}}\right\}\lesssim d\exp\left(-\frac{cC_{f}^{2}|{R_{l,h}^{b}}|^{2}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{2}d+\epsilon C_{f}|{R_{l,h}^{b}}|R_{0}\sqrt{d}}\right)+n^{-\tau}\ \,.$
     2. (b)
       For any $h\in\mathcal{H}_{l}$, the estimation error of the significant vector can be bounded as,
-      $\mathbb{P}\left\{\left\|{{\widehat{v}_{l,h}^{b}}}-{{v_{l,h}^{b}}}\right\|>\epsilon\ \big{|}\ {n_{l,h}^{b}}\right\}\lesssim d\exp\left(-\frac{cC_{f}^{2}|{R_{l,h}^{b}}|^{2}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{2}d+\epsilon C_{f}|{R_{l,h}^{b}}|R_{0}\sqrt{d}}\right)+d\exp\left(-\frac{cC_{f}^{4}|{R_{l,h}}|^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)+n^{-\tau}$
+      $\mathbb{P}\left\{\left\|{{\widehat{v}_{l,h}^{b}}}-{{v_{l,h}^{b}}}\right\|>\epsilon\ \big{|}\ {n_{l,h}^{b}}\right\}\lesssim d\exp\left(-\frac{cC_{f}^{2}|{R_{l,h}^{b}}|^{2}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{2}d+\epsilon C_{f}|{R_{l,h}^{b}}|R_{0}\sqrt{d}}\right)+d\exp\left(-\frac{cC_{f}^{4}|{R_{l,h}}|^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)+n^{-\tau}\ \,.$
     3. (c)
@@
       $\mathbb{P}\left(\left|{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{1}(\Sigma_{l,h}^{b})}\right|>u^{2}C_{f}^{2}|{R_{l,h}^{b}}|^{2}\ \bigg{|}\ {n_{l,h}^{b}}\right)$
-      $\qquad\qquad\lesssim d\exp\left(-c\frac{u^{2}C_{f}^{4}|{R_{l,h}^{b}}|^{4}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{4}d^{2}+uC_{f}^{2}|{R_{l,h}^{b}}|^{2}R_{0}^{2}d}\right)+d\exp\left(-\frac{cC_{f}^{4}|{R_{l,h}}|^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)+n^{-\tau}$
+      $\qquad\qquad\lesssim d\exp\left(-c\frac{u^{2}C_{f}^{4}|{R_{l,h}^{b}}|^{4}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{4}d^{2}+uC_{f}^{2}|{R_{l,h}^{b}}|^{2}R_{0}^{2}d}\right)+d\exp\left(-\frac{cC_{f}^{4}|{R_{l,h}}|^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)+n^{-\tau}\ \,.$
     Moreover, we have the following probability bound For fixed constant $\beta<\frac{R_{0}^{2}d}{C_{f}^{2}|{R_{l,h}}|^{2}}$,
-    $\mathbb{P}\left(\max\left(\left\|{\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|,\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|^{2}\right)\gtrsim\beta C_{f}^{2}|{R_{l,h}^{b}}|^{2}\right)\lesssim d\exp\left(-\frac{c\beta^{2}C_{f}^{4}|{R_{l,h}^{b}}|^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)$
+    $\mathbb{P}\left(\max\left(\left\|{\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|,\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|^{2}\right)\gtrsim\beta C_{f}^{2}|{R_{l,h}^{b}}|^{2}\right)\lesssim d\exp\left(-\frac{c\beta^{2}C_{f}^{4}|{R_{l,h}^{b}}|^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)\ \,.$
     Consider the event $\bm{\mathfrak{S}}$ that we have a small estimation error for information in all slices:
-    $\bm{\mathfrak{S}}(\epsilon,\beta,u)=\left\{\begin{split}&\text{ For all }h\in{\mathcal{H}_{l}^{b}},\left|\langle{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}},{v_{l,h}^{b}}\rangle\right|\lesssim\frac{\epsilon|{R_{l,h}^{b}}|^{2}\epsilon}{R_{0}\sqrt{d}},\ \left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|\lesssim\frac{R_{0}\sqrt{d}}{2},\left\|{v_{l,h}^{b}}-{\widehat{v}_{l,h}^{b}}\right\|\lesssim\epsilon\\ &\left|{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{1}(\Sigma_{l,h}^{b})}\right|\lesssim u^{2}C_{f}^{2}|{R_{l,h}^{b}}|^{2},\max\left(\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|^{2},\left\|{\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|\right)\lesssim\beta^{2}C_{f}^{2}|{R_{l,h}^{b}}|^{2}\end{split}\right\}$
+    $\bm{\mathfrak{S}}(\epsilon,\beta,u)=\left\{\begin{split}&\text{ For all }h\in{\mathcal{H}_{l}^{b}},\left|\langle{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}},{v_{l,h}^{b}}\rangle\right|\lesssim\frac{\epsilon|{R_{l,h}^{b}}|^{2}\epsilon}{R_{0}\sqrt{d}},\ \left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|\lesssim\frac{R_{0}\sqrt{d}}{2},\left\|{v_{l,h}^{b}}-{\widehat{v}_{l,h}^{b}}\right\|\lesssim\epsilon\\ &\left|{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{1}(\Sigma_{l,h}^{b})}\right|\lesssim u^{2}C_{f}^{2}|{R_{l,h}^{b}}|^{2},\max\left(\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|^{2},\left\|{\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}\right\|\right)\lesssim\beta^{2}C_{f}^{2}|{R_{l,h}^{b}}|^{2}\end{split}\right\}\ \,.$
     We claim that the conditions $\epsilon^{\prime}\asymp\beta^{\prime}\asymp\frac{C_{f}^{2}|{R_{l,h}^{b}}|^{2}}{R_{0}^{2}d}$ and $u\asymp\frac{C_{f}|{R_{l,h}^{b}}|}{R_{0}\sqrt{d}}$ are sufficient to perform almost correct classification.
     Recall that similar to the “thin” slice scenario, in order to obtain correct classification, we want the estimation error of the distance function to be small, such that for all $|h^{\prime}-{h_{x}}|\geqslant 2$,
-    $\left|{\widehat{\mathrm{dist}}}(x,h^{\prime})-{\mathrm{dist}}(x,h^{\prime})\right|+\left|{\widehat{\mathrm{dist}}}(x,{h_{x}})-{\mathrm{dist}}(x,{h_{x}})\right|<\left|{\mathrm{dist}}(x,h^{\prime})-{\mathrm{dist}}(x,{h_{x}})\right|$
+    $\left|{\widehat{\mathrm{dist}}}(x,h^{\prime})-{\mathrm{dist}}(x,h^{\prime})\right|+\left|{\widehat{\mathrm{dist}}}(x,{h_{x}})-{\mathrm{dist}}(x,{h_{x}})\right|<\left|{\mathrm{dist}}(x,h^{\prime})-{\mathrm{dist}}(x,{h_{x}})\right|\ \,.$
     Indeed, this will imply that ${\widehat{h}_{x}}=\operatorname*{{arg\,min}}_{h^{\prime}\in{\mathcal{H}_{l}^{b}}}{\widehat{\mathrm{dist}}}(x,h^{\prime})$ is the correct or adjacent classification, i.e.
@@
     Now, we are going to analyze the distance function. Notice that in the “wide” slice scenario, the difference in distance can be bounded as follows:
-    $\sqrt{{\mathrm{dist}}(x,{h_{x}}+k)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim C_{f}|{R_{l,h}^{b}}|\ \ \text{ for any }|k|\geqslant 2$
+    $\sqrt{{\mathrm{dist}}(x,{h_{x}}+k)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim C_{f}|{R_{l,h}^{b}}|\ \ \text{ for any }|k|\geqslant 2\ \,.$
     (Recall that for large $k$, we can bound this by $\mathrm{reach}_{\gamma}\gtrsim C_{f}|{R_{l,h}^{b}}|$ by assumption (SC).) Thus this gives a lower bound for the right-hand side.
@@
     $\left|{\widehat{\mathrm{dist}}}(x,h)-{\mathrm{dist}}(x,h)\right|$
-    $\lesssim$ $C_{f}^{2}|{R_{l,h}^{b}}|^{2}\beta^{2}+C_{f}|{R_{l,h}^{b}}|\beta\sqrt{{\mathrm{dist}}(x,h)}+u^{2}R_{0}^{2}d+\epsilon^{2}R_{0}^{2}d+C_{f}^{2}|{R_{l,h}^{b}}|^{2}\beta^{2}\epsilon^{2}+\frac{C_{f}^{4}|{R_{l,h}^{b}}|^{4}}{R_{0}^{2}d}\epsilon^{2}\lesssim C_{f}^{2}|{R_{l,h}^{b}}|^{2}$
+    $\lesssim$ $C_{f}^{2}|{R_{l,h}^{b}}|^{2}\beta^{2}+C_{f}|{R_{l,h}^{b}}|\beta\sqrt{{\mathrm{dist}}(x,h)}+u^{2}R_{0}^{2}d+\epsilon^{2}R_{0}^{2}d+C_{f}^{2}|{R_{l,h}^{b}}|^{2}\beta^{2}\epsilon^{2}+\frac{C_{f}^{4}|{R_{l,h}^{b}}|^{4}}{R_{0}^{2}d}\epsilon^{2}\lesssim C_{f}^{2}|{R_{l,h}^{b}}|^{2}\ \,,$
     when $\epsilon^{\prime}\asymp\beta^{\prime}\asymp\frac{C_{f}^{2}|{R_{l,h}^{b}}|^{2}}{R_{0}^{2}d}$ and $u\asymp\frac{C_{f}|{R_{l,h}^{b}}|}{R_{0}\sqrt{d}}$.
 Therefore, we derive the conclusion that the event of misclassification by at least two slices has a small probability:
-    $\mathbb{P}\left(\left|{\widehat{h}_{x}}-{h_{x}}\right|\geqslant 2\right)\lesssim ld\exp\left(-c\frac{C_{f}^{6}\max(\sigma_{\zeta},\omega_{f})^{7}n}{C_{Y}R_{0}^{7}d^{3}\sqrt{\log n}}\right)+ln^{-\tau}$
+    $\mathbb{P}\left(\left|{\widehat{h}_{x}}-{h_{x}}\right|\geqslant 2\right)\lesssim ld\exp\left(-c\frac{C_{f}^{6}\max(\sigma_{\zeta},\omega_{f})^{7}n}{C_{Y}R_{0}^{7}d^{3}\sqrt{\log n}}\right)+ln^{-\tau}\ \,.$
     This finishes the proof of Proposition 14. Theorem 5 follows easily because we mainly utilize the high accuracy of classification without performing one-dimensional regression.
@@
 Then
-    $\mathbb{P}\left\{|\hat{\rho}(E)-\rho(E)|>t\right\}\leqslant 2\exp\left\{-\frac{nt^{2}/2}{\rho(E)+t/3}\right\}$
+    $\mathbb{P}\left\{|\hat{\rho}(E)-\rho(E)|>t\right\}\leqslant 2\exp\left\{-\frac{nt^{2}/2}{\rho(E)+t/3}\right\}\ \,.$
     In particular, for $t=\rho(E)/2$, we have
-    $\mathbb{P}\left\{\hat{\rho}(E)\not\in\left[\frac{1}{2}\rho(E),\frac{3}{2}\rho(E)\right]\right\}\leqslant\mathbb{P}\left\{|\hat{\rho}(E)-\rho(E)|>\frac{1}{2}\rho(E)\right\}\leqslant 2\exp(-\frac{3}{28}n\rho(E))$
+    $\mathbb{P}\left\{\hat{\rho}(E)\not\in\left[\frac{1}{2}\rho(E),\frac{3}{2}\rho(E)\right]\right\}\leqslant\mathbb{P}\left\{|\hat{\rho}(E)-\rho(E)|>\frac{1}{2}\rho(E)\right\}\leqslant 2\exp(-\frac{3}{28}n\rho(E))\ \,.$
     Lemma 20
@@
     Lemma 23 ((Lanteri et al., 2022, Proposition 4 ))
-    Assume (X subG), ($\bm{\zeta}$ subG), and (P).
-Suppose function $\widetilde{f}\in\mathcal{C}^{s}$ with $s\in[\frac{1}{2},1]$ is defined on interval $I$ with diameter $r\geqslant 1$.
-Let integer $m=\lfloor s\rfloor$ be the largest integer smaller than or equal to $s$, so $m=0$ if $s\in[\frac{1}{2},1)$ and $m=1$ if $s=1$.
-Given $j\geqslant 1$, construct $\{I_{j,k}\}_{k=1}^{j}$ the uniform partition of $I$ into $j$ intervals.
-Let ${\widehat{v}}\in\mathbb{S}^{d-1}$ be an estimator of $v\in\mathbb{S}^{d-1}$ such that $\langle{\widehat{v}},v\rangle\geqslant 0$.
-Given data $\{(X_{i}^{\prime},Y_{i}^{\prime})\}_{i=1}^{n^{\prime}}$.
-For $u\in\{v,{\widehat{v}}\}$, denote $n_{j,k|u}=\#\{X_{i}^{\prime}:\langle u,X_{i}^{\prime}\rangle\in I_{j,k}\}$.
-For each $k\in\mathcal{K}_{j|u}:=\{k:n_{j,k|u}\geqslant n^{\prime}/j\}$, compute
-    ${\widehat{f}_{j,k|u}}=\operatorname*{{arg\,min}}_{\deg(p)\leqslant m}\sum_{i\leqslant n^{\prime}}\left|Y_{i}-p(\langle u,X_{i}^{\prime}\rangle)\right|^{2}\mathbbm{1}_{I_{j,k}}(\langle u,X_{i}^{\prime}\rangle)$
-    and compute the piecewise constant $(s<1)$ or linear $(s=1)$ estimator of $\widetilde{f}$ at scale $j$ conditioned on $u$ and truncated at $\left|\widetilde{f}\right|_{L^{\infty}}$
-    ${\widehat{f}_{j|u}}(r)=\sum_{k\in\mathcal{K}_{j}}{\widehat{f}_{j,k|u}}(r)\mathbbm{1}_{I_{j,k}}(r)$
+[…]
+    ${\widehat{f}_{j,k|u}}=\operatorname*{{arg\,min}}_{\deg(p)\leqslant m}\sum_{i\leqslant n^{\prime}}\left|Y_{i}-p(\langle u,X_{i}^{\prime}\rangle)\right|^{2}\mathbbm{1}_{I_{j,k}}(\langle u,X_{i}^{\prime}\rangle)\ \,,$
+    and compute the piecewise constant $(s<1)$ or linear $(s=1)$ estimator of $\widetilde{f}$ at scale $j$ conditioned on $u$ and truncated at $\left|\widetilde{f}\right|_{L^{\infty}}$:
+    ${\widehat{f}_{j|u}}(r)=\sum_{k\in\mathcal{K}_{j}}{\widehat{f}_{j,k|u}}(r)\mathbbm{1}_{I_{j,k}}(r)\ \,.$
     Conditioned on $\left\|{\widehat{v}}-v\right\|\leqslant t/j$ for some $t\geqslant 1$. For every $\epsilon_{0}>0$, we have
-    $\mathbb{E}_{X}\left[\left|{\widehat{f}_{j,k|v}}(\langle v,X\rangle)-{\widehat{f}_{j,k|{\widehat{v}}}}(\langle{\widehat{v}},X\rangle)\right|^{2}\mathbbm{1}_{B(0,r)}(X)\right]\lesssim\left[\widetilde{f}\right]_{\mathcal{C}^{s}}^{2}t^{2}r^{\frac{2}{2-s}}\left\|{\widehat{v}}-v\right\|^{\frac{2}{2-s}}+\epsilon_{0}^{2}$
-    with probability higher than $1-C\#\mathcal{K}_{j}\exp\left(-\frac{cn^{\prime}\epsilon_{0}^{2}}{\#\mathcal{K}_{j}\left[\widetilde{f}\right]_{\mathcal{C}^{s}}^{2}t^{2}r^{2s}}\right)$
+    $\mathbb{E}_{X}\left[\left|{\widehat{f}_{j,k|v}}(\langle{\widehat{v}},X\rangle)-{\widehat{f}_{j,k|{\widehat{v}}}}(\langle{\widehat{v}},X\rangle)\right|^{2}\mathbbm{1}_{B(0,r)}(X)\right]\lesssim\left[\widetilde{f}\right]_{\mathcal{C}^{s}}^{2}t^{2}r^{\frac{2}{2-s}}\left\|{\widehat{v}}-v\right\|^{\frac{2}{2-s}}+\epsilon_{0}^{2}\ \,.$
+    with probability higher than $1-C\#\mathcal{K}_{j}\exp\left(-\frac{cn^{\prime}\epsilon_{0}^{2}}{\#\mathcal{K}_{j}\left[\widetilde{f}\right]_{\mathcal{C}^{s}}^{2}t^{2}r^{2s}}\right)$.
+    Note that the assumption ($\bm{\text{P}_{\gamma}}$) uses the push-forward of the distribution of $X$ along the map $x\mapsto\Pi_{\gamma}(x)$, while in this Lemma we use the push-forward along the orthogonal projection onto the tangent vector at $\Pi_{\gamma}(x)$.
+This second measure however does inherit a lower-bound on its density by the regularity of the curve and of the two maps, the second one being a local linearization of the first one.
     Appendix C Case Analysis: Meyer helix
@@
 Explicitly, we take the square root of the Gaussian density centered at $\mu\in I$ and truncate it to $t\in I$:
-    $\begin{split}I\to L^{2}(I)\quad:\quad\mu\mapsto g_{\mu}(t):=\frac{1}{\sqrt[4]{2\pi}\sqrt{\delta}}\exp\left(-\frac{|t-\mu|^{2}}{4\delta^{2}}\right)\end{split}$ (10)
+    $\begin{split}I\to L^{2}(I)\quad:\quad\mu\mapsto g_{\mu}(t):=\frac{1}{\sqrt[4]{2\pi}\sqrt{\delta}}\exp\left(-\frac{|t-\mu|^{2}}{4\delta^{2}}\right)\end{split}\ \,.$ (10)
     By discretizing $I$, we may sample this manifold and project it into a finite-dimensional space.
@@
 Explicitly, the evaluation function is
-    $\begin{split}L^{2}(I)\to\mathbb{R}^{d}\quad:\quad g_{\mu}(t)\mapsto(g_{\mu}(1/d),\dots,g_{\mu}(1))^{\intercal}\end{split}$ (11)
+    $\begin{split}L^{2}(I)\to\mathbb{R}^{d}\quad:\quad g_{\mu}(t)\mapsto(g_{\mu}(1/d),\dots,g_{\mu}(1))^{\intercal}\end{split}\ \,.$ (11)
     Thus, combining the above two maps (10) and (11) together produces an embedding of interval $I$ into $\mathbb{R}^{d}$ (which is equivalent to a curve in $\mathbb{R}^{d}$).
@@
 For the standard Meyer’s staircase, it follows that the expression for $x_{k}(t)$ is
-    $x_{k}(t)=\frac{1}{\sqrt[4]{2\pi}\sqrt{\delta}}\exp\left(-\frac{|k/d-t|^{2}}{4\delta^{2}}\right)$ (12)
+    $x_{k}(t)=\frac{1}{\sqrt[4]{2\pi}\sqrt{\delta}}\exp\left(-\frac{|k/d-t|^{2}}{4\delta^{2}}\right)\ \,.$ (12)
     Note that this expression differs from Definition 1 because it does not use the unit-speed parameterization.
@@
 For example, suppose we perform the linear projection of modified Meyer’s staircase onto its first few principal components: the projected image is a simple curve, and the learning problem can be significantly simplified if we study the regression problem on the projected curve.
-Hence, to test the performance algorithm for the regression problem, we aim to test curves that are so complex that there is no trivial dimension reduction, e.g.
-via standard techniques such as Principal Component Analysis.
+Hence, to test the performance algorithm for the regression problem, we aim to test curves that are so complex that there is no trivial dimension reduction, e.g., via standard techniques such as Principal Component Analysis.
     C.0.2 Meyer helix
@@
     We introduce the following curve, called Meyer helix, as an analogy of the standard and modified Meyer’s staircase:
-    $x_{k}(t)=\frac{1}{\sqrt[4]{2\pi}\sqrt{\delta_{d}}}\cos\left(ak+\frac{t-k/d}{\delta_{d}^{\prime}}\right)G\left(\frac{\left|k/d-t\right|}{\delta_{d}}\right)$ (14)
+    $x_{k}(t)=\frac{1}{\sqrt[4]{2\pi}\sqrt{\delta_{d}}}\cos\left(ak+\frac{t-k/d}{\delta_{d}^{\prime}}\right)G\left(\frac{\left|k/d-t\right|}{\delta_{d}}\right)\ \,,$ (14)
     where $\delta_{d}=(1+0.3\cos(ak))/d\ \text{ and }\ \delta_{d}^{\prime}=(1+0.3\sin(ak))/d$, with constant $a=10$ and function $G$ is taken to be Bernstein-type decay $G(z)=\exp\left(-\frac{z^{2}}{1+z}\right)$.
@@
     Figure 10: Behavior of geometric features of the Meyer helix in $\mathbb{R}^{d}$ as a function of $d$.
-    Moreover, the above properties of complexity are pretty robust: we can also choose Gaussian-type to decay $G(z)=\exp(-z^{2})$ or choose another constant in the variants (e.g.,, $a,\delta_{d}$, and $\delta_{d}^{\prime}$) and the above properties persist.
+    Moreover, the above properties of complexity are pretty robust: we can also choose Gaussian-type to decay $G(z)=\exp(-z^{2})$ or choose another constant in the variants (e.g., $a,\delta_{d}$, and $\delta_{d}^{\prime}$) and the above properties persist.
 This indicates that this collection of Meyer helix curves indeed has its complexity growing with ambient dimension $d$.
@@
 If $M\leqslant N$, then with probability at least $1-\rho$, the following statement holds: for every pair of points $x,y\in\mathcal{M}$,
-    $(1-\epsilon)\sqrt{\frac{M}{N}}\leqslant\frac{\left\|\Phi x-\Phi y\right\|_{2}}{\left\|x-y\right\|_{2}}\leqslant(1+\epsilon)\sqrt{\frac{M}{N}}$
+    $(1-\epsilon)\sqrt{\frac{M}{N}}\leqslant\frac{\left\|\Phi x-\Phi y\right\|_{2}}{\left\|x-y\right\|_{2}}\leqslant(1+\epsilon)\sqrt{\frac{M}{N}}\ \,.$
     If we apply this Johnson-Lindenstrauss lemma to a curve $\gamma$ on $\mathbb{R}^{d}$ with length $\mathrm{len}_{\gamma}$ and reach $\mathrm{reach}_{\gamma}$, then we can take ambient dimension $N=d$, the condition number $1/\tau=1/\mathrm{reach}_{\gamma}$, volume $V=\mathrm{len}_{\gamma}$, and geodesic covering regularity $R=\mathcal{O}(1)$, and thus $M=\mathcal{O}\left(\epsilon^{-2}\log(1/\rho)\log\left(\frac{d}{\epsilon}\frac{\mathrm{len}_{\gamma}}{\mathrm{reach}_{\gamma}}\right)\right)$, which suggests the possibility of dimension reduction for the curve via linear projection.